POLARIS
Network Activity Log; Last Updated: Tue Jul 10 20:49:45 EDT 2007
|
|||||
No | Date | Time UT | Location | Activity | Comments |
1 | 2021-07-16 | UWO | Hub Start | First POLARIS Hub started at UWO. | |
2 | 2021-07-27 | TYNO | New Station Start | First POLARIS station. HRD digitizer and CMG-3ESP. SOSN/POLARIS. AC Powered. | |
3 | 2021-08-20 | OTTO | New Station Start | This test station is not registered in Station Book. | |
4 | 2021-09-07 | YMBN | New Station Start | First station of NWT array. DC (Solar Panels). | |
5 | 2021-09-09 | ACKN | New Station Start | DC (Solar Panels). | |
6 | 2021-09-13 | LDGN | New Station Start | DC (Solar Panels). | |
7 | 2021-09-14 | EKTN | New Station Start | First AC station of NWT array. | |
8 | 2021-11-09 | STCO | New Station Start | SOSN/POLARIS. AC Powered. | |
9 | 2022-01-29 | WLVO | New Station Start | SOSN/POLARIS. AC Powered. | |
10 | 2022-01-30 | PKRO | New Station Start | SOSN/POLARIS. AC Powered. | |
11 | 2022-01-31 | ACTO | New Station Start | SOSN/POLARIS. AC Powered. | |
12 | 2022-02-01 | ELAO | New Station Start | First DC (Solar Panels) station of Ontario array. | |
13 | 2022-02-04 | HGVO | New Station Start | DC (Solar Panels). | |
14 | 2022-02-05 | PTCO | New Station Start | DC (Solar Panels). | |
15 | 2022-02-06 | BRCO | New Station Start | SOSN/POLARIS. AC Powered. | |
16 | 2022-02-07 | ELAO | Seismo Swap | Changed from CMG-3ESP to CMG-40T. | |
No | Date | Time UT | Location | Activity | Comments |
17 | 2022-02-07 | PTCO | Seismo Swap | Changed from CMG-3ESP to CMG-40T. | |
18 | 2022-02-08 | ACTO | Seismo Swap | Changed from CMG-3ESP to CMG-40T. | |
19 | 2022-02-08 | BRCO | Seismo Swap | Changed from CMG-3ESP to CMG-40T. | |
20 | 2022-02-08 | ELGO | Name Change | Ringbuffer and station code changed from ELAO to ELGO. | |
21 | 2022-02-21 | PTCO | Seismo Swap | Changed from CMG-40T to CMG-3ESP. | |
22 | 2022-02-22 | ACTO | Seismo Swap | Changed from CMG-40T to CMG-3ESP. | |
23 | 2022-02-22 | ELGO | Seismo Swap | Changed from CMG-40T to CMG-3ESP. | |
24 | 2022-02-26 | PGCB | New Station Start | First station of BC array. AC powered. | |
25 | 2022-02-26 | UWOO | New Station Start | AC powered. Noisy site on UWO campus. | |
26 | 2022-03-01 | BRCO | Seismo Swap | Changed from CMG-40T to CMG-3ESP. | |
27 | 2022-04-04 | ENGB | New Station Start | AC powered. | |
28 | 2022-04-06 | COQB | New Station Start | AC powered. | |
29 | 2022-04-07 | MGCB | New Station Start | DC (Solar Panels). | |
30 | 2022-04-08 | KELB | New Station Start | AC powered. | |
31 | 2022-04-10 | GOWB | New Station Start | AC powered. | |
32 | 2022-04-27 | KSVO | New Station Start | DC (Solar Panels). Installed in NY state. | |
No | Date | Time UT | Location | Activity | Comments |
33 | 2022-05-08 | UWOO | Closed | Finished with test station. | |
34 | 2021-06-26 | STCO | Hardware Swap | Replaced Cygnus 155/HRD 185 to Cygnus 182/Trident 82. | |
35 | 2021-07-18 | 21:00 | EKTN | Temp Stop | Stopped recording in EKTN ringbuffers to swap electronics. |
36 | 2021-07-18 | 21:10 | EKTN | Hardware Swap | Cygnus 156/HRD 207 replaced with Cygnus 222/Trident 143. |
37 | 2021-07-19 | PGCB | Closed | Finished with training station. | |
38 | 2021-07-19 | 00:00 | EKTN | SR Change | Sampling rate changed from 20 sps to 40 sps. |
39 | 2021-07-19 | 00:00 | P24T | New temp start | Started recording in P24T ringbuffers at EKTN. |
40 | 2021-07-19 | 15:00 | EKTN | Restart | Started recording in new ringbuffers at EKTN. |
41 | 2021-07-19 | 22:00 | P30T | New Station Start | Start recording in P30T ringbuffers at COWN. |
42 | 2021-07-19 | 23:00 | COWN | New Station Code | Start recording in COWN ringbuffers. |
43 | 2021-07-20 | 20:20 | ACKN | Temp Stop | Stopped recording in ACKN ringbuffers to swap electronics. |
44 | 2021-07-20 | 20:21 | ACKN | Hardware Swap | Cygnus 160/HRD 211 replaced with Cygnus 220/Trident 141. |
45 | 2021-07-20 | 21:00 | ACKN | SR Change | Sampling rate from to 20 sps to 40 sps |
46 | 2021-07-20 | 21:00 | P23T | New temp start | Started recording in P23T ringbuffers at ACKN. |
47 | 2021-07-20 | 21:00? | ACKN | Restart | Started recording in new ACKN ringbuffers. |
48 | 2021-07-21 | YMBN | Hardware Swap | Cygnus 157/HRD 210 replaced with Cygnus 246/Trident 235. | |
No | Date | Time UT | Location | Activity | Comments |
49 | 2021-07-21 | YMBN | Temp Stop | Stopped recording in ringbuffers YMBN to swap electronics. | |
50 | 2021-07-21 | 19:00 | P29T | New temp start | Started recording in ringbuffers P29T at YMBN. |
51 | 2021-07-21 | 19:00 | YMBN | SR Change | Sampling rate from to 20 sps to 40 sps. |
52 | 2021-07-21 | 23:00 | YMBN | Restart | Started recording in new YMBN ringbuffers. |
53 | 2021-07-22 | 17:00 | LDGN | Hardware Swap | Cygnus 158/HRD 209 replaced with Cygnus 245/Trident 237. |
54 | 2021-07-22 | 17:00 | LDGN | Temp Stop | Stopped recording in ringbuffers LDGN to swap electronics. |
55 | 2021-07-23 | 01:00 | LDGN | SR Change | Sampling rate from to 20 sps to 40 sps. |
56 | 2021-07-23 | 01:00 | P28T | New temp start | Started recording in ringbuffers P28T at LDGN. |
57 | 2021-07-23 | 15:00 | P31T | New Station Start | Start recording in P31T ringbuffers at GLWN. |
58 | 2021-07-23 | 16:00 | LDGN | Restart | started recording in new LDGN ringbuffers. |
59 | 2021-07-23 | 17:00 | GLWN | New Station Code | Start recording in GLWN ringbuffers. 2021-07-24 16:00 P25T New Station Start Start recording in P25T ringbuffers at YNEN with CMG-40T. |
60 | 2021-07-24 | 18:00 | YNEN | LVD Check | The Low Volt Disconnect set to 11.0 VDC instead of 11.5. |
61 | 2021-07-24 | 18:00 | YNEN | New Station Code | Start recording in YNEN ringbuffers. |
62 | 2021-07-24 | 22:00 | P26T | New Station Start | Start recording in P26T ringbuffers at LGSN with CMG-40T. 2021-07-25 00:00 LGSN New Station Code Start recording in LGSN ringbuffers. |
63 | 2021-07-25 | 01:00 | LGSN | New channel prototype | Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3). |
64 | 2021-07-25 | 01:00 | YNEN | New channel prototype | Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3). |
No | Date | Time UT | Location | Activity | Comments |
65 | 2021-07-26 | 00:00 | P27T | New Station Start | Start recording in P27T ringbuffers DVKN with CMG-40T. |
66 | 2021-07-26 | 03:00 | DVKN | New Station Code | Start recording in DVKN ringbuffers. |
67 | 2021-07-26 | 04:52 | DVKN | New channel prototype | Channel prototype to Broad band, Trident and CMG-40T seismometer (BHZ-2 to BHZ-3, BHN-2 to BHN-3, BHE-2 to BHE-3). |
68 | 2021-07-26 | 17:00 | P22T | New Station Start | Start recording in P22T ringbuffers at MCKN with CMG-40T. |
69 | 2021-07-26 | 20:00 | MCKN | New Station Code | Start recording in MCKN ringbuffers. |
70 | 2021-07-26 | 20:01 | MCKN | LVD Check | The Low Volt Disconnect set to 11.0 VDC instead of 11.5. |
71 | 2021-07-29 | 17:00 | P40T | New Station Start | Start recording in P40T ringbuffers at NODN. |
72 | 2021-07-29 | 20:30 | NODN | New Station Code | Start recording in NODN ringbuffers. |
73 | 2021-07-30 | 00:00 | P33T | New Station Start | Start recording in P33T ringbuffers CAMN. |
74 | 2021-07-30 | 03:45 | CAMN | New Station Code | Start recording in CAMN ringbuffers. |
75 | 2021-07-31 | 18:00 | P35T | New Station Start | Start recording in P35T ringbuffers at SNPN. |
76 | 2021-07-31 | 19:15 | SNPN | New Station Code | Start recording in SNPN ringbuffers. |
77 | 2021-07-31 | 22:00 | P34T | New Station Start | Start recording in P34T ringbuffers at IHLN. |
78 | 2021-07-31 | 23:25 | IHLN | New Station Code | Start recording in IHLN ringbuffers. |
79 | 2021-08-02 | 22:00 | P39T | New Station Start | Start recording in P39T ringbuffers BOXN. |
80 | 2021-08-02 | 23:50 | BOXN | New Station Code | Start recording in BOXN ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
81 | 2021-08-03 | 18:00 | P36T | New Station Start | Start recording in P36T ringbuffers at MLON. |
82 | 2021-08-03 | 20:30 | MLON | New Station Code | Start recording in MLON ringbuffers. |
83 | 2021-08-03 | 21:00 | P41T | New Station Start | Start recording in P41T ringbuffers at MGTN. |
84 | 2021-08-04 | 01:25 | MGTN | New Station Code | Start recording in MGTN ringbuffers. |
85 | 2021-08-04 | 13:00 | P37T | New Station Start | Start recording in P37T ringbuffers at KNDN. |
86 | 2021-08-04 | 16:20 | MGTN | Trident swap | Changed Trident from 122 to 227. |
87 | 2021-08-04 | 16:30 | KNDN | New Station Code | Start recording in KNDN ringbuffers. |
88 | 2021-08-09 | 21:30 | STCO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
89 | 2021-08-12 | 21:25 | GLWN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
90 | 2021-08-12 | 21:45 | CAMN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
91 | 2021-08-12 | 22:45 | BRCO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
92 | 2021-08-13 | 18:28 | ACKN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
93 | 2021-08-13 | 18:35 | IHLN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
94 | 2021-08-13 | 18:44 | PKRO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
95 | 2021-08-13 | 20:11 | GOWB | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
96 | 2021-08-13 | 22:50 | ACTO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
No | Date | Time UT | Location | Activity | Comments |
97 | 2021-08-13 | 23:08 | LGSN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
98 | 2021-08-13 | 23:28 | SNPN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
99 | 2021-08-13 | 23:42 | PTCO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
100 | 2021-08-14 | 20:12 | ENGB | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
101 | 2021-08-14 | 22:14 | UWO-Hub | Hub Down | UWO Hub Power Failure. Loss of at least 1/2 hour of data on the network. |
102 | 2021-08-21 | HGVO | Firmware Upgrade | Upgraded to Libra V5.20 | |
103 | 2021-08-21 | 19:18 | ELGO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
104 | 2021-08-21 | 19:24 | KNDN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
105 | 2021-08-21 | 19:45 | YNEN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
106 | 2021-08-21 | 23:45 | P44T | New Station Start | Start recording in P44T ringbuffers at MPPO. |
107 | 2021-08-21 | 23:55 | MPPO | New Station Code | Start recording in MPPO ringbuffers. |
108 | 2021-08-23 | 21:58 | BOXN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
109 | 2021-08-23 | 22:03 | COWN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
110 | 2021-08-23 | 22:17 | KELB | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
111 | 2021-08-24 | 23:00 | P45T | New Station Start | Start recording in P45T ringbuffers at BANO. |
112 | 2021-08-25 | 17:30 | BANO | New Station Code | Start recording in BANO ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
113 | 2021-08-26 | 20:30 | MGCB | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
114 | 2021-08-26 | 20:31 | DVKN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
115 | 2021-08-27 | COQB | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 | |
116 | 2021-08-27 | LDGN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 | |
117 | 2021-08-27 | 15:00 | P43T | New Station Start | Start recording in P43T ringbuffers at DELO. |
118 | 2021-08-27 | 17:15 | DELO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
119 | 2021-08-27 | 17:15 | DELO | Gain Check | Kept gain at 1.048576. |
120 | 2021-08-27 | 17:15 | DELO | New Station Code | Start recording in DELO ringbuffers. |
121 | 2021-08-27 | 21:42 | MGTN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
122 | 2021-08-28 | 18:53 | NODN | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
123 | 2021-09-03 | 18:01 | MLON | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
124 | 2021-09-06 | 16:30 | KELB | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
125 | 2021-09-06 | 16:30 | KELB | Gain Change | From 1.048576 to 0.5. Vpp unchanged at 40. |
126 | 2021-09-06 | 17:00 | P21T | New Station Start | Start recording in P21T ringbuffers at PECO. |
127 | 2021-09-06 | 17:00 | P21T | New Station Start | Start recording in P21T ringbuffers at PECO. |
128 | 2021-09-06 | 19:10 | PECO | New Station Code | Start recording in PECO ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
129 | 2021-09-06 | 19:10 | PECO | New Station Code | Start recording in PECO ringbuffers. |
130 | 2021-09-06 | 19:10 | PECO | Seismo Check | Started the station with CMG-40T. |
131 | 2021-09-07 | 22:00 | P42T | New Station Start | Start recording in P42T ringbuffers at LINO. |
132 | 2021-09-08 | 17:00 | LINO | New Station Code | Start recording in LINO ringbuffers. |
133 | 2021-09-10 | 16:00 | P47T | New Station Start | Start recording in P47T ringbuffers at BRPO. |
134 | 2021-09-10 | 16:50 | BRPO | New Station Code | Start recording in BRPO ringbuffers. |
135 | 2021-09-11 | 21:00 | P46T | New Station Start | Start recording in P46T ringbuffers at ALGO. |
136 | 2021-09-11 | 23:05 | ALGO | New Station Code | Start recording in ALGO ringbuffers. |
137 | 2021-09-12 | 09:25 | KELB | DC Filter Change | Unchanged from on at 0.001 mHz corner frequency. |
138 | 2021-09-12 | 09:25 | KELB | Gain Change | From 0.5 to 0.1. Vpp unchanged at 40. |
139 | 2021-09-12 | 09:36 | ENGB | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
140 | 2021-09-12 | 09:36 | ENGB | Gain Change | From 1.048576 to 0.1. |
141 | 2021-09-12 | 09:44 | MGCB | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
142 | 2021-09-12 | 09:44 | MGCB | Gain Change | From 1.048576 to 0.1. |
143 | 2021-09-12 | 10:12 | GOWB | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
144 | 2021-09-12 | 10:12 | GOWB | Gain Change | From 1.048576 to 0.1. |
No | Date | Time UT | Location | Activity | Comments |
145 | 2021-09-14 | 15:30 | EKTN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
146 | 2021-09-14 | 15:30-16:30 | ACKN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
147 | 2021-09-14 | 18:40 | LGSN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
148 | 2021-09-14 | 19:17 | COWN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
149 | 2021-09-14 | 21:10 | KNDN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
150 | 2021-09-14 | 21:30 | SNPN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
151 | 2021-09-14 | 21:50 | IHLN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
152 | 2021-09-14 | 22:10 | CAMN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
153 | 2021-09-14 | 22:30 | MLON | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
154 | 2021-09-14 | 22:45 | BOXN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
155 | 2021-09-14 | 23:00 | NODN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
156 | 2021-09-16 | 21:10 | BRCO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
157 | 2021-09-16 | 21:10 | BRCO | Gain Change | From 1.048576 to 0.1. |
158 | 2021-09-16 | 21:20 | MPPO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
159 | 2021-09-16 | 21:20 | MPPO | Gain Change | From 1.048576 to 0.1. |
160 | 2021-09-16 | 21:55 | STCO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
No | Date | Time UT | Location | Activity | Comments |
161 | 2021-09-16 | 21:55 | STCO | Gain Change | From 1.048576 to 0.1. |
162 | 2021-09-16 | 22:15 | PECO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
163 | 2021-09-16 | 22:15 | PECO | Gain Change | From 1.048576 to 0.1. |
164 | 2021-09-16 | 23:25 | WLVO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
165 | 2021-09-16 | 23:25 | WLVO | Gain Change | From 1.048576 to 0.1. |
166 | 2021-09-16 | 23:35 | WLVO | Firmware Upgrade | Upgraded to Libra V5.20, TS V1.20, TR V1.50 EPLD V1.08 |
167 | 2021-09-17 | 00:15 | PTCO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
168 | 2021-09-17 | 00:15 | PTCO | Gain Change | From 1.048576 to 0.1. |
169 | 2021-09-17 | 00:25 | ALGO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
170 | 2021-09-17 | 00:25 | ALGO | Gain Change | From 1.048576 to 0.1. |
171 | 2021-09-17 | 01:35 | BPRO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
172 | 2021-09-17 | 01:35 | BPRO | Gain Change | From 1.048576 to 0.1. |
173 | 2021-09-17 | 01:50 | LINO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
174 | 2021-09-17 | 01:50 | LINO | Gain Change | From 1.048576 to 0.1. |
175 | 2021-09-17 | 02:00 | PKRO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
176 | 2021-09-17 | 02:00 | PKRO | Gain Change | From 1.048576 to 0.1. |
No | Date | Time UT | Location | Activity | Comments |
177 | 2021-09-17 | 02:20 | ELGO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
178 | 2021-09-17 | 02:20 | ELGO | Gain Change | From 1.048576 to 0.1. |
179 | 2021-09-17 | 02:20 | Ontario | Resynch check | Stations in 133a-133w: Trident Resynchronization mode = Discard samples |
180 | 2021-09-19 | 16:30 | EKTN | Site Visit | All ok. |
181 | 2021-09-19 | 16:31 | EKTN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
182 | 2021-09-19 | 18:30 | DELO | DC Filter Check | Kept on at 0.001 mHz corner frequency. |
183 | 2021-09-19 | 18:30 | DELO | Gain Change | From 1.048576 to 0.1. |
184 | 2021-09-19 | 22:40 | YMBN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
185 | 2021-09-19 | 22:40 | YMBN | Repair | Wildlife damage repaired. Station re-started. |
186 | 2021-09-20 | 00:20 | DVKN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
187 | 2021-09-20 | 00:49 | SNPN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
188 | 2021-09-20 | 14:45 | LGSN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
189 | 2021-09-20 | 16:30 | MCKN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
190 | 2021-09-20 | 16:30 | MCKN | Repair | Wildlife damage repaired. Station re-started. |
191 | 2021-09-20 | 17:04 | PECO | Gain Check | Software gain was 0.1, IIR filter on at 0.001 |
192 | 2021-09-20 | 17:04 | PECO | Seismo Swap | Changed seismometer from CMG-40T to CMG-3ESP. |
No | Date | Time UT | Location | Activity | Comments |
193 | 2021-09-20 | 19:15 | ALGO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
194 | 2021-09-20 | 19:30 | PTCO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
195 | 2021-09-20 | 21:10 | MLON | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
196 | 2021-09-20 | 21:10 | MLON | Site Visit | All appear to be ok at this station. |
197 | 2021-09-20 | 22:10 | BRPO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
198 | 2021-09-20 | 22:12 | GOWB | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
199 | 2021-09-21 | 12:23 | ACKN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
200 | 2021-09-21 | 12:37 | BOXN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
201 | 2021-09-21 | 15:56 | LDGN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
202 | 2021-09-21 | 16:40 | IHLN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
203 | 2021-09-21 | 22:26 | KNDN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
204 | 2021-09-22 | 16:51 | YNEN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
205 | 2021-09-22 | 16:51 | YNEN | Repair | Wildlife damage repaired. Station re-started. |
206 | 2021-09-22 | 20:55 | COWN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
207 | 2021-09-22 | 20:55 | COWN | Re-Point | VSAT antenna repointed, tightened. Wildlife damage repaired. |
208 | 2021-09-23 | 17:39 | WLVO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
No | Date | Time UT | Location | Activity | Comments |
209 | 2021-09-23 | 17:51 | BRCO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
210 | 2021-09-23 | 18:01 | GLWN | Repair | Snowstorm at the site. RF cables repaired. |
211 | 2021-09-23 | 19:44 | NODN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
212 | 2021-09-23 | 20:12 | PKRO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
213 | 2021-09-23 | 20:14 | ENGB | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
214 | 2021-09-23 | 21:?? | MGTN | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
215 | 2021-09-23 | 21:?? | MGTN | Software bug. | Station had lost TDMA. Stopped. |
216 | 2021-09-23 | 22:15 | ACTO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
217 | 2021-09-23 | 22:25 | KELB | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
218 | 2021-09-23 | 23:40 | BANO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
219 | 2021-09-24 | KSVO | Closed | Plan to move station. | |
220 | 2021-09-24 | 09:22 | ELGO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
221 | 2021-09-24 | 09:24 | MGCB | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
222 | 2021-09-24 | 12:00 | KSVO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
223 | 2021-09-24 | 12:16 | COQB | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
224 | 2021-09-24 | 21:45 | STCO | Firmware Upgrade | Upgraded to Libra Firmware version 5.31. |
No | Date | Time UT | Location | Activity | Comments |
225 | 2021-09-25 | 01:50 | COQB | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
226 | 2021-09-25 | 01:50 | COQB | Gain Change | From 1.048576 to 0.1. |
227 | 2021-09-25 | 03:00 | HGVO | DC Filter Change | From off to on. At 0.001 mHz corner frequency. |
228 | 2021-09-25 | 03:00 | HGVO | Gain Change | From 1.048576 to 0.1. |
229 | 2021-09-25 | 16:40 | MCKN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
230 | 2021-09-25 | 17:10 | GLWN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
231 | 2021-09-25 | 18:10 | YNEN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
232 | 2021-09-25 | 18:55 | YMBN | SSPB Check | Corrected SSPB temperature profile from flat 45 dB to manufacturer specs. |
233 | 2021-10-01 | 18:10 | BANO | Hardware Swap | Cygnus 281 from 269, Trident 292 from 256 |
234 | 2021-10-11 | 18:00 | P54T | New Station Start | Start recording in P54T ringbuffers at CLPO. |
235 | 2021-10-11 | 20:15 | CLPO | New Station Code | Start recording in CLPO ringbuffers. |
236 | 2021-10-11 | 20:16 | CLPO | Seismo Check | Started with CMG-ESP (standard vault), Trillium (Test vault). |
237 | 2021-10-11 | 20:17 | CLPO | Gain Check | This station started with gain 0.1, iir on at 0.001 on both Tridents? Check with CA |
238 | 2021-10-15 | 09:?? | BRPO | Closed | Moving station to safer location. |
239 | 2021-10-24 | DVKN | VSAT icing fixed. | Icing formed on vsat feed window were cleared. Station Started transmitting again. | |
240 | 2021-10-31 | 22:00 | TYNO | Hardware Swap | Replaced Cygnus 154/HRD 184 to Cygnus 282/Trident 259. |
No | Date | Time UT | Location | Activity | Comments |
241 | 2021-11-07 | 18:35 | WLVO | Hardware Swap | Replaced Cygnus 171/Trident 63 to Cygnus 276/Trident 258. |
242 | 2021-11-09 | 23:45 | P20T | New Station Start | Start recording in P20T ringbuffers at HSLB. |
243 | 2021-11-10 | 19:00 | HLSB | New Station Code | Start recording in HLSB ringbuffers. |
244 | 2021-11-12 | 00:21 | P55T | New Station Start | Start recording in P55T ringbuffers at HRMB. |
245 | 2021-11-12 | 02:35 | HRMB | New Station Code | Start recording in HRMB ringbuffers. |
246 | 2021-11-13 | 22:00 | P58T | New Station Start | Start recording in P58T ringbuffers at PEMO. |
247 | 2021-11-13 | 22:45 | PEMO | New Station Code | Start recording in PEMO ringbuffers. |
248 | 2021-11-14 | 22:45 | P53T | New Station Start | Start recording in P53T ringbuffers at PIMB. |
249 | 2021-11-14 | 23:05 | PIMB | New Station Code | Start recording in PIMB ringbuffers. |
250 | 2021-11-15 | 20:30 | KSVO | Hardware Swap | Removed Cygnus 184, installed 171. Kept same Trident. |
251 | 2021-11-15 | 22:00 | KSVO | Reopened | Libra 5.31, Tsr1.21, Trd. 16202, epsd 123, gain 0.1, DC removal filter on at 0.001 Hz. |
252 | 2021-11-16 | 23:37 | P51T | New Station Start | Start recording in P51T ringbuffers at CLVB. |
253 | 2021-11-17 | 18:50 | CLVB | New Station Code | Start recording in CLVB ringbuffers. |
254 | 2021-11-18 | 18:00 | P32T | New Station Start | Start recording in P32T ringbuffers at ANMB. |
255 | 2021-11-18 | 19:35 | ANMB | New Station Code | Start recording in ANMB ringbuffers. |
256 | 2021-11-20 | 22:00 | P49T | New Station Start | Start recording in P49T ringbuffers at TWKB. |
No | Date | Time UT | Location | Activity | Comments |
257 | 2021-11-20 | 23:30 | TWKB | Gain Check | Starting with software gain of 0.1. |
258 | 2021-11-20 | 23:30 | TWKB | DC Filter Check | Starting with DC filter on at 0.001 Hz. |
259 | 2021-11-20 | 23:30 | TWKB | New Station Code | Start recording in TWKB ringbuffers. |
260 | 2021-11-23 | 01:33 | MGCB | Hardware Swap | Trident 080 installed, 083 removed. |
261 | 2021-11-25 | TWGB | New Station Code | Start recording in TWGB ringbuffers. | |
262 | 2021-11-25 | 19:40 | P50T | New Station Start | Start recording in P50T ringbuffers at TWGB. |
263 | 2021-11-25 | 19:40 | TWGB | Gain Check | Starting with software gain of 0.1. |
264 | 2021-11-25 | 19:40 | TWGB | DC Filter Check | Starting with DC filter on at 0.001 Hz. |
265 | 2021-11-27 | 00:10 | P56T | New Station Start | Start recording in P56T ringbuffers at TWBB. |
266 | 2021-11-27 | 00:11 | TWBB | New Station Code | Start recording in TWBB ringbuffers. |
267 | 2021-11-27 | 00:12 | TWBB | Gain Check | Starting with software gain of 0.1. |
268 | 2021-11-27 | 00:12 | TWBB | DC Filter Check | Starting with DC filter on at 0.001 Hz. |
269 | 2021-12-04 | 17:29 | P57T | New Station Start | Start recording in P57T ringbuffers at SSIB. |
270 | 2021-12-04 | 17:30 | SSIB | DC Filter Check | Starting with DC filter on at 0.001 Hz. |
271 | 2021-12-04 | 17:30 | SSIB | Gain Check | Starting with software gain of 0.1. |
272 | 2021-12-04 | 17:30 | SSIB | New Station Code | Start recording in SSIB ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
273 | 2021-12-06 | 21:30 | CLVB | Battery swap. | All six batteries swapped with new ones. All new firmware installed. |
274 | 2022-01-28 | 18:45 | COQB | Hardware Swap | Cygnus 187 replaced with 269, TRD 77 with 214. |
275 | 2022-02-14 | 16:50 | PEMO | Hardware Swap | Trident 076 replaced with TRD 122. |
276 | 2022-02-19 | 23:30 | ALGO | Hardware Swap | Cygnus 266 replaced with Cygnus 185. |
277 | 2022-02-19 | 23:30 | PEMO | Hardware Swap | Cygnus 174 replaced with Cygnus 184. |
278 | 2022-02-22 | 21:30 | P60T | New Station Start | Start recording in P60T ringbuffers at SILB. SILB New Station Code Start recording in SILB ringbuffers. |
279 | 2022-04-10 | 17:10 | WLVO | Site Visit | Maintenance visit. |
280 | 2022-04-10 | 17:10 | WLVO | VSAT Check | Realigned. Co-pol 28, X-pol 0 dB. |
281 | 2022-04-11 | TYNO | Site Visit | Maintenance visit. | |
282 | 2022-04-11 | 16:56 | ACTO | Hardware Swap | Now running CYG174 TRD77. |
283 | 2022-04-11 | 16:56 | ACTO | Site Visit | Maintenance visit. |
284 | 2022-04-15 | HGVO | Power Upgrade | Added 2 more batteries. | |
285 | 2022-04-15 | HGVO | Site Visit | Maintenance visit. | |
286 | 2022-04-15 | PTCO | Power Upgrade | Added 2 more batteries. | |
287 | 2022-04-15 | PTCO | Site Visit | Maintenance visit. | |
288 | 2022-04-15 | PTCO | Wildlife Alert | Snake visit (below battery box). Wasps in battery box. | |
No | Date | Time UT | Location | Activity | Comments |
289 | 2022-05-05 | 22:15 | PKRO | Hardware Swap | Now running CYG172 TRD66. |
290 | 2022-05-05 | 22:15 | PKRO | Site Visit | Maintenance visit. |
291 | 2022-05-06 | 14:49 | CLPO | Seismo Check | Trillium Seismometer and its Trident removed. |
292 | 2022-05-06 | 14:49 | CLPO | Site Visit | Maintenance visit. |
293 | 2022-05-06 | 14:49 | CLPO | VSAT Check | Co-pol 32, X-pol 9dB. Adjusted, co-pol 34, X-pol 2 dB. |
294 | 2022-05-06 | 17:39 | MPPO | Site Visit | Maintenance visit. Site in perfect condition. |
295 | 2022-05-06 | 17:39 | MPPO | VSAT Check | Co-pol 32, X-pol 9dB. Adjusted, co-pol 33, X-pol 3 dB. |
296 | 2022-05-07 | 17:39 | ALGO | Site Visit | Cygnus installed in battery box. 2 more batteries. GPS cable in conduit. Blanket removed from Trident. |
297 | 2022-05-07 | 17:39 | ALGO | Site Visit | Maintenance visit. |
298 | 2022-05-07 | 17:39 | ALGO | VSAT Check | Co-pol 33, X-pol 2dB. Adjustment not needed. |
299 | 2022-05-08 | 15:34 | PEMO | Site Visit | Cygnus installed in battery box. more batteries? GPS cable in conduit. Battery box dug in a bit more. |
300 | 2022-05-08 | 15:34 | PEMO | Site Visit | Maintenance visit. |
301 | 2022-05-08 | 15:34 | PEMO | VSAT Check | Co-pol 34, X-pol 4dB. Adjusted, co-pol 33, X-pol 2 dB. IFL loss set to 2 dB. |
302 | 2022-05-14 | 20:30 | P59T | New Station Start | Start recording in P59T ringbuffers at YOUB. |
303 | 2021-05-15 | 15:20 | YOUB | Hardware List | Guralp ESP T3D81 Cygnus 187 Trident 81 |
304 | 2021-05-15 | 15:20 | YOUB | New Station Code | Start recording in YOUB ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
305 | 2021-05-17 | 01:49 | P62T | New Station Start | Start recording in P62T ringbuffers at TSJB. |
306 | 2021-05-18 | 14:?? | TSJB | Hardware List | Guralp ESP T3D84 Cygnus 266 Trident 76 |
307 | 2021-05-18 | 14:?? | TSJB | New Station Code | Start recording in TSJB ringbuffers. |
308 | 2021-05-29 | 21:00 | HBDB | Hardware List | Guralp ESP T3A81?? Cygnus 394 Trident 425 |
309 | 2021-05-29 | 21:00 | HBDB | Site Geology | Soft Clay Soil |
310 | 2021-05-29 | 21:00 | HBDB | VSAT Start | X-pol 2, Co-pol 34 dB. |
311 | 2021-05-29 | 21:00 | P75T | New Station Start | Start recording in P57T ringbuffers at HBDB. |
312 | 2021-05-29 | 21:?? | HBDB | New Station Code | Start recording in HBDB ringbuffers. |
313 | 2021-06-09 | 00:21 | O81T | New Station Start | Start recording in O81T ringbuffers at SILO. |
314 | 2021-06-09 | 00:21 | Ottawa | Hub Start | Start recording at Ottawa Hub. |
315 | 2021-06-09 | 12:31 | SILO | Hardware List | Cyg 378, TRD365, ESP T3E88. SPS 100, Gain 1. |
316 | 2021-06-09 | 12:31 | SILO | New Station Code | Start recording in SILO ringbuffers. |
317 | 2021-06-09 | 12:31 | SILO | Site Geology | Hard solid bedrock. |
318 | 2021-06-09 | 12:31 | SILO | VSAT Start | X-pol 0, Co-pol 27 dB. |
319 | 2021-06-10 | 21:45 | O72T | New Station Start | Start recording in O72T ringbuffers at VIMO. |
320 | 2021-06-11 | 13:00 | VIMO | Hardware List | Cyg 364, Trd 370, ESP T3F09. SPS 100, Gain 1. |
No | Date | Time UT | Location | Activity | Comments |
321 | 2021-06-11 | 13:00 | VIMO | New Station Code | Start recording in VIMO ringbuffers. |
322 | 2021-06-11 | 13:00 | VIMO | Site Geology | Limestone. |
323 | 2021-06-11 | 13:00 | VIMO | VSAT Start | X-pol 0, Co-pol 32 dB. |
324 | 2021-06-14 | 19:10 | O71T | New Station Start | Start recording in O71T ringbuffers at MUMO. |
325 | 2021-06-15 | 00:01 | MUMO | Hardware List | Cyg390 Trd364 3ESP T3F10 check w CA. SPS 100, Gain 1. |
326 | 2021-06-15 | 00:01 | MUMO | New Station Code | Start recording in MUMO ringbuffers. |
327 | 2021-06-15 | 00:01 | MUMO | Site Geology | Solid bedrock, about 100 m off the end of the runway. |
328 | 2021-06-15 | 00:01 | MUMO | VSAT Start | X-pol 0 Co-pol 35 dB |
329 | 2021-06-17 | HSLB | Service Visit | ||
330 | 2021-06-17 | 18:34 | ENGB | Gain Change | Software gain is set to 1.0 now but sps kept at 100 as before. |
331 | 2021-06-17 | 18:34 | ENGB | Hardware Swap | Installed Cyg398, Trd418, kept seismo the same, T3B20. |
332 | 2021-06-17 | 18:34 | ENGB | Hub Swap | Station is served by Ottawa Hub. |
333 | 2021-06-18 | ANMB | Service Visit | All ok. Fixed fence. | |
334 | 2021-06-18 | PIMB | Service Visit | All OK. Screw fixed in junction box. | |
335 | 2021-06-18 | 19:58 | O77T | New Station Start | Start recording in O77T ringbuffers at OTRO. |
336 | 2021-06-18 | 21:45 | COQB | Service Visit | Grounded both Cygnus & Trident. All else ok. |
No | Date | Time UT | Location | Activity | Comments |
337 | 2021-06-18 | 23:47 | OTRO | Hardware List | Cyg391 Trd466 CMG-40T T4767. SPS 40, Gain 1. |
338 | 2021-06-18 | 23:47 | OTRO | New Station Code | Start recording in OTRO ringbuffers. |
339 | 2021-06-18 | 23:47 | OTRO | Site Geology | Glacial Moraine (similar to NWT) about 100 m from river |
340 | 2021-06-18 | 23:47 | OTRO | VSAT Start | Xpol 5, Co-pol 35 |
341 | 2021-06-19 | 00:33 | OTRO | SR Change | SR changed to 100 sps, Gain 1. |
342 | 2021-06-20 | 15:04 | O73T | New Station Start | Start recording in O73T ringbuffers at MALO. |
343 | 2021-06-20 | 19:18 | MALO | Hardware List | Cyg392 Trd361 CMG-40T T4774, SPS 40, Gain 1. |
344 | 2021-06-20 | 19:18 | MALO | New Station Code | Start recording in MALO ringbuffers. |
345 | 2021-06-20 | 19:18 | MALO | SR Change | SR change to100 sps did not register. SR is still at 40. |
346 | 2021-06-20 | 19:18 | MALO | Site Geology | Bedrock. About 1.5 km from lake. |
347 | 2021-06-20 | 19:18 | MALO | VSAT Start | Xpol 1, Co-pol 31 |
348 | 2021-06-21 | 19:56 | O70T | New Station Start | Start recording in O70T ringbuffers at KILO. |
349 | 2021-06-22 | 13:04 | KILO | Hardware List | CYG397 TRD369 CMG-40T T4168. SPS 100, Gain 1. |
350 | 2021-06-22 | 13:04 | KILO | New Station Code | Start recording in KILO ringbuffers. |
351 | 2021-06-22 | 13:04 | KILO | Site Geology | Bedrock |
352 | 2021-06-22 | 13:04 | KILO | VSAT Start | x/c ?? |
No | Date | Time UT | Location | Activity | Comments |
353 | 2021-06-23 | 20:35 | O74T | New Station Start | Start recording in O74T ringbuffers at SUNO. |
354 | 2021-06-23 | 21:15 | SUNO | New Station Code | Start recording in SUNO ringbuffers. |
355 | 2021-06-23 | 21:15 | SUNO | Hardware List | CYG388 TRD360 CMG-40T T4776, SPS 100, Gain 1. |
356 | 2021-06-23 | 21:15 | SUNO | Site Geology | Bedrock |
357 | 2021-06-23 | 21:15 | SUNO | VSAT Start | co-pol 34 x-pol 2 |
358 | 2021-06-27 | 16:04 | MGCB | Gain Change | Software gain at 1.0 (IA: check with IAK records) |
359 | 2021-06-27 | 16:04 | MGCB | Hardware Swap | Cygn180 replaced with Cyg327. All othe hardware the same. |
360 | 2021-07-06 | 16:22 | GLWN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
361 | 2021-07-06 | 16:22 | GLWN | Feed Window | Small bite in Feed Window repaired. |
362 | 2021-07-06 | 16:22 | GLWN | Firmware Upgrade | Libra 5.6201, TS 1.23, Trd 1.7201, EPLD 1.23 |
363 | 2021-07-06 | 16:22 | GLWN | Hardware Swap | Cyg 251replaced with Cyg421. All other hardware the same. |
364 | 2021-07-06 | 16:22 | GLWN | Hub Swap | Served by Ottawa Hub. |
365 | 2021-07-06 | 16:22 | GLWN | Re-Point | Co-pol 20 dB. Low signal. |
366 | 2021-07-06 | 16:22 | GLWN | Service Visit | Could not communicate with Cygnus 251. Something wrong with Cygnus. Need to be checked later. SSPB was loose. |
367 | 2021-07-06 | 20:11 | YNEN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
368 | 2021-07-06 | 20:11 | YNEN | Firmware Upgrade | Libra 5.6201, TS 1.23, Trd 1.7201, EPLD 1.23 |
No | Date | Time UT | Location | Activity | Comments |
369 | 2021-07-06 | 20:11 | YNEN | Hardware Swap | New SSPB installed. S/N H0017A. |
370 | 2021-07-06 | 20:11 | YNEN | Hub Swap | Served by Ottawa Hub. |
371 | 2021-07-06 | 20:11 | YNEN | Re-Point | Co-pol 25, X-pol 0 dB. |
372 | 2021-07-06 | 20:11 | YNEN | SOH Monitor | Working OK. |
373 | 2021-07-06 | 20:11 | YNEN | Service Visit | Wildlife damage to SSPB. |
374 | 2021-07-06 | 22:00 | COWN | Service Visit | Site is working. Stopped for service. |
375 | 2021-07-06 | 23:15 | COWN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
376 | 2021-07-06 | 23:15 | COWN | Feed Window | New Feed Window installed. |
377 | 2021-07-06 | 23:15 | COWN | Firmware Upgrade | Libra 5.6201 in both partitions. |
378 | 2021-07-06 | 23:15 | COWN | Re-Point | Co-pol 28, X-pol 0 dB. |
379 | 2021-07-07 | 17:14 | Ottawa | Hub Down | Hub and Naqs shut down 17:14-17:35 to install AC switch on Inverter-charger. |
380 | 2021-07-07 | 18:11 | MCKN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
381 | 2021-07-07 | 18:11 | MCKN | Firmware Upgrade | Libra 5.6201 installed. |
382 | 2021-07-07 | 18:11 | MCKN | Re-Point | Co-pol 22, X-pol 0 dB. IFL loss at 3 db. |
383 | 2021-07-07 | 18:11 | MCKN | Service Visit | Cables were damaged but feed window and sspb ok. |
384 | 2021-07-07 | 18:30 | MCKN | Hub Swap | Served by Ottawa Hub. |
No | Date | Time UT | Location | Activity | Comments |
385 | 2021-07-07 | 21:34 | NODN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
386 | 2021-07-07 | 21:34 | NODN | Feed Window | Feed window replaced. |
387 | 2021-07-07 | 21:34 | NODN | Firmware Upgrade | All new firware (due to out-of-sync LibGui and firmware, initially had trouble sending test signal) |
388 | 2021-07-07 | 21:34 | NODN | Hardware Swap | Cygnus 395 installed, 265 removed. |
389 | 2021-07-07 | 21:34 | NODN | Re-Point | Co-pol 38, X-pol 2 dB. Strong Ontario-like signal. |
390 | 2021-07-07 | 21:34 | NODN | SSPB Check | New 2 watts SSPB SN 16464 installed. Profile (25/46.93, -40/45.73,50/47.73). |
391 | 2021-07-07 | 21:34 | NODN | Service Visit | Damage to feed horn mount. Looked like bear visit. |
392 | 2021-07-07 | 22:49 | NODN | Hub Swap | Served by Ottawa Hub. |
393 | 2021-07-08 | 16:00 | LGSN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
394 | 2021-07-08 | 16:00 | LGSN | Feed Window | Feed window replaced. |
395 | 2021-07-08 | 16:00 | LGSN | Firmware Upgrade | Libra 5.6201 installed. |
396 | 2021-07-08 | 16:00 | LGSN | Re-Point | Co-pol 25, X-pol 0 dB. |
397 | 2021-07-08 | 16:00 | LGSN | Service Visit | Site was working OK. |
398 | 2021-07-08 | 18:52 | MLON | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
399 | 2021-07-08 | 18:52 | MLON | Firmware Upgrade | Libra 5.6201 installed. |
400 | 2021-07-08 | 18:52 | MLON | Re-Point | Co-pol 22 X-pol 2 dB. |
No | Date | Time UT | Location | Activity | Comments |
401 | 2021-07-08 | 18:52 | MLON | Service Visit | Site was working OK. |
402 | 2021-07-08 | 22:10 | DVKN | Before Re-Point | Before any change to the station, it was noted that the elevation nut on antenna was loose. Before any re-adjustment, co-pol was 28 and x-pol was 5 dB. This could explain bad TX in recent times. |
403 | 2021-07-08 | 22:10 | DVKN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
404 | 2021-07-08 | 22:10 | DVKN | Firmware Upgrade | Libra 5.6201 installed. SSPB S/N H0016A. |
405 | 2021-07-08 | 22:10 | DVKN | Re-Point | After antenna adjustments and re-pointing, co-pol was 38 with x-pol of 2 dB. Much improved. |
406 | 2021-07-08 | 22:10 | DVKN | Seismo Check | 40T seismometer was off-level probably due to shifting of ground. Re-leveled. |
407 | 2021-07-08 | 22:10 | DVKN | Service Visit | Site was working generally OK. |
408 | 2021-07-09 | 19:30 | LDGN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
409 | 2021-07-09 | 19:30 | LDGN | Feed Window | Replaced. |
410 | 2021-07-09 | 19:30 | LDGN | Firmware Upgrade | Libra 5.6201 installed and other firmware. |
411 | 2021-07-09 | 19:30 | LDGN | Gain Check | Software gain 1.0, DC removal filter on at 1 mHz, Input range 40 vPP. |
412 | 2021-07-09 | 19:30 | LDGN | Hardware Swap | Trident 423 installed, 237 removed. |
413 | 2021-07-09 | 19:30 | LDGN | Hub Swap | Served by Ottawa Hub. |
414 | 2021-07-09 | 19:30 | LDGN | Re-Point | SSPB was not replaced (no spare available). Cleaned-up old SSPB installed. Co-pol 17-20, X-pol 0 dB. IP 245. Pinged ok. |
415 | 2021-07-09 | 19:30 | LDGN | Service Visit | Site was down. Memberane on feed window was damaged. SSPB full of water. |
416 | 2021-07-09 | 20:42 | YMBN | Service Visit | Site was working OK. Stopped for service. |
No | Date | Time UT | Location | Activity | Comments |
417 | 2021-07-09 | 21:48 | YMBN | Before Re-Point | Before any re-adjustment, co-pol was 27 and x-pol was 0 dB, indicating stable antenna base. |
418 | 2021-07-09 | 21:48 | YMBN | Clean up | Garbage and old battery box prepared for removal. |
419 | 2021-07-09 | 21:48 | YMBN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
420 | 2021-07-09 | 21:48 | YMBN | Firmware Upgrade | Libra 5.6201 installed. SSPB S/N H0013A. |
421 | 2021-07-09 | 21:48 | YMBN | Re-Point | After re-pointing, co-pol was 28 with x-pol of 0 dB. Not much improved as expected. |
422 | 2021-07-09 | 21:48 | YMBN | Seismo Check | Blue tarp was off the vault. Buried the vault. No mass-centering needed. Calibrated ok. |
423 | 2021-07-11 | 22:00 | O79T | New Station Start | Start recording in O79T ring buffers at LUPN. Station co-installed with Canadian Space Agency (CSA) CANOPUS station at Contort Lake, Lupin where Magnetometer and Riometer are operating. |
424 | 2021-07-11 | 22:22 | LUPN | Hardware List | CYG389 TRD437 CMG-3ESP T3E89, SPS 40, Gain 1., DC filter on at 1 mHz. |
425 | 2021-07-11 | 22:22 | LUPN | New Station Code | Start recording in LUPN ringbuffers. |
426 | 2021-07-11 | 22:22 | LUPN | SSPB Check | 2 Watts SSPB, S/N 16462 |
427 | 2021-07-11 | 22:22 | LUPN | Site Geology | |
428 | 2021-07-11 | 22:22 | LUPN | VSAT Start | co-pol 25 x-pol 0 |
429 | 2021-07-12 | 19:17 | LUPN | DC Filter Check | DC filter is turned off. |
430 | 2021-07-12 | 21:31 | LUPN | Restart | Initial installation of battery charger at the back of the antenna created spikes on CSA CANOPUS Riometer nearby. The battery charger was moved away from antenna to inside the CSA building and station restarted at 21:31 UT. Data prior to this time are disrupted frequently. |
431 | 2021-07-13 | 18:30? | EKTN | Service Visit | Staion was not working OK even after a previous service visit last week. The LNB cable was found to be loose. All cables re-checked. |
432 | 2021-07-13 | 19:20 | EKTN | Firmware Upgrade | Libra 5.6201, TS 1.23, TRD 1.701 running. |
No | Date | Time UT | Location | Activity | Comments |
433 | 2021-07-13 | 19:20 | EKTN | Gain Check | Software gain 1.0, DC removal filter off, Input range 40 Vpp. |
434 | 2021-07-13 | 19:20 | EKTN | Hardware Swap | Trident 417 installed, 143 removed. 2 Watts SSPB installed, S/N 16465. |
435 | 2021-07-13 | 19:20 | EKTN | Hub Swap | Served by Ottawa Hub. |
436 | 2021-07-13 | 19:20 | EKTN | Re-Point | 24 dB co-pol 0 x-pol. |
437 | 2021-07-13 | 19:30 | EKTN | Feed Horn Box | New Feed Horn protective box installed. RF cables near feed horn placed in conduit. |
438 | 2021-07-14 | 20:10 | ACKN | Clean up | Garbage removed from site. |
439 | 2021-07-14 | 20:10 | ACKN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
440 | 2021-07-14 | 20:10 | ACKN | Feed Window | Replaced. |
441 | 2021-07-14 | 20:10 | ACKN | Firmware Upgrade | Libra 5.6201, TS 1.23, TRD 1.701 running. |
442 | 2021-07-14 | 20:10 | ACKN | Gain Check | Software gain 1.048576, DC removal filter off, Input range 40 Vpp. |
443 | 2021-07-14 | 20:10 | ACKN | Hub Swap | Served by Ottawa Hub. |
444 | 2021-07-14 | 20:10 | ACKN | Re-Point | 35 dB co-pol 1 x-pol. |
445 | 2021-07-14 | 20:10 | ACKN | SSPB Check | Two watts sspb installed. S/N 16461. |
446 | 2021-07-14 | 20:10 | ACKN | Service Visit | Station was down. GPS cable was chewed up, feed window memberane damaged, water in feed horn assembly. |
447 | 2021-07-17 | 20:55 | MALO | Service Visit | Station was OK but Hub communication to it was not possible due to unknown problem. Power was turned off. |
448 | 2021-07-17 | 21:01 | MALO | Restart | Restarted without any hardware change. |
No | Date | Time UT | Location | Activity | Comments |
449 | 2021-07-17 | 21:15 | MALO | Hub Login | Hub communication restored. Logged in an change sample rate from 40 to 100 sps on Cygnus 392. |
450 | 2021-07-17 | 22:00 | MALO | Temp Stop | Station stopped. End of Cygnus 392. |
451 | 2021-07-17 | 22:22 | MALO | Hardware Swap | Cygnus 422 installed, 392 removed. |
452 | 2021-07-17 | 22:34 | MALO | Temp Stop | Stopped to re-connect temperature sensor. |
453 | 2021-07-17 | 22:39 | MALO | Restart | Restarted. |
454 | 2021-07-17 | 22:52 | MALO | Gain Check | Software gain 1.0 , DC removal filter ON at 1 mHz, Input range 40 Vpp. |
455 | 2021-07-17 | 22:52 | MALO | Sample Rate Change | Sample rate is set to 100 SPS, on Cygnus 422. This is start of good data with new sample rate. |
456 | 2021-07-18 | 03:02 | DSMN | Hardware List | CYG396 TRD513 CMG-3ESP T3F12, SPS 40, Gain 1., DC filter On at 1 mHz. |
457 | 2021-07-18 | 03:02 | DSMN | New Station Code | Start recording in DSMN ringbuffers. |
458 | 2021-07-18 | 03:02 | DSMN | New Station Start | Start recording in DSMN ringbuffers at DSMN. No data from this location in temp name O78T which was previously allocated to electronics. |
459 | 2021-07-18 | 03:02 | DSMN | SSPB Check | 2 Watts SSPB, S/N 16463 |
460 | 2021-07-18 | 03:02 | DSMN | Site Geology | Solid Bedrock. Station rock-bolted. Discovery camp nearby, about 300 yards, generator running. |
461 | 2021-07-18 | 03:02 | DSMN | VSAT Start | Co-pol 35 X-pol 3 |
462 | 2021-07-18 | 13:28 | DSMN | DC Filter Check | DC filter is turned off. |
463 | 2021-07-18 | 16:00 | DSMN | Seismo Check | Vertical component calibration signal clipped. Probably due to bad calibration coil. Seismometer locked/unlocked a few times. Calibration appear to unsettle seismometer. DO NOT Calibrate this seismometer until further notice. Noisy test data between 16:00 and 17:30. |
464 | 2021-07-19 | 17:28 | IHLN | Feed Horn Box | New Feed Horn protective box installed. New cables and conduits installed. |
No | Date | Time UT | Location | Activity | Comments |
465 | 2021-07-19 | 17:28 | IHLN | Feed Window | Broken, replaced. |
466 | 2021-07-19 | 17:28 | IHLN | Firmware Upgrade | Libra 5.6201 installed. |
467 | 2021-07-19 | 17:28 | IHLN | Re-Point | Co-pol 25, x-pol 0 dB. |
468 | 2021-07-19 | 17:28 | IHLN | Service Visit | Site was working generally OK. Stopped for service. |
469 | 2021-07-19 | 19:38 | IHLN | Restart | Calibrated OK. |
470 | 2021-07-19 | 19:54 | SNPN | Before Re-Point | Before any re-adjustment, co-pol was 28 and x-pol was 0 dB, indicating stable antenna base. |
471 | 2021-07-19 | 19:54 | SNPN | Feed Horn Box | New Feed Horn protective box installed. New cables and conduits installed. |
472 | 2021-07-19 | 19:54 | SNPN | Firmware Upgrade | Libra 5.6201 installed. |
473 | 2021-07-19 | 19:54 | SNPN | Re-Point | After re-pointing, co-pol was 32 with x-pol of 1 dB. |
474 | 2021-07-19 | 19:54 | SNPN | Service Visit | Site was working OK. Stopped for service. Solar panel rack was loose, fixed. |
475 | 2021-07-19 | 20:45 | SNPN | Restart | Calibrated ok, mass-centred ok. |
476 | 2021-07-19 | 21:47 | CAMN | Before Re-Point | Before any re-adjustment, co-pol was 35 and x-pol was 15 dB, indicating antenna move. |
477 | 2021-07-19 | 21:47 | CAMN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
478 | 2021-07-19 | 21:47 | CAMN | Firmware Upgrade | Libra 5.6201 installed? |
479 | 2021-07-19 | 21:47 | CAMN | Re-Point | After re-pointing, co-pol was 35 with x-pol of 2 dB. Improved. |
480 | 2021-07-19 | 21:47 | CAMN | Service Visit | Site was working OK. Stopped for service. |
No | Date | Time UT | Location | Activity | Comments |
481 | 2021-07-19 | 22:54 | CAMN | Restart | Calibrated & mass-centred ok. |
482 | 2021-07-20 | 00:19 | BOXN | Before Re-Point | Before any re-adjustment, co-pol was 34 and x-pol was 10 dB, indicating antenna move. |
483 | 2021-07-20 | 00:19 | BOXN | Feed Horn Box | New Feed Horn protective box installed. New conduits installed. |
484 | 2021-07-20 | 00:19 | BOXN | Firmware Upgrade | Libra 5.6201 installed? |
485 | 2021-07-20 | 00:19 | BOXN | Re-Point | After re-pointing, co-pol was 38 with x-pol of 5 dB. Much improved. |
486 | 2021-07-20 | 00:19 | BOXN | Service Visit | Site was working OK. Stopped for service. Solar panel rack was loose, fixed. |
487 | 2021-07-20 | 01:29 | BOXN | Restart | Calibrated and mass-centred ok. |
488 | 2021-07-20 | 15:31 | MGTN | Before Re-Point | Before any re-adjustment, co-pol was low at 15 dB and no x-pol detected, indicating unstable antenna base. |
489 | 2021-07-20 | 15:31 | MGTN | Feed Horn Box | New Feed Horn protective box installed. New cables and conduits installed. |
490 | 2021-07-20 | 15:31 | MGTN | Firmware Upgrade | Libra 5.6201 installed? |
491 | 2021-07-20 | 15:31 | MGTN | Re-Point | After re-pointing, co-pol was 21 with x-pol of 0 dB, improved as expected. |
492 | 2021-07-20 | 15:31 | MGTN | SSPB Check | 2 Watts SSPB, S/N 16460 installed. Profile (25/46.9, -40/46.0, 50/47.5 entered). |
493 | 2021-07-20 | 15:31 | MGTN | Service Visit | Site was working OK. Stopped for service. SSPB cable was loose, connector fell off. Minor damage to tarp on battery box. |
494 | 2021-07-20 | 16:39 | MGTN | Restart | Calibrated & mass-centred ok. |
495 | 2021-07-20 | 19:19 | KNDN | DC Filter Check | DC removal filter is off. |
496 | 2021-07-20 | 19:19 | KNDN | Feed Horn Box | New Feed Horn protective box installed. New calbes and conduits installed. |
No | Date | Time UT | Location | Activity | Comments |
497 | 2021-07-20 | 19:19 | KNDN | Feed Window | Replaced. |
498 | 2021-07-20 | 19:19 | KNDN | Firmware Upgrade | Libra 5.6201, TS 1.23, Trident 1.7201, 1.23. |
499 | 2021-07-20 | 19:19 | KNDN | Gain Change | Software gain changed to 1.0. |
500 | 2021-07-20 | 19:19 | KNDN | Hardware Swap | Cygnus 419 installed, 260 removed. |
501 | 2021-07-20 | 19:19 | KNDN | Hub Swap | Served by Ottawa Hub. |
502 | 2021-07-20 | 19:19 | KNDN | Re-Point | Co-pol 30, X-pol 0 dB. |
503 | 2021-07-20 | 19:19 | KNDN | Restart | Calibrated & mass-centred ok. |
504 | 2021-07-20 | 19:19 | KNDN | Service Visit | Site was down (dead Cygnus). Pinhole in feed window memberane. |
505 | 2021-07-25 | 17:25 | CLWO | DC Filter Check | On at 0.001 Hz. |
506 | 2021-07-25 | 17:25 | CLWO | Gain Check | Station Started at software gain of 0.5, sample rate 100 sps. |
507 | 2021-07-25 | 17:25 | CLWO | Hardware List | CYG183 TRD85 CMG-40T 4595, SSPB A0306A |
508 | 2021-07-25 | 17:25 | CLWO | New Station Code | Start recording in CLWO ringbuffers. |
509 | 2021-07-25 | 17:25 | CLWO | Site Geology | Limestone bedrock beneath shallow overburden. |
510 | 2021-07-25 | 17:25 | CLWO | VSAT Start | Co-pol 25 X-pol 0 |
511 | 2021-07-25 | 17:25 | P82T | New Station Start | Start recording in P82T ringbuffers at CLWO. |
512 | 2021-07-27 | 16:47 | P83T | DC Filter Check | On at 0.001 Hz. |
No | Date | Time UT | Location | Activity | Comments |
513 | 2021-07-27 | 16:47 | P83T | Gain Check | Station Started at software gain of 0.5, sample rate 100 sps. |
514 | 2021-07-27 | 16:47 | P83T | Hardware List | CYG417 TRD501 CMG-40T 4567, SSPB A0305A |
515 | 2021-07-27 | 16:47 | P83T | New Station Start | Start recording in P83T ringbuffers at KLBO. |
516 | 2021-07-27 | 16:47 | P83T | Site Geology | Bedrock |
517 | 2021-07-27 | 16:47 | P83T | VSAT Start | Co-pol 25 X-pol 0 |
518 | 2021-07-28 | 13:45 | KLBO | New Station Code | Start recording in KLBO ringbuffers. |
519 | 2021-07-29 | 14:29 | P81T | New Station Start | Start recording in P81T ringbuffers at TOBO. |
520 | 2021-07-29 | 15:43 | TOBO | DC Filter Check | On at 0.001 Hz. |
521 | 2021-07-29 | 15:43 | TOBO | Gain Check | Station Started at software gain of 0.5, sample rate 100 sps. |
522 | 2021-07-29 | 15:43 | TOBO | Hardware List | CYG413 TRD363 CMG-40T 4401, SSPB A0308A |
523 | 2021-07-29 | 15:43 | TOBO | New Station Code | Start recording in TOBO ringbuffers. |
524 | 2021-07-29 | 15:43 | TOBO | Site Geology | Limestone bedrock. |
525 | 2021-07-29 | 15:43 | TOBO | VSAT Start | Co-pol 28 X-pol 2 |
526 | 2021-08-11 | 18:31 | ENGB | DC Filter Check | Unchanged from ON at 0.001 Hz. |
527 | 2021-08-11 | 18:31 | ENGB | Gain Change | Software gain changed from 1.0 to 0.5. |
528 | 2021-08-11 | 18:54 | MGCB | DC Filter Check | Unchanged from ON at 0.001 Hz. |
No | Date | Time UT | Location | Activity | Comments |
529 | 2021-08-11 | 18:54 | MGCB | Gain Change | Software gain changed from 0.1 to 0.5. |
530 | 2021-08-11 | 19:24 | SILO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
531 | 2021-08-11 | 19:24 | SILO | Gain Check | Software gain unchanged from 1.0. |
532 | 2021-08-11 | 19:30 | VIMO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
533 | 2021-08-11 | 19:30 | VIMO | Gain Check | Software gain unchanged from 1.0. |
534 | 2021-08-11 | 19:35 | MUMO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
535 | 2021-08-11 | 19:35 | MUMO | Gain Check | Software gain unchanged from 1.0. |
536 | 2021-08-11 | 19:42 | MALO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
537 | 2021-08-11 | 19:42 | MALO | Gain Check | Software gain unchanged from 1.0. |
538 | 2021-08-11 | 19:46 | KILO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
539 | 2021-08-11 | 19:46 | KILO | Gain Check | Software gain unchanged from 1.0. |
540 | 2021-08-11 | 19:50 | SUNO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
541 | 2021-08-11 | 19:50 | SUNO | Gain Check | Software gain unchanged from 1.0. |
542 | 2021-08-11 | 19:56 | OTRO | DC Filter Check | Unchanged from ON at 0.001 Hz. |
543 | 2021-08-11 | 19:56 | OTRO | Gain Check | Software gain unchanged from 1.0. |
544 | 2021-08-12 | 11:55 | GLWN | DC Filter Check | Unchanged from Off. |
No | Date | Time UT | Location | Activity | Comments |
545 | 2021-08-12 | 11:55 | GLWN | Gain Change | Software gain changed from 1.048576 to 1.0. |
546 | 2021-08-12 | 12:18 | YNEN | DC Filter Check | Unchanged from Off. |
547 | 2021-08-12 | 12:18 | YNEN | Gain Change | Software gain changed from 1.048576 to 1.0. |
548 | 2021-08-12 | 12:30 | MCKN | DC Filter Check | Unchanged from Off. |
549 | 2021-08-12 | 12:30 | MCKN | Gain Change | Software gain changed from 1.048576 to 1.0. |
550 | 2021-08-12 | 12:40 | NODN | DC Filter Check | Unchanged from Off. |
551 | 2021-08-12 | 12:40 | NODN | Gain Change | Software gain changed from 1.048576 to 1.0. |
552 | 2021-08-12 | 13:00 | LDGN | DC Filter Change | Changed from On to Off. |
553 | 2021-08-12 | 13:00 | LDGN | Gain Check | Unchanged from 1.0. |
554 | 2021-08-12 | 13:21 | EKTN | DC Filter Check | Unchanged from Off. |
555 | 2021-08-12 | 13:21 | EKTN | Gain Check | Software gain unchanged from 1. |
556 | 2021-08-12 | 13:37 | DSMN | DC Filter Check | Unchanged from Off. |
557 | 2021-08-12 | 13:37 | DSMN | Gain Check | Software gain unchanged from 1. |
558 | 2021-08-12 | 17:57 | KNDN | DC Filter Check | Unchanged from Off. |
559 | 2021-08-12 | 17:57 | KNDN | Gain Check | Software gain unchanged from 1. |
560 | 2021-08-12 | 18:06 | LUPN | DC Filter Check | Unchanged from Off. |
No | Date | Time UT | Location | Activity | Comments |
561 | 2021-08-12 | 18:06 | LUPN | Gain Check | Software gain unchanged from 1. |
562 | 2021-08-12 | 18:38 | ACKN | DC Filter Check | Unchanged from Off. |
563 | 2021-08-12 | 18:38 | ACKN | Gain Change | Software gain changed from 1.048576 to 1.0. |
564 | 2021-08-12 | 19:20 | KILO | Seismo Check | Confirmed to be T4168. |
565 | 2021-08-21 | 16:17 | P84T | DC Filter Check | On at 0.001 Hz. |
566 | 2021-08-21 | 16:17 | P84T | Gain Check | Station Started at software gain of 0.5, sample rate 100 sps. |
567 | 2021-08-21 | 16:17 | P84T | Hardware List | CYG180 TRD319 CMG-3ESP T3D83, SSPB A0303A |
568 | 2021-08-21 | 16:17 | P84T | New Station Start | Start recording in P84T ringbuffers at PLIO. |
569 | 2021-08-21 | 16:17 | P84T | Site Geology | Limestone with shallow overburden. |
570 | 2021-08-21 | 16:17 | P84T | VSAT Start | Co-pol 25 X-pol 0 |
571 | 2021-08-22 | 16:00 | PLIO | New Station Code | Start recording in PLIO ringbuffers. |
572 | 2021-08-25 | 17:15 | TWKB | Service Visit | Station shut down to check seismometer orientation. |
573 | 2021-08-25 | 18:00 | TWKB | Orientation Change | Seismometer orientation was 179 degrees worng. Corrected. |
574 | 2021-08-25 | 18:15 | TWKB | Restart | Restarted. |
575 | 2021-08-25 | 20:36 | KLBO | Service Visit | Station was not transmitting, Cygnus found in Test mode. |
576 | 2021-08-25 | 21:30 | KLBO | Seismo Swap | CMG-40T T4567 removed, 40T T4317 installed. |
No | Date | Time UT | Location | Activity | Comments |
577 | 2021-08-25 | 21:30 | KLBO | VSAT Start | Co-pol 29, X-pol 0. |
578 | 2021-08-26 | 12:25 | KLBO | Service Visit | Stopped to replace Trident due to noise on n/s comp. |
579 | 2021-08-26 | 12:40 | KLBO | DC Filter Check | Unchanged at 0.001 Hz and on. |
580 | 2021-08-26 | 12:40 | KLBO | Gain Check | Unchanged at 0.5, sample rate 100. |
581 | 2021-08-26 | 12:40 | KLBO | Hardware Swap | Trident 490 installed. TRD501 removed. Restarted. Still noisy. |
582 | 2021-08-26 | 12:58 | KLBO | Seismo Swap | CMG-40T T4317 removed, 40T T4567 installed. |
583 | 2021-08-27 | 16:30 | HGVO | DC Filter Check | Unchanged at 0.001 Hz and on. |
584 | 2021-08-27 | 16:30 | HGVO | Gain Check | Unchanged at 0.5, sample rate 100. |
585 | 2021-08-27 | 16:30 | HGVO | Hardware Swap | CYG175, TRD073 removed, CYG173, TRD059 installed. |
586 | 2021-08-27 | 16:30 | HGVO | Service Visit | |
587 | 2021-09-24 | 20:40 | O82T | Hardware List | Cyg 260, Trd 143, CMG-40T T4769, SSPB H0028A. |
588 | 2021-09-24 | 20:40 | O82T | New Station Start | Start recording in O82T ringbuffers at GBLN. |
589 | 2021-09-24 | 20:40 | O82T | Site Geology | |
590 | 2021-09-24 | 20:40 | O82T | VSAT Start | Co-pol 20 X-pol 0 |
591 | 2021-09-24 | 21:28 | GBLN | DC Filter Check | Started with DC filter off. |
592 | 2021-09-24 | 21:28 | GBLN | Gain Check | Started with software gain of 1.0 at 40 sps. |
No | Date | Time UT | Location | Activity | Comments |
593 | 2021-09-24 | 21:28 | GBLN | New Station Start | Start recording in GBLN ringbuffers. |
594 | 2021-09-27 | 18:10 | O84T | Hardware List | Cyg 251, Trd 501, CMG-40T T4316, SSPB A0147A. |
595 | 2021-09-27 | 18:10 | O84T | New Station Start | Start recording in O84T ringbuffers at ILKN. |
596 | 2021-09-27 | 18:10 | O84T | Site Geology | |
597 | 2021-09-27 | 18:10 | O84T | VSAT Start | Co-pol 27 X-pol 0 |
598 | 2021-09-27 | 22:13 | SNPN | Bird-protection | No bird damage observed but additional tie-wraps were installed on feed arms. |
599 | 2021-09-27 | 22:13 | SNPN | Re-Point | X-pol 0 Co-pol 20 dB. |
600 | 2021-09-27 | 22:13 | SNPN | Service Visit | Visit initiated due to poor VSAT performance. |
601 | 2021-09-27 | 22:13 | SNPN | VSAT Check | Bad co-pol due probably to shifting of antenna. X-pol 0 Co-pol 7dB. |
602 | 2021-09-27 | 22:36 | SNPN | Cold-Boot | Restarted from cold boot. Eirp increase adjusted to 5 dB at Hub. |
603 | 2021-09-27 | 23:54 | ILKN | DC Filter Check | Started with DC filter off. |
604 | 2021-09-27 | 23:54 | ILKN | Gain Check | Started with software gain of 1.0 at 40 sps. |
605 | 2021-09-27 | 23:54 | ILKN | New Station Start | Start recording in ILKN ringbuffers. |
606 | 2021-09-30 | 13:50 | COWN | Re-Point | Due to lack of communication to GE, x-pol/co-pol not available. |
607 | 2021-09-30 | 13:50 | COWN | Service Visit | Station was down due to shifting of antenna. No wildlife damage observed but additional bird protection (shark teeth) installed on feed horn box. |
608 | 2021-09-30 | 18:00 | ACKN | DC Filter Check | DC filter unchanged at off. |
No | Date | Time UT | Location | Activity | Comments |
609 | 2021-09-30 | 18:00 | ACKN | Gain Check | Software gain unchanged at 1.0 at 40 sps. |
610 | 2021-09-30 | 18:00 | ACKN | Hardware Swap | Cygnus 220 removed, 265 installed. 2 W SSPB removed, 1 W installed (SN A0208A). |
611 | 2021-09-30 | 18:00 | ACKN | Re-Point | Much improved. X-pol 0 Co-pol 28 dB. Additional bird protection (shark teeth) installed. |
612 | 2021-09-30 | 18:00 | ACKN | SSPB Check | All looked ok at station and at Hub (on analyzer), but no data received, even after a Cygnus swap. It was concluded that the new 2 W sspb was not working. |
613 | 2021-09-30 | 18:00 | ACKN | Service Visit | Visit initiated due to poor VSAT performance. |
614 | 2021-09-30 | 18:00 | ACKN | VSAT Check | Before re-pointing: X-pol 0 Co-pol 20 dB. |
615 | 2021-10-02 | 18:58 | P85T | Hardware List | Cyg 175, Trd 073, Trillium 0075, SSPB A0171A. |
616 | 2021-10-02 | 18:58 | P85T | New Station Start | Start recording in ALFO ringbuffers. Cold Boot ok. |
617 | 2021-10-02 | 18:58 | P85T | New Station Start | Start recording in P85T ringbuffers at ALFO. |
618 | 2021-10-02 | 18:58 | P85T | Site Geology | Bedrock |
619 | 2021-10-02 | 18:58 | P85T | VSAT Start | Co-pol 31 X-pol 1. |
620 | 2021-10-02 | 19:25 | EKTN | Hardware Swap | Cygnus 222 removed, 392 installed. All else the same. |
621 | 2021-10-02 | 19:25 | EKTN | Service Visit | Visit initiated to swap Cygnus. |
622 | 2021-10-02 | 20:01 | EKTN | Gain Check | Software gain unchanged at 1.0 at 40 sps. |
623 | 2021-10-02 | 20:18 | ALFO | DC Filter Check | Started with DC filter on at 0.001 Hz. |
624 | 2021-10-02 | 20:18 | ALFO | Gain Check | Started with software gain of 0.5 at 100 sps, input range 16 Vp-p. |
No | Date | Time UT | Location | Activity | Comments |
625 | 2021-10-19 | 15:30 | PKRO | Service Visit | Station was down due to bad Cygnus. |
626 | 2021-10-19 | 15:56 | PKRO | Hardware Swap | Cyg 222 installed, 172 removed. |
627 | 2021-10-22 | 16:15 | HGVO | Seismo Swap | ESP-T3B16 Removed. 40T-T4770 insatalled. |
628 | 2021-10-29 | 16:30 | OTRO | Service Visit | Just to check VSAT. Solar panel racks and antenna loose, tightened. Added bird deterent on feed horn. Local person checks once in a while. |
629 | 2021-10-29 | 16:32 | OTRO | VSAT Check | Before 30/0, after 30/0. |
630 | 2021-10-29 | 20:30 | MALO | Service Visit | Station not working ok due to bad Cygnus? Visit indicated old Cygnus not working in Test mode. No physical damage observed at structures. |
631 | 2021-10-29 | 20:31 | MALO | Hardware Swap | Cyg 422 removed, 220 installed. |
632 | 2021-10-29 | 20:32 | MALO | VSAT Check | 28/1. No change from installation time. |
633 | 2021-10-29 | 20:34 | MALO | Restart | All OK with new Cygnus. |
634 | 2021-10-29 | 20:40 | MALO | Gain Check | Unchanged from 1. |
635 | 2021-10-29 | 20:41 | MALO | DC Filter Check | Unchaged from On at 1 mHz. |
636 | 2021-10-30 | 16:08 | KILO | Service Visit | Station working ok but Trident did not connect after reboot. |
637 | 2021-10-30 | 16:09 | KILO | Wildlife | Live and dead mouce in seismo vault. |
638 | 2021-10-30 | 16:10 | KILO | Comments | First time Trident failed to connect due to bad 1.7201 firmware. |
639 | 2021-10-30 | 16:18 | KILO | Hardware Swap | Trident 516 installed, 369 removed. |
640 | 2021-10-30 | 17:35 | KILO | SR Change | Incorrectly set to 40 sps. |
No | Date | Time UT | Location | Activity | Comments |
641 | 2021-10-30 | 17:36 | KILO | VSAT Check | Before 30/4 now 30/1 dB. |
642 | 2021-10-30 | 17:37 | KILO | Gain Check | Remained at 1, as before. |
643 | 2021-10-30 | 17:38 | KILO | DC Filter Check | Remained on at 1 mHz as before. |
644 | 2021-10-30 | 17:40 | KILO | Repair | Seismo vault sealed. |
645 | 2021-10-31 | 13:04 | SUNO | Firmware Upgrade | Trident 1.7210 in B. B is active. |
646 | 2021-10-31 | 14:41 | PCTO | Firmware Upgrade | Installed Trident 1.7210 in A. |
647 | 2021-10-31 | 15:06 | SUNO | Service Visit | Station was working, some hardware needed tigthening. |
648 | 2021-10-31 | 15:08 | SUNO | VSAT Check | Before 32/5 after 32/1 dB. |
649 | 2021-10-31 | 15:09 | SUNO | VSAT Start | Cold boot and restarted VSAT. |
650 | 2021-10-31 | 18:07 | PTCO | Closed | Moving to more secure location (BUKO). |
651 | 2021-10-31 | 18:08 | PTCO | Comments | 2 solar panel stolen, one shattered. |
652 | 2021-10-31 | 18:09 | PTCO | Comments | Old large seismo valut and fence left at the site. |
653 | 2021-11-01 | 13:55 | KILO | SR Change | Restored from 40 to 100 sps. |
654 | 2021-11-01 | 13:56 | KILO | Firmware Upgrade | Trident 1.7210 in B. |
655 | 2021-11-01 | 14:23 | KILO | Firmware Upgrade | Trident 1.7210 in A and active. |
656 | 2021-11-02 | 18:00 | BUKO | New Station Start | Started new statation (moved hardware from PTCO). |
No | Date | Time UT | Location | Activity | Comments |
657 | 2021-11-02 | 18:01 | BUKO | VSAT Start | 28/1 dB. |
658 | 2021-11-02 | 18:02 | BUKO | Hardware List | CYG179 TRD071 ESP-T3B18 |
659 | 2021-11-02 | 18:03 | BUKO | Site Geology | Bedrock (outcrop). |
660 | 2021-11-02 | 18:04 | BUKO | Power Upgrade | Needed on next visit. Only 5 solar panels installed. |
661 | 2021-11-03 | 13:50 | BUKO | Gain Check | Started with software gain of 0.5. |
662 | 2021-11-03 | 13:51 | BUKO | DC Filter Check | Started with DC filter on at 1 mHz. |
663 | 2021-11-03 | 14:16 | DSMN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
664 | 2021-11-03 | 14:53 | ENGB | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
665 | 2021-11-03 | 15:12 | EKTN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
666 | 2021-11-03 | 16:04 | MALO | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
667 | 2021-11-03 | 16:05 | MUMO | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
668 | 2021-11-03 | 17:33 | SUNO | Firmware Upgrade | Trident 1.7210 in A. A is active. |
669 | 2021-11-03 | 17:34 | SUNO | Comments | Libra firmware is out of date (5.50) |
670 | 2021-11-03 | 17:35 | OTRO | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
671 | 2021-11-03 | 18:09 | VIMO | Comments | Libra firmware is out of date (5.50) |
672 | 2021-11-03 | 18:35 | VIMO | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
No | Date | Time UT | Location | Activity | Comments |
673 | 2021-11-03 | 19:31 | ACKN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
674 | 2021-11-03 | 20:23 | GBLN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
675 | 2021-11-03 | 21:02 | GLWN | Firmware Upgrade | Trident EPLD 1.23 installed in B. |
676 | 2021-11-03 | 21:03 | GLWN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
677 | 2021-11-03 | 21:21 | ILKN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
678 | 2021-11-04 | 14:48 | LDGN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
679 | 2021-11-04 | 15:08 | MCKN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
680 | 2021-11-04 | 15:24 | KNDN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
681 | 2021-11-04 | 15:26 | LUPN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
682 | 2021-11-04 | 16:24 | NODN | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
683 | 2021-11-04 | 16:42 | MGCB | Firmware Upgrade | Trident 1.7210 in A and B. A is active. |
684 | 2021-11-04 | 16:44 | YNEN | Firmware Upgrade | Trident EPLD 1.23 installed in B. |
685 | 2021-11-04 | 16:54 | YNEN | Firmware Upgrade | Trident 1.7210 in A and B. B is active. |
686 | 2021-11-05 | 18:45 | GOWB | Service Visit | Station was going down repeatedly. Visit to find the problem. |
687 | 2021-11-05 | 18:46 | GOWB | VSAT Check | Before repoint and after repoint attemp 27/0. |
688 | 2021-11-05 | 18:47 | GOWB | Comments | No problem found at the site. |
No | Date | Time UT | Location | Activity | Comments |
689 | 2021-11-05 | 19:00 | KSVO | Service Visit | Station went down after days of acting up. |
690 | 2021-11-05 | 19:01 | KSVO | VSAT Check | Before repoint attemp, 32/2, after 32/2 dB. No chagne. |
691 | 2021-11-05 | 19:02 | KSVO | Battery Check | All batteries and charge controller ok. |
692 | 2021-11-05 | 19:03 | KSVO | Cables | LNB cable found loose. All cables tightened. |
693 | 2021-11-05 | 19:05 | KSVO | Comments | Was LNB cable the only reason for problems?? |
694 | 2021-11-10 | 15:57 | HGVO | Temp Halt | Cyg, Tri, seismo returned to UWO for repair. |
695 | 2021-11-10 | 21:05 | MALO | Comments | Rebooted Time Server to fix gps sat problem. |
696 | 2021-11-10 | 21:20 | P86T | New Station Start | Electronics from HGVO under test at UWO. |
697 | 2021-11-15 | 20:37 | ELGO | Comments | Rebooted Time Server to fix GPS SAT not showing problem. |
698 | 2021-11-15 | 20:40 | ELGO | Comments | Now showing Instrument, Timing and GPS Sat SOH correctly. Needed a reboot. |
699 | 2021-11-17 | 19:08 | BUKO | Service Visit | Added one solar panel, replaced some battery terminals |
700 | 2021-11-17 | 19:12 | BUKO | Battery Check | Starting to charge batteries with generator. |
701 | 2021-11-17 | 19:13 | BUKO | Restart | Site restarting. |
702 | 2021-11-18 | 13:00 | BUKO | Battery Check | Generator stopped. Lasted about 18 hours?? |
703 | 2021-11-18 | 16:00 | HGVO | Service Visit | New hardware, trying to find source of problem. |
704 | 2021-11-18 | 16:10 | HGVO | Cables | Bus cable checks ok. Not replaced. Grounding ok. |
No | Date | Time UT | Location | Activity | Comments |
705 | 2021-11-18 | 16:20 | HGVO | Hardware List | CYG422 TIM528 TRD237 T3B16(ESP) |
706 | 2021-11-18 | 16:21 | HGVO | VSAT Check | 35/1.5. |
707 | 2021-11-18 | 16:24 | HGVO | Hub Swap | Served by Ottawa Hub. |
708 | 2021-11-18 | 16:25 | HGVO | Restart | Restarted with new hardware. |
709 | 2021-11-18 | 16:26 | HGVO | SR Change | Incorrect at 40 sps. |
710 | 2021-11-18 | 16:35 | HGVO | Gain Check | Unchanged at 0.5. |
711 | 2021-11-18 | 16:36 | HGVO | DC Filter Check | Unchaged at on and 1 mHz. |
712 | 2021-11-18 | 16:37 | HGVO | SR Change | Corrected from 40 to 100 sps. |
713 | 2021-11-18 | 17:01 | HGVO | Comments | VSAT log in ok at ip 10.6.255.33 |
714 | 2021-11-18 | 17:02 | HGVO | Battery Check | DC volts ok at 13.5 |
715 | 2021-11-19 | 22:24 | BUKO | Battery Check | Generator started again. Battery volts 12.6 VDC |
716 | 2021-11-21 | 00:47 | CLVB | Battery Check | Charged for many hours using generator. |
717 | 2021-11-21 | 00:48 | CLVB | Restart | Station restarted after battery charged. |
718 | 2021-11-21 | 00:49 | CLVB | Comments | LVD set to 10.5 VDC. |
719 | 2021-11-21 | 00:50 | CLVB | Battery Check | Batteries charged for 5 hours. |
720 | 2021-11-21 | 13:59 | OTTAWA | Hub Login | NTP server is now 132.156.100.63. |
No | Date | Time UT | Location | Activity | Comments |
721 | 2021-11-21 | 14:01 | HGVO | VSAT Check | VSAT power level increased by 2 dB at Ottawa Hub. |
722 | 2021-11-21 | 14:10 | OTTAWA | Hub Login | Firewall NTP server set to 132.156.100.63. |
723 | 2021-11-22 | 20:45 | TOBO | Temp Halt | Station stopped. No AC power, Battery volts at 8.7 VDC. |
724 | 2021-11-22 | 21:32 | COWN | Battery Check | Low battery volts, 11.92 VDC. |
725 | 2021-11-22 | 21:33 | CLVB | Battery Check | Low battery volts, 11.92 VDC. 11.67 VDC. |
726 | 2021-11-23 | 18:05 | CLVB | Temp Halt | Stopped at 11.52 VDC due to low battery power. |
727 | 2021-11-23 | 21:55 | COQB | Comments | Station off. VSAT problem?? |
728 | 2021-11-24 | 08:11 | OTRO | Temp Halt | Stopped due to heavy snow & ice on antenna? |
729 | 2021-11-24 | 19:20 | OTRO | VSAT Check | Increased VSAT power at hub by 6 dB and changed modulation to 64 KB to get the site back on. |
730 | 2021-11-24 | 19:42 | OTRO | VSAT Check | Remote Tx power too low. |
731 | 2021-11-25 | 19:01 | OTRO | VSAT Check | Restored VSAT power at Hub to original level of 2dB and modulation to 112KB. |
732 | 2021-11-25 | 19:03 | OTRO | Comments | VSAT is back to normal, all appear OK. |
733 | 2021-11-26 | 17:44 | OTTAWA | Carina Hubs | Carina 62 and 63 were powered down briefly to switch to temp gps antenna. |
734 | 2021-11-26 | 19:25 | OTTAWA | Carina Hubs | Carina 62 and 63 were powered down briefly to switch to new gps antenna. |
735 | 2021-11-26 | 20:06 | OTTAWA | Carina Hubs | Original GPS antenna cables were found to have been damaged during Hub installation. All cables were replaced. Improvements observed in GPS S/N on all Carinas. |
736 | 2021-12-01 | 12:39 | HGVO | VSAT Check | Hub power level increased from 5 to 7 dB, throughput from 8000 to 11000. |
No | Date | Time UT | Location | Activity | Comments |
737 | 2021-12-01 | 12:40 | HGVO | Comments | Hub tweaked to improve gap situation. |
738 | 2021-12-08 | 20:00 | SUNO | Service Visit | Added 6 more batteries to power system. |
739 | 2021-12-08 | 20:30 | SUNO | Battery Check | Started charging batteries with generator. |
740 | 2021-12-08 | 20:54 | SUNO | Battery Check | Battery volt is now at 12.58VDC. |
741 | 2021-12-09 | 15:20 | SUNO | Battery Check | Stopped generator. Batteries are at 12.94 VDC. |
742 | 2021-12-09 | 15:28 | SUNO | Seismo Check | NS component has large DC offset, looks odd. Bad Seismo?? |
743 | 2021-12-09 | 19:45 | CLVB | Service Visit | 8 New fully charged Batteries installed, old batteries removed. |
744 | 2021-12-09 | 20:03 | CLVB | Battery Check | All batteries are ok now. VDC is 12.92. |
745 | 2021-12-09 | 20:03 | CLVB | Restart | Site restarting. |
746 | 2021-12-09 | 20:05 | CLVB | Battery Check | Solar panel is charging ok too. 20 VDC. |
747 | 2021-12-19 | 15:00 | HGVO | Service Visit | Station down about 1/2 hour to swap bus cable. |
748 | 2021-12-19 | 16:00 | HGVO | Cables | Bus cable replaced to fix gap problems. |
749 | 2021-12-19 | 18:00 | TYNO | Service Visit | Station down about 1/2 hour to swap batteries. |
750 | 2021-12-19 | 18:35 | TYNO | Battery Check | AC-type batteries and LVD installed. |
751 | 2021-12-19 | 18:44 | HBDB | Service Visit | Station shut down to swap electronics. |
752 | 2021-12-19 | 18:45 | HBDB | Hardware List | Old hardware to be removed: CYG394 TRD425 |
No | Date | Time UT | Location | Activity | Comments |
753 | 2021-12-19 | 19:20 | HBDB | Hardware Swap | CYG394 TRD425 removed, CYG161 TRD064 installed |
754 | 2021-12-19 | 20:01 | HBDB | Restart | Back on air. |
755 | 2021-12-19 | 20:02 | HBDB | Hub Swap | Moved to Ottawa Hub. |
756 | 2021-12-19 | 20:10 | HBDB | DC Filter Check | Unchanged, On at 1 mHz. |
757 | 2021-12-19 | 20:12 | HBDB | Gain Check | Unchanged at 0.5. |
758 | 2021-12-19 | 20:13 | HBDB | Comments | IP is at 161 on Hub 5. |
759 | 2021-12-19 | 21:00 | ACTO | Service Visit | Station down about 1/2 hour to swap Trident. |
760 | 2021-12-19 | 21:30 | ACTO | Hardware Swap | Trident 77 removed, 295 installed. |
761 | 2021-12-19 | 21:31 | ACTO | DC Filter Check | Unchanged from on at 1 mHz. |
762 | 2021-12-19 | 21:32 | ACTO | Gain Check | Unchanged from 0.5. |
763 | 2022-01-16 | 18:00 | CLVB | Service Visit | To swap batteries. |
764 | 2022-01-16 | 18:10 | CLVB | Battery Check | 8 charged batteries installed. |
765 | 2022-01-16 | 18:20 | CLVB | VSAT Check | Co/X-pol 33/4 dB. |
766 | 2022-01-16 | 19:30 | CLVB | Cold-Boot | All started after cold boot. |
767 | 2022-01-16 | 19:36 | CLVB | Restart | Station restarted. |
768 | 2022-02-02 | 17:41 | O91T | Comments | Temp Station O91T, at least, on Jan. 31 and Feb. 1, 04 shows Trident SOH temperature is about 10 deg warmer than the SSPB temperature. Both sensors are in the same test vault. |
No | Date | Time UT | Location | Activity | Comments |
769 | 2022-02-12 | 19:20 | TWKB | Seismo Check | HHN Component is bust. Mass-center did not help. Calibration shows odd signal on HHN but other two chans. ok. Seismo stuck? |
770 | 2022-02-20 | 11:00 | PGCB | VSAT Start | Restarted test station, X/Co: 2/31dB. CMG40T 4688, cyg 172, trd 369 |
771 | 2022-02-21 | 00:01 | SILO | VSAT Check | Modulation increased from 64 to 112KB to improve bandwidth use. |
772 | 2022-02-21 | 00:03 | VIMO | VSAT Start | Modulation reduced from 112 to 64 KB to improve chance of connecting to the station. |
773 | 2022-02-21 | 00:04 | VIMO | VSAT Check | EIRP level increased from 6 to 8 dB to get the station up again. |
774 | 2022-02-21 | 22:00 | VIMO | VSAT Check | Station back up after VSAT adjustments. |
775 | 2022-02-21 | 22:01 | SUNO | VSAT Check | EIRP level reduced from 10 to 6 dB balance levels. |
776 | 2022-02-23 | 19:03 | MGTN | Comments | Station was not sending seismic signal. VSAT login showed no Trident. Rebooted. |
777 | 2022-02-23 | 19:05 | MGTN | Comments | Reboot did not solve problem. Trident not showing. No seismic signal. |
778 | 2022-02-27 | 15:05 | B7-VAULT | New Station Code | O88T, Internet station. Started. |
779 | 2022-02-27 | 15:07 | B7-VAULT | Hardware List | CYG284 TIM228 TRD231, 3ESP |
780 | 2022-02-27 | 15:09 | B7-VAULT | Gain Check | Started at 100 SPS, 40 VPP, Gain of 0.5 |
781 | 2022-02-27 | 15:10 | B7-VAULT | DC Filter Check | Started with NO DC filter. |
782 | 2022-02-27 | 15:11 | O88T | New Station Start | O88T, Ottawa Vault is now under Test. |
783 | 2022-03-02 | 17:06 | ENGB | Service Visit | Stopped station for maintenance visit. |
784 | 2022-03-02 | 17:13 | ENGB | VSAT Check | X/Co: 0/27 dB. |
No | Date | Time UT | Location | Activity | Comments |
785 | 2022-03-02 | 17:17 | ENGB | Restart | Station restarted, all OK. |
786 | 2022-03-02 | 19:44 | HLSB | Service Visit | Stopped station for maintenance visit. |
787 | 2022-03-02 | 19:55 | HLSB | VSAT Check | X/Co: 5/34 dB. |
788 | 2022-03-02 | 19:58 | HLSB | Cold-Boot | Restarted OK. |
789 | 2022-03-02 | 20:00 | HLSB | DC Filter Check | Unchanged at On. |
790 | 2022-03-02 | 20:00 | HLSB | Gain Check | Unchanged at 0.5. |
791 | 2022-03-02 | 20:58 | HRMB | Service Visit | Stopped station for maintenance visit. |
792 | 2022-03-02 | 21:00 | HRMB | VSAT Check | X/Co. 3/33 dB. |
793 | 2022-03-02 | 21:08 | HRMB | Cold-Boot | Restarted OK. |
794 | 2022-03-02 | 21:10 | HRMB | DC Filter Check | Unchaged from On at 1 mHz. |
795 | 2022-03-02 | 21:10 | HRMB | Gain Check | Unchanged at 0.5. |
796 | 2022-03-02 | 22:13 | HBDB | Service Visit | Stopped station for maintenance visit. |
797 | 2022-03-02 | 22:15 | HBDB | VSAT Check | X/Co. 2/32 dB. |
798 | 2022-03-02 | 22:18 | HBDB | Cold-Boot | Restarted OK. |
799 | 2022-03-15 | 16:48 | ALGO | DC Filter Change | Changed from ON @0.0010 Hz to OFF. Requested by Savka D. for calibration test. Will be restored to on later. |
800 | 2022-03-15 | 20:30 | ALGO | DC Filter Change | Changed from OFF to ON at 0.001 Hz. Savka's test finished. |
No | Date | Time UT | Location | Activity | Comments |
801 | 2022-03-15 | 21:40 | TWKB | Hardware Swap | TRD071, CMG-3ESP T3D91 removed. TRD425, CMG-3ESP T3G52 installed. |
802 | 2022-03-15 | 21:41 | TWKB | DC Filter Check | DC filter on at 0.0010 Hz |
803 | 2022-03-15 | 21:42 | TWKB | Gain Check | Software gain = 0.5 |
804 | 2022-03-26 | 17:00 | CLWO | Service Visit | To replace seismo. Station found in good and secure condition. Snow in the field. Walked to the site. |
805 | 2022-03-26 | 17:10 | CLWO | Hardware Swap | 40T removed, ESP (T3G91) installed. |
806 | 2022-03-26 | 17:20 | CLWO | VSAT Check | Antenna re-pointed. X/Co: 1/29 dB. |
807 | 2022-03-26 | 17:25 | CLWO | DC Filter Check | No change. |
808 | 2022-03-26 | 17:30 | CLWO | Gain Check | No change. |
809 | 2022-03-26 | 19:30 | CLWO | Comments | End of Service. Sericed by KA. Database entry IA. |
810 | 2022-03-26 | 22:45 | TOBO | Service Visit | To replace seismo. Station found in good and secure condition. |
811 | 2022-03-26 | 22:55 | TOBO | Hardware Swap | 40T removed, ESP (T3G89) installed. |
812 | 2022-03-26 | 23:10 | TOBO | Comments | Laptop did not work so no re-pointing possible. No software changes. Serviced by KA, database entry IA. |
813 | 2022-04-01 | 17:15 | KLBO | Seismo Swap | CMG-40T removed, ESP T3H26 installed. |
814 | 2022-04-01 | 17:20 | KLBO | VSAT Check | X/Co pol 3/34 dB. |
815 | 2022-04-01 | 17:30 | KLBO | Comments | No change to gain of dc filter, station was good and secure. |
816 | 2022-04-19 | 22:00 | BRCO | Firmware Upgrade | TimeServer V2.00.00 upgrade to fix GPS SOH problem. Not sure if upgrade fixed the problem or simply the reboot did it. Served by BD, data entry IA. |
No | Date | Time UT | Location | Activity | Comments |
817 | 2022-05-04 | 18:05 | P99T | New Station Start | Started recording at TORO in temp P99T ringbuffers. |
818 | 2022-05-04 | 18:10 | P99T | VSAT Check | Co/X Pol 35/2 dB. |
819 | 2022-05-04 | 19:00 | TORO | New Station Code | Started recording at TORO ringbuffres. |
820 | 2022-05-04 | 19:10 | TORO | Hardware List | Cyg173 TRD059 CMG-40T T4565 SSPB A0157A |
821 | 2022-05-04 | 19:11 | TORO | Gain Check | Software gain 0.5 at 100 SPS. |
822 | 2022-05-04 | 19:12 | TORO | DC Filter Check | Filter on at 1 mHZ (check 1 mHz again). |
823 | 2022-05-04 | 19:15 | TORO | Site Geology | Man-made fill. |
824 | 2022-05-07 | 21:00 | MEDO | New Station Code | Started recording at MEDO ringbuffers. |
825 | 2022-05-07 | 21:01 | MEDO | Hardware List | Cyg 394, TRD095, TS375, ESP T3B91, SSPB a0117a. |
826 | 2022-05-07 | 21:02 | MEDO | VSAT Check | Co/X pol 36/6 dB. |
827 | 2022-05-07 | 21:03 | MEDO | Gain Check | Started at Gain 0.5, sps 100. |
828 | 2022-05-07 | 21:05 | MEDO | DC Filter Check | Started with DC filter on at 1 mHz. |
829 | 2022-05-07 | 21:07 | MEDO | Comments | Fence installed to stop cows. |
830 | 2022-05-10 | 15:10 | TORO | DC Filter Change | DC filter incorrectly set to 0.1 Hz. |
831 | 2022-05-10 | 15:10 | TORO | Hardware Swap | Trd. 077 installed, 059 removed. |
832 | 2022-05-10 | 18:33 | TORO | DC Filter Change | Corrected cutoff freq. to 0.001 HZ. |
No | Date | Time UT | Location | Activity | Comments |
833 | 2022-05-10 | 20:00 | MEDO | Firmware Upgrade | Libra firmware to V6.65.01. |
834 | 2022-05-10 | 20:05 | MEDO | Firmware Upgrade | TS upgraded to V2.00.00. |
835 | 2022-05-10 | 20:10 | MEDO | SSPB Check | SSPB gain table set ftat to 45.0 dB. |
836 | 2022-05-10 | 20:15 | MEDO | Comments | TX IFL loss set to 2.0 dB for 6 m cable. |
837 | 2022-05-10 | 20:20 | MEDO | Comments | Ringbuffer sizes changed (BDunn). |
838 | 2022-05-13 | 20:30 | MGTN | Temp Halt | Station was removed from carina 31, TDMA. No data coming from Trident. |
839 | 2022-05-13 | 20:30 | PLVO | DC Filter Check | DC filter is on at 0.001 HZ, startup. |
840 | 2022-05-13 | 20:30 | PLVO | Gain Check | Gain is at 0.5 for startup. |
841 | 2022-05-13 | 20:30 | PLVO | Hardware List | CYG541, TS694, TRD733, SSPB A0172A, Seis. ESP T3H25 (from UWO). |
842 | 2022-05-13 | 20:30 | PLVO | VSAT Check | Co/X 32/.4 dB. |
843 | 2022-05-13 | 22:30 | PLVO | New Station Code | Started recording at PLVO ringbuffers. |
844 | 2021-05-26 | 14:30 | PEMO | Service Visit | To upgrade seismometer. |
845 | 2021-05-26 | 14:31 | PEMO | Seismo Swap | CMG-40T s/n T4185 removed. ESP s/n T3J31 installed. |
846 | 2021-05-26 | 14:32 | PEMO | Comments | Evidence of mice in seismometer vault. |
847 | 2021-05-26 | 14:33 | PEMO | Seismo Check | Mass centered and calibrated ok. |
848 | 2021-05-26 | 14:34 | PEMO | Gain Check | Unchaged at 0.5. |
No | Date | Time UT | Location | Activity | Comments |
849 | 2021-05-26 | 14:35 | PEMO | DC Filter Check | Unchaged from On at 1 mHz. |
850 | 2021-05-26 | 14:36 | PEMO | Firmware Upgrade | Trident upgraded to 1.7201, epld 1.23. TS was running current v 1.23. Unable to upgrade Libra, kept old version 5.42 in A. Newer 5.6201 loaded in B but fails to test. |
851 | 2021-05-27 | 14:47 | ALFO | Service Visit | To replace seismometer. Site looked ok. |
852 | 2021-05-27 | 14:50 | ALFO | Seismo Swap | Trillium 0075 removed. ESP T3J04 installed. |
853 | 2021-05-27 | 14:51 | ALFO | Gain Check | Unchaged at 0.5. |
854 | 2021-05-27 | 14:52 | ALFO | DC Filter Check | Unchaged at on and 1 mHz. |
855 | 2021-05-27 | 14:53 | ALFO | Comments | Trident Input range changed from 16 (for Trillium) to 40 Vpp(for ESP). |
856 | 2021-05-27 | 16:10 | ALFO | Cold-Boot | Standard procedure. |
857 | 2021-05-27 | 16:15 | ALFO | Comments | For next visit: Security cable for battery box and Proper upper pedestal to fit over leveling plate. |
858 | 2021-06-03 | 17:00 | KLNB | Comments | VSAT errors must be due to antenna pointed while Eirp power was at -7dB. Power is restored to 0 dB at Hub and a re-point is needed. |
859 | 2021-06-07 | 13:32 | TORO | Cables | Seismo cable replaced to fix EW component. Some noise heard before cable swap, not after. |
860 | 2021-06-07 | 13:34 | TORO | Comments | No system shut down. Cable swap took a couple of minutes. Waveform show good traces for now. |
861 | 2021-06-09 | 18:30 | HLSB | Service Visit | To investigate VSAT problems. No damage to site but tall grass in view of antenna. Will take pics. |
862 | 2021-06-09 | 18:31 | HLSB | VSAT Check | No signal from station. Increased VSAT power at hub by 2 dB and did cold boot on station. No signal yet. |
863 | 2021-06-09 | 18:37 | HLSB | VSAT Check | Increased VSAT by another 5 dB (in total, changed from -7dB to 0 dB). No change. Asked operator to check all RF cables and VSAT. |
864 | 2021-06-09 | 18:40 | HLSB | VSAT Check | RF cable appears to have been damaged. Moving cable a little helped and station came back up. Shutting down to fix cable. |
No | Date | Time UT | Location | Activity | Comments |
865 | 2021-06-09 | 18:50 | HLSB | VSAT Check | SSPB cable has small scratch on the tape at the sspb end. Suspecting damage to feed horn. Plan to re-point the antenna. |
866 | 2021-06-09 | 19:26 | HLSB | VSAT Check | Replace feed horn. Contacted GE. They report signal is trash. Suspect bad Cygnus. |
867 | 2021-06-09 | 19:50 | HLSB | Temp Halt | Shut down to replace Cygnus. Need to bring a Cygnus from test site at PGC tomorrow. |
868 | 2021-06-09 | 21:30 | O96T | New Station Start | Started first of FedNor2 stations at Atikokan Iron Mine, ATKO, recoding in O96T ringbuffers for the next 23 minutes. |
869 | 2021-06-09 | 21:34 | ATKO | VSAT Check | Co/X 29/0 dB. Ajusted IFL to 2 dB for long Cable. Correct SSPB profile entred. |
870 | 2021-06-09 | 21:36 | ATKO | Hardware List | CYG539, TS696, TRD059, Seismo. ESP T377, SSPB A0395A |
871 | 2021-06-09 | 21:37 | ATKO | Gain Check | 100 SPS, Gain 0.5, startup gain. |
872 | 2021-06-09 | 21:38 | ATKO | DC Filter Check | DC filter on at 0.001Hz. Startup. |
873 | 2021-06-09 | 21:39 | ATKO | Site Geology | Solid bedrock on hill top about 30 ft high about 40 yrds from access road. |
874 | 2021-06-09 | 21:40 | ATKO | Comments | Old (may be inactive) hydro wire nearby. |
875 | 2021-06-09 | 21:53 | ATKO | New Station Code | Started recording on Naqs2 in ATKO ringbuffers. |
876 | 2021-06-09 | 21:58 | ATKO | New Station Code | Started recording in Naqs1 in ATKO ringbuffers. |
877 | 2021-06-10 | 17:25 | LONDON HUB | Comments | Moved NAQS1 and Polaris firewall, 2 network switches, modem, monitor and KVM switch to 24 hour UPS. |
878 | 2021-06-10 | 17:26 | LONDON | Carina Hubs | 3 out of 4 were hung due to power switch. Cold booted to fix problem. |
879 | 2021-06-10 | 17:27 | LONDON | Carina Hubs | Carina Hubs and fans will be moved to UPS tomorrow. |
880 | 2021-06-10 | 19:00 | HLSB | VSAT Check | Co/X pol 25/0 dB. |
No | Date | Time UT | Location | Activity | Comments |
881 | 2021-06-10 | 19:40 | HLSB | Comments | New feedhorn installed with new rf cables. |
882 | 2021-06-10 | 19:40 | HLSB | Hardware Swap | Cygnus 217 removed, 172 installed to fix VSAT problem. |
883 | 2021-06-10 | 19:41 | HLSB | Comments | Station operation on London Carina 3 now. |
884 | 2021-06-10 | 19:45 | HLSB | Hardware Swap | New SSPB H0018A installed. Temperature profile needs to be updated. |
885 | 2021-06-10 | 19:46 | HLSB | Cold-Boot | Done. |
886 | 2021-06-10 | 19:47 | HLSB | Comments | Cygnus was from test site at PGC, was configured for GPS on serial port. Serial port was removed. |
887 | 2021-06-10 | 19:50 | HLSB | DC Filter Check | On at 0.001 Hz. Unchanged. |
888 | 2021-06-10 | 19:51 | HLSB | Gain Check | Unchanged at 0.5. |
889 | 2021-06-11 | 17:12 | KLNB | Service Visit | To improve VSAT. |
890 | 2021-06-11 | 17:14 | KLNB | VSAT Check | Re-point. Co/X now at 26/0 dB. Was 22/0 dB before. |
891 | 2021-06-11 | 17:15 | KLNB | Comments | Earlier pointing was done when raining. No good. Today is cloudy but ok. |
892 | 2021-06-11 | 17:16 | KLNB | Comments | In Trident Operationl screen, RX overrun increased from 22 to 30 in about 10 mins. No good. Bus problem? |
893 | 2021-06-11 | 17:43 | KLNB | Cold-Boot | All is ok. Trident screen was checked. Overruns look ok. |
894 | 2021-06-11 | 17:44 | KLNB | DC Filter Check | DC filter is on at 0.001 HZ. Unchanged. |
895 | 2021-06-11 | 17:45 | KLNB | Comments | Construction work in underway near the site. Plan to finish in a few months. |
896 | 2021-06-11 | 17:45 | KLNB | Gain Check | Is 0.5, unchanged. |
No | Date | Time UT | Location | Activity | Comments |
897 | 2021-06-11 | 17:46 | KLNB | Comments | When construction is done, school buses will park near the site. Expect noise during their work hours. |
898 | 2021-06-11 | 18:12 | O93T | Hardware List | Sspb a0340a, cyg 523, ts 653, trd 693. seis t379 esp. |
899 | 2021-06-11 | 18:12 | O93T | New Station Start | Started new site at Exp. Lake. |
900 | 2021-06-11 | 18:13 | O93T | VSAT Check | Co/x 25/0 dB. |
901 | 2021-06-11 | 18:14 | O93T | DC Filter Check | On is at 0.001 starting. |
902 | 2021-06-11 | 18:14 | O93T | Gain Check | Sps 100, gain is 0.5 starting. |
903 | 2021-06-11 | 18:14 | O93T | Site Geology | Cleared forested area, exposed bedrock. |
904 | 2021-06-11 | 18:15 | O93T | Comments | Trident input range is at 40 pp. |
905 | 2021-06-11 | 18:16 | O93T | Comments | Finishing off the site in the next hour. |
906 | 2021-06-11 | 18:17 | O93T | Comments | Cold boot was done before latest checks. |
907 | 2021-06-11 | 18:33 | EPLO | New Station Start | Started on Naqs2, Ottawa Hub, in EPLO buffers. |
908 | 2021-06-11 | 18:34 | EPLO | New Station Start | Started on Naqs1, Ottawa Hub, in EPLO buffers. |
909 | 2021-06-11 | 19:42 | LDGN | VSAT Check | Problem with VSAT: Dish out or Cygnus problem. TX is down by 5 dB but shape is ok. No data coming. |
910 | 2021-06-11 | 19:44 | LDGN | Comments | Site probably need a visit. |
911 | 2021-06-13 | 16:51 | O94T | New Station Start | Installed and starting at Red Lake but not connected to Hub yet. VSAT problems. |
912 | 2021-06-13 | 17:21 | O94T | New Station Start | Cygnus 536 is installed, 535(not working) removed. VSAT is on now. |
No | Date | Time UT | Location | Activity | Comments |
913 | 2021-06-13 | 17:35 | O94T | Comments | Mass centered ok. |
914 | 2021-06-13 | 17:37 | O94T | Site Geology | Bedrock, spotty. Seismo on bedrock. VSAT is bolted. Solar panels not. |
915 | 2021-06-13 | 17:38 | O94T | Hardware List | Cyg 536, ts 706, trd 695, seism esp T3762, 3T seismo. sspb a0389a. |
916 | 2021-06-13 | 17:38 | O94T | VSAT Check | Co/X 32/0 dB. |
917 | 2021-06-13 | 17:40 | O94T | DC Filter Check | On at 0.001Hz startup. |
918 | 2021-06-13 | 17:40 | O94T | Gain Check | 100 sps, gain 0.5, startup gain. |
919 | 2021-06-13 | 17:41 | O94T | Comments | On a decommissioned gold mine. About a KM from some workshop with minor activity. |
920 | 2021-06-13 | 17:47 | RDLO | New Station Code | Started recordoring on Naqs2 in RDLO ringbuffers. |
921 | 2021-06-13 | 17:51 | RDLO | New Station Code | Started recording on Naqs1 in RDLO ringbuffers. |
922 | 2021-06-15 | 14:54 | O99T | New Station Start | Starting in O99T buffers at Pickle Lake, PKLO |
923 | 2021-06-15 | 14:55 | O99T | Hardware List | SSPB A0393A, CYG542, TS693, TRD738, Seism. 3T, T3249 |
924 | 2021-06-15 | 14:56 | O99T | DC Filter Check | On at 0.001 Hz. |
925 | 2021-06-15 | 14:56 | O99T | Gain Check | SPS 100, gain 0.5 starting. |
926 | 2021-06-15 | 14:57 | O99T | Cold-Boot | Done. |
927 | 2021-06-15 | 14:57 | O99T | Site Geology | Exposed bedrock near old quarry (being used for gravel, no crushing at this time). |
928 | 2021-06-15 | 14:57 | O99T | VSAT Check | Co/x 32/2 dB. |
No | Date | Time UT | Location | Activity | Comments |
929 | 2021-06-15 | 14:58 | O99T | Comments | Gate, but open mostly. MNR people to check on the site. Appear not to be a security issue. |
930 | 2021-06-15 | 14:59 | O99T | Comments | Mass Centred ok. |
931 | 2021-06-15 | 15:00 | O99T | Comments | Driving to next site today. Be there tonight. |
932 | 2021-06-15 | 15:10 | PKLO | New Station Code | Started on Naqs2, Ottawa Hub, in PKLO buffers. |
933 | 2021-06-15 | 15:14 | PKLO | New Station Code | Started on Naqs1, Ottawa Hub, in PKLO buffers. |
934 | 2021-06-16 | 21:03 | O97T | New Station Start | Started at LDIO in O97T ringbuffers. |
935 | 2021-06-16 | 21:04 | O97T | Hardware List | Cyg540, ts697, trd 736, sspb a0394a, seism. 3T, T3190. |
936 | 2021-06-16 | 21:05 | O97T | Gain Check | Gain of 0.5 at sps of 100 at start. |
937 | 2021-06-16 | 21:05 | O97T | VSAT Check | Co/x: 35/4 dB. |
938 | 2021-06-16 | 21:06 | O97T | DC Filter Check | On at 0,001 Hz. |
939 | 2021-06-16 | 21:06 | O97T | Site Geology | Exposed bedrock (cleared for exploration, no longer in use). |
940 | 2021-06-16 | 21:07 | O97T | Comments | Hilty drill broke on last hole (overheated). Truck window shattered. Trailer tire burst. |
941 | 2021-06-16 | 21:14 | O97T | Seismo Check | Seismo calibration ok. |
942 | 2021-06-16 | 21:30 | LDIO | New Station Code | Started recording on Naqs2 in LDIO ringbuffers. |
943 | 2021-06-16 | 21:35 | LDIO | New Station Code | Started recording on Ottawa Naqs1 in LDIO ringbuffers. |
944 | 2021-06-16 | 21:38 | LDIO | Seismo Check | Low frequency ride on Vertical comp. |
No | Date | Time UT | Location | Activity | Comments |
945 | 2021-06-17 | 18:00 | PKRO | Service Visit | To improve VSAT power. |
946 | 2021-06-17 | 18:10 | PKRO | Comments | Weeds were blocking feedhorn and antenna. All cleaned up. Power went up by 5 dB. |
947 | 2021-06-17 | 18:20 | PKRO | VSAT Check | Antenna re-pointed. X/Co: 0/30 dB. |
948 | 2021-06-17 | 18:21 | PKRO | Comments | No additional improvement in power level. |
949 | 2021-06-17 | 18:22 | PKRO | Gain Check | Unchanged at 0.5 |
950 | 2021-06-17 | 18:23 | PKRO | DC Filter Check | Unchanged at on and 0.001 Hz. |
951 | 2021-06-17 | 18:25 | PKRO | Firmware Upgrade | All firmeware are up-to-date. |
952 | 2021-06-17 | 18:40 | PKRO | Cold-Boot | Cold boot is standard procedure at the end of visit. |
953 | 2021-06-17 | 18:50 | PKRO | Comments | Covered ground with rubber mat to prevent weeds growing. |
954 | 2021-06-18 | 18:03 | EPLO | VSAT Check | Increased throughput from 9000 to 12000 in TDMA. |
955 | 2021-06-18 | 18:03 | EPLO | VSAT Check | ReTX too high, getting too many bad packets. Why? |
956 | 2021-06-18 | 18:10 | EPLO | VSAT Check | RX: good/bad packes is now 1075/143. |
957 | 2021-06-18 | 18:20 | EPLO | VSAT Check | RX: good/bad packes is now 1782/841. In 10 mins, we had about 700 bad packets (but about 10 bad bursts only). |
958 | 2021-06-18 | 18:24 | EPLO | VSAT Check | TDMA Eirp was reduced by 2 dB. |
959 | 2021-06-18 | 18:46 | EPLO | VSAT Check | Too many bad bursts. TDMA Erip was increased by 6 dB. Now it is at 8 dB, 3 dB higher than original value before this set of tests. |
960 | 2021-06-18 | 18:54 | EPLO | VSAT Check | Far too many bad bursts. Eirp was reduced to 0.0 dB. |
No | Date | Time UT | Location | Activity | Comments |
961 | 2021-06-18 | 19:07 | EPLO | VSAT Check | Giving up on EPLO. VSAT behaving badly. |
962 | 2021-06-18 | 20:34 | O95T | New Station Start | New station at pnpo Pukaskwa park. |
963 | 2021-06-18 | 20:35 | O95T | Comments | Test name is not matched correctly with Cyg. Trident. |
964 | 2021-06-18 | 20:36 | O95T | Hardware List | Cyg 178, TS 043, Trd. 730, sspb a0396a, seismo. 3T, T3189. |
965 | 2021-06-18 | 20:36 | O95T | VSAT Check | Co/x 35/5 dB. |
966 | 2021-06-18 | 20:37 | O95T | DC Filter Check | On at 0.001 Hz. |
967 | 2021-06-18 | 20:37 | O95T | Gain Check | Gain is 0.5 startup. sps is at 100. |
968 | 2021-06-18 | 20:38 | O95T | Cold-Boot | Issued. OK. |
969 | 2021-06-18 | 20:38 | O95T | Comments | One of the 12 solar panels broken but usable, it is in use. |
970 | 2021-06-18 | 20:38 | O95T | Site Geology | Solid bedrock. |
971 | 2021-06-18 | 20:48 | PNPO | New Station Code | Started recording at Ottawa Naqs2 in PNPO ringbuffers. |
972 | 2021-06-18 | 20:57 | PNPO | Comments | Sensor Calibration ok. |
973 | 2021-06-18 | 21:03 | PNPO | New Station Code | Started recording at Ottawa Naqs1 in PNPO ringbuffers. |
974 | 2021-06-18 | 21:04 | PNPO | Comments | Observation: If Cyg/Trd do not match in naqs.stn, then the code attached to the correct Trident will tranport the data. |
975 | 2021-06-18 | 21:58 | FEDNOR2 | Comments | Team just finished the last of FedNor2 installation. On their way home CA, JD, MP, CB. Already home: AS, David and Coleman and IA. |
976 | 2021-06-19 | 17:00 | SUNO | Service Visit | To upgrade seismometer. VSAT kept on. Bus cable disconnected to swap seismo. |
No | Date | Time UT | Location | Activity | Comments |
977 | 2021-06-19 | 17:14 | SUNO | Seismo Swap | CMG-40T s/n T4776 removed, ESP s/n T3J32 installed. |
978 | 2021-06-20 | 23:00 | PNPO | Comments | Stopped sending GPS soh data. Why? |
979 | 2021-06-22 | 18:24 | ACTO | Require GPS Lock | GPS receiver installed on Libra com port, recording at London Hub. |
980 | 2021-07-07 | 12:04 | RDLO | Seismo Check | HHN component was noisy yesterday, looks ok today. |
981 | 2021-07-11 | 19:18 | PTCO | Reopened | In original location. Started recording in PTCO ringbuffers at Ottawa Hub on both NaqsServers. |
982 | 2021-07-11 | 19:23 | PTCO | Hardware List | CYG171,TRD070, TS034, SSPB a0307A, SEISM. 40T, 4401 |
983 | 2021-07-11 | 19:30 | PTCO | VSAT Check | Co/x 35/3 dB. |
984 | 2021-07-11 | 19:32 | PTCO | Cold-Boot | Will screen capture. |
985 | 2021-07-11 | 19:32 | PTCO | Comments | Lots of people on road about 50 m from seism. Station visible from the road. |
986 | 2021-07-11 | 19:32 | PTCO | Site Geology | Solid bedrock. |
987 | 2021-07-11 | 19:42 | PTCO | Gain Check | Restarted with Ontario default gain of 0.5 at 100 sps. |
988 | 2021-07-11 | 19:43 | PTCO | DC Filter Check | Restarted with Ontario default DC filter on at 0.001Hz. |
989 | 2021-07-18 | 01:40 | O28T | New Station Start | At location about 10 miles N of Castor Lake |
990 | 2021-07-18 | 01:44 | O28T | Hardware List | Cyg 560, TRD 752, TS757, SSPB A093A,Seis. esp T3J33 |
991 | 2021-07-18 | 01:45 | O28T | DC Filter Check | Off at 0.001 startup. |
992 | 2021-07-18 | 01:45 | O28T | Gain Check | Gain 1.0 at 40 sps startup. |
No | Date | Time UT | Location | Activity | Comments |
993 | 2021-07-18 | 01:45 | O28T | VSAT Check | Co/x 33/2 dB. |
994 | 2021-07-18 | 01:46 | O28T | Comments | No good location near original Castor lake, so moved to this location. |
995 | 2021-07-18 | 01:46 | O28T | Site Geology | Sand, few trees around. A few trees cut to produce view. |
996 | 2021-07-18 | 01:47 | O28T | Seismo Check | Mass centered ok. |
997 | 2021-07-18 | 01:49 | O28T | Cold-Boot | No problem. |
998 | 2021-07-18 | 01:52 | O28T | Comments | On Sunday, plan to visit ILKN and GBLN and DSMN, then to Gameti lake to install new site. |
999 | 2021-07-18 | 16:00 | ILKN | VSAT Check | Co/x was unchanged at 27/0. No re-point needed. |
1000 | 2021-07-18 | 16:19 | ILKN | Gain Check | Unchanged at 1.0. |
1001 | 2021-07-18 | 16:21 | ILKN | DC Filter Check | Unchanged off. |
1002 | 2021-07-18 | 16:30 | CTLN | New Station Code | Started recording at Ottawa Naqs2 in CTLN ringbuffers (temp site was O28T). |
1003 | 2021-07-18 | 16:35 | ILKN | Cold-Boot | Station started after cold-boot test. |
1004 | 2021-07-18 | 16:40 | CTLN | New Station Code | Started recording at Ottawa Naqs1 in CTLN ringbuffers (temp site was O28T). |
1005 | 2021-07-18 | 17:50 | ILKN | Service Visit | To check VSAT. Site was working before. All apper ok. |
1006 | 2021-07-18 | 20:00 | DSMN | Service Visit | To swap seismometer. Problem loging into Cygnus due to IP conflict (cyg ip same as laptop). |
1007 | 2021-07-18 | 20:28 | DSMN | Seismo Swap | Installed ESP T3J66 removed ESP T3F12(confirm later). |
1008 | 2021-07-18 | 20:30 | DSMN | Seismo Check | Mass centered ok. |
No | Date | Time UT | Location | Activity | Comments |
1009 | 2021-07-18 | 20:31 | DSMN | VSAT Check | Repointed. Before co/x 35/10 after 36/5 dB. It is a 2 Watt sspb. |
1010 | 2021-07-18 | 20:49 | DSMN | Seismo Check | Seismo calibration ok. New seismo ok, then. |
1011 | 2021-07-18 | 20:53 | DSMN | Seismo Check | Calibration at 10 sec worked as well as 5 sec cal. All seem to be ok with new seism. |
1012 | 2021-07-18 | 20:55 | DSMN | Comments | Refer to Activity line 539 on July 18, 03 for seismometer history. |
1013 | 2021-07-20 | 19:59 | O21T | New Station Start | Started new station at Gameti Lake. |
1014 | 2021-07-20 | 20:02 | O21T | Hardware List | Cyg 558, TS 759, TRD 757, SSPB A0407A, seism. ESP, T3J18. |
1015 | 2021-07-20 | 20:04 | O21T | DC Filter Check | Started with DC Filter off, NWT default. |
1016 | 2021-07-20 | 20:04 | O21T | Gain Check | Started with sps 40 and gain 1.0 default nwt. |
1017 | 2021-07-20 | 20:04 | O21T | VSAT Check | Co/x 25/2 dB. |
1018 | 2021-07-20 | 20:05 | O21T | Site Geology | Near airport, sand, rock, esker. |
1019 | 2021-07-20 | 20:06 | O21T | Comments | Airport nearby is used a few times/day. |
1020 | 2021-07-20 | 20:06 | O21T | Comments | One of the 20 amp breaker on solar input is faulty. |
1021 | 2021-07-20 | 20:07 | O21T | SSPB Check | SSPB gain table set according to sspb profile. |
1022 | 2021-07-20 | 20:24 | GALN | New Station Code | Started on Naqs2, Ottawa Hub in GALN ringbuffers. |
1023 | 2021-07-20 | 20:29 | GALN | New Station Code | Started on Naqs1, Ottawa Hub in GALN ringbuffers. |
1024 | 2021-07-21 | 17:55 | PTCO | Service Visit | Stoped to change seismo and trident. |
No | Date | Time UT | Location | Activity | Comments |
1025 | 2021-07-21 | 18:00 | PTCO | Seismo Swap | ESP T3K14 installed. CMG-40T 4401 removed. |
1026 | 2021-07-21 | 18:10 | PTCO | DC Filter Check | DC filter is on at 0.001 HZ, unchanged. |
1027 | 2021-07-21 | 18:10 | PTCO | Gain Check | Gais is 0.5 is unchanged. |
1028 | 2021-07-21 | 18:10 | PTCO | Hardware Swap | Trident 070 removed, TRD747 installed. |
1029 | 2021-07-21 | 18:47 | PTCO | Comments | Cygnus ringbuffers restarted. |
1030 | 2021-07-21 | 18:49 | PTCO | Cold-Boot | Cold Boot. |
1031 | 2021-07-21 | 18:50 | PTCO | Comments | Problem restarting VSAT after cold boot. Source unknown. Increasing TX power did not help. |
1032 | 2021-07-21 | 19:10 | PTCO | Restart | All OK and station started. |
1033 | 2021-07-21 | 19:12 | PTCO | Comments | Trident gain and DC filter checked again, they are unchanged after cold boot. |
1034 | 2021-07-23 | 16:50 | ELFO | New Station Code | New Internet station started, served by UWO hub. |
1035 | 2021-07-23 | 16:51 | ELFO | Hardware List | Cyg 564, Trd 775, ESP seism. T3J88. |
1036 | 2021-07-23 | 16:52 | ELFO | Gain Check | Started with standard gain of 0.5 at 100 sps. |
1037 | 2021-07-23 | 16:53 | ELFO | DC Filter Check | Started with standard Ontario DC filter on at 0.001 Hz. |
1038 | 2021-07-23 | 16:54 | ELFO | Comments | Mass center and calibration ok but not able to log to the station. |
1039 | 2021-07-30 | 20:17 | CBRQ | New Station Code | Started first Quebec station at Cabonga Reservoir. |
1040 | 2021-07-30 | 20:18 | CBRQ | Comments | Served by Ottawa Hub. Naqs2 recording started 20:17, Naqs1 at 20:21. |
No | Date | Time UT | Location | Activity | Comments |
1041 | 2021-07-30 | 20:19 | CBRQ | Hardware List | Cyg 561, TS 713, TRD 754, ESP Seism. T3K12, sspb a0496a. |
1042 | 2021-07-30 | 20:20 | CBRQ | VSAT Check | Co/X 35/4 dB. |
1043 | 2021-07-30 | 20:21 | CBRQ | Site Geology | Solid bedrock. |
1044 | 2021-07-30 | 20:22 | CBRQ | Cold-Boot | Before taking screen captures. |
1045 | 2021-07-30 | 20:23 | CBRQ | Gain Check | Started with gain of 0.5 at 100 sps. |
1046 | 2021-07-30 | 20:24 | CBRQ | DC Filter Check | Started with dc filter on at 0.001 hz. |
1047 | 2021-07-30 | 20:25 | CBRQ | Comments | Ringbuffers restarted. |
1048 | 2021-07-31 | 18:54 | O30T | Service Visit | Visited HBDB to swap bad cygnus. Recording at O30T ringbuffer for now. |
1049 | 2021-07-31 | 18:55 | O30T | VSAT Check | Co/X 33/2 dB. Minor adjustment was done. |
1050 | 2021-07-31 | 18:56 | HBDB | Hardware Swap | Installed Cyg 562, TS 711, Trd 760. Removed Cyg 161, Trd 064. |
1051 | 2021-07-31 | 18:57 | HBDB | DC Filter Check | Unchaged at on at 0.001 Hz. |
1052 | 2021-07-31 | 18:57 | HBDB | Gain Check | Unchaged at 0.5. |
1053 | 2021-07-31 | 18:58 | HBDB | Comments | Condition of the site was ok at the time of visit. |
1054 | 2021-07-31 | 18:58 | HBDB | Comments | Ringbuffers restarted. |
1055 | 2021-07-31 | 19:10 | HBDB | Comments | Switched O30T to HBDB in naqs.stn files on Ottawa Naqs2 and Naqs1. Station currently down for final adjustments. |
1056 | 2021-07-31 | 19:13 | HBDB | Comments | Cold boot. Ground slump near the base on the King Post. May be problematic. |
No | Date | Time UT | Location | Activity | Comments |
1057 | 2021-08-02 | 21:30 | MPPO | Service Visit | To fix VSAT problem. GPS engine suspected defective. Swapping Cygnus. |
1058 | 2021-08-02 | 21:35 | MPPO | Hardware Swap | Cyg 268 removed, Cyg 563 installed. |
1059 | 2021-08-02 | 21:35 | MPPO | Hub Swap | Moved from London to Ottawa Hub. |
1060 | 2021-08-02 | 21:45 | OTTAWA | Comments | Naqs1 is down. |
1061 | 2021-08-02 | 22:06 | MPPO | Restart | Station started at Ottawa Hub, Naqs2. |
1062 | 2021-08-02 | 22:08 | MPPO | Cold-Boot | Standard procedure. |
1063 | 2021-08-02 | 22:15 | MPPO | Comments | Site in perfect condition. Old Cygnus GPS tracking but not locking. |
1064 | 2021-08-02 | 22:15 | MPPO | DC Filter Check | Unchaged at on at 0.001 Hz. |
1065 | 2021-08-02 | 22:15 | MPPO | Gain Check | Unchanged at 0.5 |
1066 | 2021-08-03 | 00:57 | OTTAWA | Comments | Naqs1 restarted after system fixup. |
1067 | 2021-08-03 | 00:58 | MPPO | New Station Code | Started recording in Naqs1 ringbuffers. |
1068 | 2021-08-05 | 17:45 | O23T | New Station Start | Started new station at Restoule Prov. Park. |
1069 | 2021-08-05 | 17:46 | O23T | Gain Check | Cyg. 553, Trd. 750, TS 667, sspb A0399A, seis. esp. T3K46. |
1070 | 2021-08-05 | 17:47 | O23T | Gain Check | Started with gain of 0.5 sps 100. |
1071 | 2021-08-05 | 17:47 | O23T | VSAT Check | Co/x: 34/4 dB. |
1072 | 2021-08-05 | 17:48 | O23T | DC Filter Check | Started with On at 0.001 Hz. |
No | Date | Time UT | Location | Activity | Comments |
1073 | 2021-08-05 | 17:49 | O23T | Cold-Boot | Did as standard procedure. |
1074 | 2021-08-05 | 17:49 | O23T | Comments | Cleared ringbuffers. |
1075 | 2021-08-05 | 17:49 | O23T | Site Geology | Solid bedrock. |
1076 | 2021-08-05 | 17:50 | O23T | Comments | Started at Ottawa Hub. |
1077 | 2021-08-05 | 17:50 | O23T | Power Check | 8 solar panels and 10 batteries installed. |
1078 | 2021-08-05 | 17:59 | RSPO | New Station Code | Started at Ottawa Hub, Naqs2 ringbuffers. |
1079 | 2021-08-05 | 18:00 | TWBB | Service Visit | Station vandalized, battery box left open and was full of water. Battery positive terminal corroded. Positive lead disconnected. Box drained. Batteries cleaned and connected. Appear to be working. |
1080 | 2021-08-05 | 18:03 | RSPO | New Station Code | Started at Ottawa Hub, Naqs1 ringbuffers. |
1081 | 2021-08-05 | 19:40 | TWBB | VSAT Check | Co/X 34/5 dB. |
1082 | 2021-08-05 | 19:42 | TWBB | Cold-Boot | Standard procedure. |
1083 | 2021-08-05 | 19:49 | TWBB | Restart | Station restarted, all OK. |
1084 | 2021-08-05 | 20:15 | TWGB | Service Visit | On the way back from TWBB, power cable from junction box bitten but OK. Loose connector fixed. |
1085 | 2021-08-07 | 19:21 | O25T | New Station Start | Started new station at Morin Heights, MRHQ. |
1086 | 2021-08-07 | 19:22 | O25T | Hardware List | Cyg 556, ts 615, trd 772, sspb a0391a, esp. T3K13. |
1087 | 2021-08-07 | 19:23 | O25T | Site Geology | Solid bedrock. |
1088 | 2021-08-07 | 19:24 | O25T | Cold-Boot | Standard procedure. |
No | Date | Time UT | Location | Activity | Comments |
1089 | 2021-08-07 | 19:25 | O25T | VSAT Check | Co/x 35/5 dB. |
1090 | 2021-08-07 | 19:26 | O25T | Gain Check | Started with gain of 0.5 at 100 sps. |
1091 | 2021-08-07 | 19:27 | O25T | DC Filter Check | Started with dc filter on at 0.001 hz. |
1092 | 2021-08-07 | 19:28 | O25T | Comments | This is an AC site, served by Ottawa Hub. |
1093 | 2021-08-07 | 19:29 | O25T | Comments | Short trees planted to onscure the VSAT dish. |
1094 | 2021-08-07 | 19:45 | MRHQ | New Station Code | Started recording at Ottawa Naqs2 in MRHQ ringbuffers. |
1095 | 2021-08-07 | 19:50 | MRHQ | New Station Code | Started recording at Ottawa Naqs1 in MRHQ ringbuffers. |
1096 | 2021-08-07 | 23:24 | OTTAWA HUB | Comments | ReTX turned on on Naqs1. It was already on on Naqs2. |
1097 | 2021-08-09 | 17:59 | OTTAWA HUB | Comments | Naqs1 was down for service from 13:59 to 14:35 ET. |
1098 | 2021-08-13 | 14:59 | LGSN | Hardware Swap | Switched from London to Ottawa Hub. |
1099 | 2021-08-13 | 16:00 | LGSN | Service Visit | Animal damage. Feed horn bent, feed window pecked. Both fixed. |
1100 | 2021-08-13 | 16:01 | LGSN | VSAT Check | Co/X 23/0 dB. |
1101 | 2021-08-13 | 16:02 | LGSN | Comments | Need large wrench for next visit to tighten antenna. |
1102 | 2021-08-13 | 16:43 | LGSN | Restart | Station restarted, Ottawa Hub, both Naqsservers. |
1103 | 2021-08-14 | 14:30 | MCKN | Service Visit | Trident not communicating to NMX bus, replaced. |
1104 | 2021-08-14 | 14:31 | MCKN | Comments | Trd 137 had failed on July 30, 2004. Never recovered. |
No | Date | Time UT | Location | Activity | Comments |
1105 | 2021-08-14 | 14:31 | MCKN | Hardware Swap | Trident 137 removed, 746 installed. |
1106 | 2021-08-14 | 14:32 | MCKN | Seismo Swap | ESP T3J30 installed. 40T T4938 removed |
1107 | 2021-08-14 | 15:50 | MCKN | Restart | Restarted recording, Ottawa. |
1108 | 2021-08-14 | 15:51 | MCKN | Cold-Boot | Standard procedure. |
1109 | 2021-08-14 | 15:52 | MCKN | Gain Check | Unchanged at 1.0 and sps of 40. |
1110 | 2021-08-14 | 15:53 | MCKN | DC Filter Check | Unchanged at not on. |
1111 | 2021-08-14 | 18:17 | MCKN | Comments | Started recording MT data on both Naqsservers at Ottawa. |
1112 | 2021-08-14 | 18:38 | MCKN | Comments | MT Nims instrument configured at low gain with filter on DC. |
1113 | 2021-08-14 | 18:44 | MCKN | Comments | Old 486 computer used for MT setup has failed. |
1114 | 2021-08-16 | 17:34 | ENGB | Comments | GPS sats were not visible, but system worked. |
1115 | 2021-08-16 | 17:34 | ENGB | Firmware Upgrade | Upgrade to Libra V57005B and TS2000. |
1116 | 2021-08-16 | 17:34 | ENGB | Service Visit | To figure out the gap problem. |
1117 | 2021-08-16 | 17:36 | ENGB | Comments | GPS stas are visible now. |
1118 | 2021-08-16 | 17:36 | ENGB | Comments | No change to Trident firmware. |
1119 | 2021-08-16 | 17:37 | ENGB | Cold-Boot | Standard procedure. |
1120 | 2021-08-16 | 17:37 | ENGB | VSAT Check | X/Co 0/22 dB. Unchaged but checked. |
No | Date | Time UT | Location | Activity | Comments |
1121 | 2021-08-16 | 17:47 | ENGB | Gain Check | Unchanged at 0.5. |
1122 | 2021-08-16 | 17:48 | ENGB | Comments | Firmware upgrade appear to have helped. |
1123 | 2021-08-16 | 17:48 | ENGB | DC Filter Check | Unchaged from On at 1 mHz. |
1124 | 2021-08-17 | 14:00 | MALO | DC Filter Check | Unchaged at on and 1 mHz. |
1125 | 2021-08-17 | 14:00 | MALO | VSAT Check | Co/x unchanged at 29/0 dB. |
1126 | 2021-08-17 | 14:18 | MALO | Service Visit | To upgrade seismometer. |
1127 | 2021-08-17 | 14:19 | MALO | Comments | There is a new locked gate on the way to the site. However, battery box security cable was cut open but box was ok. Wire to one solar panel was cut, fixed. |
1128 | 2021-08-17 | 14:20 | MALO | Seismo Swap | ESP T3J87 installed, 40T T4774 removed. |
1129 | 2021-08-17 | 14:52 | MALO | Cold-Boot | To make sure all config. parameters are saved correctly. |
1130 | 2021-08-17 | 14:53 | MALO | Restart | Station restarted. Both Ottawa Hub Naqsservers receive data. |
1131 | 2021-08-17 | 14:55 | MALO | Gain Check | Unchaged at 1.0 (this is default for FedNor1). |
1132 | 2021-08-17 | 15:34 | MALO | Comments | Some new mining/drilling activity near by (about 2 km). |
1133 | 2021-08-18 | 14:00 | OTRO | Service Visit | To swap seismometer. Site condition good. |
1134 | 2021-08-18 | 14:00 | OTRO | VSAT Check | Co/x at 29/0, unchanged. |
1135 | 2021-08-18 | 14:30 | OTRO | Seismo Swap | ESP T3H24 installed, 40T T4767 removed. |
1136 | 2021-08-18 | 15:06 | OTRO | Cold-Boot | Standard procedure. |
No | Date | Time UT | Location | Activity | Comments |
1137 | 2021-08-18 | 15:07 | OTRO | Restart | Station restarted, recording on Ottawa Hub Naqs2. |
1138 | 2021-08-18 | 16:02 | OTRO | Gain Check | Unchanged at 1.0, default for FedNor1 sites. |
1139 | 2021-08-18 | 16:03 | OTRO | Comments | Hunters camping nearby. |
1140 | 2021-08-18 | 16:03 | OTRO | DC Filter Check | Unchanged at on at 0.001 Hz. |
1141 | 2021-08-18 | 16:04 | OTRO | Comments | Site gurad hired to clear snow from dish in the winter, watch site. |
1142 | 2021-08-18 | 16:07 | OTRO | Restart | Station recording on Ottawa Hub Naqs2 with the new seismometer. |
1143 | 2021-08-18 | 16:08 | MALO | Comments | Security screws installed on battery boxes. |
1144 | 2021-08-18 | 16:30 | MGTN | Service Visit | Trident found dead. Swapped. |
1145 | 2021-08-18 | 17:05 | MGTN | Hardware Swap | Trident 694 installed, 178 removed. |
1146 | 2021-08-18 | 17:08 | MGTN | Hub Swap | Station moved from London to Ottawa Hub. |
1147 | 2021-08-18 | 20:00 | MGTN | Hardware Swap | Cygnus 261 removed, 557 installed. |
1148 | 2021-08-18 | 20:01 | MGTN | VSAT Check | Poor VSAT. Swapped Cygnus, no help. Did a re-point, improved Co/x from 13/0 to 26/0. |
1149 | 2021-08-18 | 20:02 | MGTN | Comments | Cyg 261 removed from MGTN might have been ok. |
1150 | 2021-08-19 | 13:11 | KILO | Service Visit | To upgrade seismometer. |
1151 | 2021-08-19 | 13:12 | KILO | Seismo Swap | ESP T3D69 installed, 40T T4186 removed. |
1152 | 2021-08-19 | 14:25 | KILO | Cold-Boot | Standard procedure. |
No | Date | Time UT | Location | Activity | Comments |
1153 | 2021-08-19 | 14:26 | KILO | Restart | With new seismometer, recording at Ottawa Hub on Naqs2. |
1154 | 2021-08-19 | 14:27 | KILO | Gain Check | Unchanged at 1. |
1155 | 2021-08-19 | 14:28 | KILO | DC Filter Check | Unchanged at on and 0.001 Hz. |
1156 | 2021-08-19 | 14:29 | KILO | VSAT Check | Co/x at 20/0 dB, unchanged. |
1157 | 2021-08-19 | 14:47 | KILO | Seismo Check | Vertical component of new seismometer appeared to have been unlocked during transportation. However, all channels mass-centred ok and calibrated ok. Calibration singals at 1, 5 and 10 seconds around 14:47 UT look normal. |
1158 | 2021-08-19 | 14:50 | KILO | Comments | Mouse droppings in seismo vault. |
1159 | 2021-08-19 | 14:51 | KILO | Comments | Seismo may have a bad bubble, used external bubble to level. Inner vault cover was not installable but insolation covers installed. Should be adequate. |
1160 | 2021-08-19 | 14:52 | KILO | Comments | Black bears spotted nearby but not bothering the site. |
1161 | 2021-08-19 | 15:13 | KILO | Comments | Naqs1 restarted with new seismomter. |
1162 | 2021-08-19 | 17:31 | ANMB | Closed | Station closed at 2004-08-19-17-31. Site owner requested removal of station since the property is sold. |
1163 | 2021-08-19 | 20:18 | MGTN | Comments | Starting MT data at Ottawa Hub on Naqs2. |
1164 | 2021-08-19 | 20:25 | MGTN | Comments | Starting MT data at Ottawa Hub on Naqs1. |
1165 | 2021-08-20 | 13:48 | SUNO | Service Visit | To add four more solar panels. |
1166 | 2021-08-20 | 14:03 | SUNO | Restart | Station restarted adding solar panels. A total of 12 panels power the system now. |
1167 | 2021-08-20 | 14:04 | SUNO | VSAT Check | Co/X 33/1 dB, unchanged. |
1168 | 2021-08-20 | 16:04 | SUNO | Cold-Boot | Standard procedure. |
No | Date | Time UT | Location | Activity | Comments |
1169 | 2021-08-20 | 16:06 | SUNO | Gain Check | Unchaged at 1.0 (this is default for FedNor1). |
1170 | 2021-08-20 | 16:07 | SUNO | DC Filter Check | Unchaged at on and 1 mHz. |
1171 | 2021-08-20 | 16:08 | SUNO | Comments | Station in good condition. |
1172 | 2021-08-20 | 17:30 | NODN | Service Visit | Site was down due to a dead Cygnus. No other damage to anything. |
1173 | 2021-08-20 | 17:31 | NODN | Hardware Swap | Dead Cygnus 395 removed, Cyg261 installed. |
1174 | 2021-08-20 | 17:32 | NODN | Comments | Com ports not working for Cyg261 but they are not in use. |
1175 | 2021-08-20 | 17:33 | NODN | Re-Point | Co/X was 18/0 before re-point and is now 23/0 dB. |
1176 | 2021-08-20 | 17:34 | NODN | Restart | Recording at Ottawa Hub, Naqs2. |
1177 | 2021-08-20 | 18:27 | NODN | Restart | Ottawa Hub Naqs1 restarted. |
1178 | 2021-08-20 | 18:37 | NODN | Comments | No change to gain or dc filter. |
1179 | 2021-08-20 | 20:20 | KNDN | Service Visit | To re-point antenna and borrow Cygnus. |
1180 | 2021-08-20 | 20:21 | KNDN | Re-Point | Co/X improved from 30/5 to 30/1 dB. |
1181 | 2021-08-20 | 20:25 | KNDN | Hardware Swap | Cygnus 419 removed, Cyg 557 installed. |
1182 | 2021-08-20 | 20:26 | KNDN | Comments | Cyg 557 has bad com ports but they are not used here. |
1183 | 2021-08-20 | 20:27 | KNDN | Comments | New Cyg added to Naqs2, Ottawa Hub. |
1184 | 2021-08-20 | 20:31 | KNDN | Comments | New Cyg added to Naqs1, Ottawa Hub. |
No | Date | Time UT | Location | Activity | Comments |
1185 | 2021-08-20 | 20:42 | KNDN | Restart | With new Cygnus 557. |
1186 | 2021-08-20 | 21:10 | ELFO | Comments | Cygnus#564 for this Internet node was incorrectly left in TDMA for Ottawa Hub. Removed. |
1187 | 2021-08-20 | 21:43 | KNDN | Comments | New Cyg added to Naqs1, Ottawa Hub. |
1188 | 2021-08-20 | 21:45 | KNDN | Comments | Lost some data and is not doing ReTX right now. Is it due to Cygnus 557 buffers not flushed? |
1189 | 2021-08-21 | 16:12 | SNPN | Service Visit | To add MT. |
1190 | 2021-08-21 | 16:13 | SNPN | Hardware Swap | Moved from London to Ottawa Hub. |
1191 | 2021-08-21 | 16:16 | SNPN | Restart | Recording at Ottawa Hub on Naqs1 and Naqs2. |
1192 | 2021-08-21 | 16:26 | SNPN | Comments | MT installation in progress. |
1193 | 2021-08-21 | 20:50 | SNPN | Comments | MT data started. Temperature option of SOH monitor disconnected. |
1194 | 2021-08-21 | 20:59 | SNPN | Comments | MT team are digging electric lines. |
1195 | 2021-08-22 | 14:56 | ACKN | Restart | Trident 141 shows bus gaps, rebooted to reset negative counter on "Bytes sent". |
1196 | 2021-08-23 | 15:11 | SNPN | Comments | MT station was found to be in GPS mode. Switched to normal mode by Zoc telenet login, and then ^C command. |
1197 | 2021-08-24 | 15:00 | GBLN | Service Visit | To bring the station up and install MT. |
1198 | 2021-08-24 | 15:01 | GBLN | Wildlife | Tarps missing on battery boxes, one shredded nearby suspect animal attack, big hole in feed window. |
1199 | 2021-08-24 | 15:30 | GBLN | Repair | Feed window replaced with a reconditioned one. |
1200 | 2021-08-24 | 15:32 | GBLN | VSAT Check | Co/x improved from 17/0 to 20/0 dB. |
No | Date | Time UT | Location | Activity | Comments |
1201 | 2021-08-24 | 15:33 | GBLN | Seismo Swap | ESP T3K45 installed, 40T T4316 removed. |
1202 | 2021-08-24 | 15:34 | GBLN | Restart | Ottawa Hub Naqs2 restarted. |
1203 | 2021-08-24 | 15:35 | GBLN | Restart | Ottawa Hub Naqs1 restarted. |
1204 | 2021-08-24 | 15:47 | GBLN | Comments | Both Naqs1 and Naqs2 of Ottawa Hub ready for MT. |
1205 | 2021-08-24 | 15:50 | GBLN | Comments | Calibration and Mass-centre ok. |
1206 | 2021-08-24 | 16:39 | GBLN | Comments | MT installation in progress. |
1207 | 2021-08-26 | 18:00 | OTTAWA HUB | VSAT Start | Started Anik E2R Service on second 3.8 m antenna, Carina 92. |
1208 | 2021-08-26 | 18:10 | OTTAWA HUB | VSAT Check | Co-Pol of 35 dB achieved. |
1209 | 2021-08-26 | 18:10 | OTTAWA | Comments | Tx power adjusted to 28.4 dB as per Telesat contract. Sent a test signal with BPSK32_1_2 and power of 39 to reach Tx gain of 4.75 for this calibration of power. |
1210 | 2021-08-26 | 20:56 | O22T | VSAT Start | Started recording test station in Ottawa on E2R satellite. |
1211 | 2021-09-01 | 18:00 | PKRO | Service Visit | To fix battery charger. Fuse was blown in charger and charger was bust. Replacing fuse did not fix the problem. Charger replaced. |
1212 | 2021-09-01 | 18:15 | PKRO | Battery Check | Old, likely bad, DC-Type batteries removed, new AC-Type batteries installed. |
1213 | 2021-09-01 | 19:14 | PKRO | Restart | Station back on. Shutting down again to check VSAT. |
1214 | 2021-09-01 | 19:40 | PKRO | VSAT Check | X pol 0 dB, co-pol not measured due to modulated signal. |
1215 | 2021-09-01 | 20:56 | PKRO | Gain Check | Unchaged at 0.5. |
1216 | 2021-09-01 | 20:57 | PKRO | DC Filter Check | Unchaged at on at 0.001 Hz. |
No | Date | Time UT | Location | Activity | Comments |
1217 | 2021-09-01 | 20:58 | PKRO | Comments | Fuses left at the site for battery charger. |
1218 | 2021-09-02 | 13:10 | TORO | Service Visit | To improve bad VSAT power. |
1219 | 2021-09-02 | 13:30 | TORO | Re-Point | Antenna was off. Before 20 dB. After, 31/0 dB. |
1220 | 2021-09-02 | 13:47 | TORO | Comments | Non-POLARIS lock on battery box. BAD!!! |
1221 | 2021-09-02 | 18:30 | O43T | New Station Start | Starting new station at Igloolik Airport. Installation in progress. |
1222 | 2021-09-02 | 18:45 | O43T | Comments | Too cold for cement to set on seismo vault. Waiting. |
1223 | 2021-09-02 | 19:00 | O43T | Power Check | Powered by AC from nearby airport shack. |
1224 | 2021-09-02 | 19:10 | O43T | Repair | 2 W SSPB failed. Replaced with one Watt unit. |
1225 | 2021-09-02 | 21:08 | IGLN | New Station Code | Started recording at Ottawa Naqs2 in IGLN ringbuffers. Served by Telesat Anik E2R. |
1226 | 2021-09-02 | 21:09 | IGLN | DC Filter Check | Started with DC filter off. |
1227 | 2021-09-02 | 21:09 | IGLN | Gain Check | Started with gain of 1.0 at 40 sps. |
1228 | 2021-09-02 | 21:10 | IGLN | VSAT Check | Co/X at 31/1 dB. |
1229 | 2021-09-02 | 21:17 | IGLN | Seismo Check | Mass Center and Calibration OK. |
1230 | 2021-09-02 | 21:18 | IGLN | Hardware List | Cyg 284, Trd 770, ESP T3K44. |
1231 | 2021-09-02 | 21:53 | IGLN | VSAT Check | Operating with BPSK, 32 KB. Eirp increase 6 dB, bad/good packet ration 118/6792. |
1232 | 2021-09-02 | 21:57 | IGLN | New Station Code | Started recording at Ottawa Naqs1 in IGLN ringbuffers. |
No | Date | Time UT | Location | Activity | Comments |
1233 | 2021-09-02 | 22:03 | IGLN | VSAT Check | Bad/good ratio 156/7703. Eirp increased by 1 dB. |
1234 | 2021-09-04 | 21:17 | O42T | New Station Start | Started new DC station at Qililugaq Exploration Camp, QILN, served by Ottawa Hub Antenna 2, Telesat. |
1235 | 2021-09-04 | 21:18 | O42T | VSAT Check | Co-pol isolation 30-32 dB. Good. |
1236 | 2021-09-04 | 21:19 | O42T | Comments | Minor damage to antenna edge and back during transportation. Front ok. |
1237 | 2021-09-04 | 22:11 | QILN | New Station Code | Started recording at Ottawa Naqs2 in QILN ringbuffers. |
1238 | 2021-09-04 | 22:14 | QILN | New Station Code | Started recording at Ottawa Naqs1 in QILN ringbuffers. |
1239 | 2021-09-04 | 22:15 | QILN | Gain Check | Gain of 1.0 at 40 sps. |
1240 | 2021-09-04 | 22:16 | QILN | DC Filter Check | DC filter not on. |
1241 | 2021-09-04 | 22:54 | QILN | VSAT Check | Eirp increase up from 4 to 6 dB to improve TX power. TX DAC increased from 10.50 to 16.25. |
1242 | 2021-09-04 | 22:56 | QILN | Hardware List | Cygnus 555, Trident 737, TimeServer 675. Seismo. ESP T3K94, SSPB A0397A. |
1243 | 2021-09-06 | 22:08 | O22T | New Station Start | Started new AC station at Rankin Inlet Airport, YRTN, served by Ottawa Hub, Antenna 2, Telesat Anik E2R. |
1244 | 2021-09-06 | 22:10 | O22T | VSAT Check | 32 dB of isolation. Good. |
1245 | 2021-09-06 | 22:31 | YRTN | New Station Code | Started recording at Ottawa Naqs2 in YRTN ringbuffers. |
1246 | 2021-09-06 | 22:33 | YRTN | New Station Code | Started recording at Ottawa Naqs1 in YRTN ringbuffers. |
1247 | 2021-09-06 | 22:35 | YRTN | Hardware List | Cygnus 550, Trident 767, TimeServer 633, sspb A0400A. ESP . |
1248 | 2021-09-06 | 22:36 | YRTN | Seismo Check | Bottom of seismometer vault was wet. |
No | Date | Time UT | Location | Activity | Comments |
1249 | 2021-09-06 | 22:37 | YRTN | Gain Check | Started with gain of 1.0 at 40 sps. |
1250 | 2021-09-06 | 22:38 | YRTN | DC Filter Check | Started with DC filter off. |
1251 | 2021-09-12 | 17:19 | ACKN | Firmware Upgrade | Libra 5.7005A installed. |
1252 | 2021-09-12 | 17:24 | ACKN | Firmware Upgrade | TS 2.0000B installed. Trident 1.7210 already running. |
1253 | 2021-09-12 | 17:26 | ACKN | Restart | Trident rebooted. Transmit errors before reboot 9640, after 0. |
1254 | 2021-09-12 | 17:29 | ACKN | Comments | Monitoring station for bus gaps. Would firmware upgrade help? |
1255 | 2021-09-14 | 03:00 | MCKN | Comments | Trident 746 is not communicating to Cygnus. No seismic data. This is second time that Cygnus 219 at this site is failing to service the NMX bus or causing Trident to die. |
1256 | 2021-09-15 | 16:41 | YMBN | MT Station | Libra ports not showing any read or write. MT data stopped. |
1257 | 2021-09-15 | 17:30 | TYNO | Require GPS Lock | GPS receiver installed on Libra com port, recording at London Hub. |
1258 | 2021-09-15 | 18:00 | MCKN | Restart | Logged to MCKN over VSAT, Trident not showing. Reboot TimeServer in an attempt to power-cycle NMX bus. |
1259 | 2021-09-15 | 18:07 | MCKN | Restart | Reboot Cygnus in attempt to restart the Trident. |
1260 | 2021-09-15 | 18:20 | MCKN | Comments | Booting TimeServer and Cygnus did not fix the problem. Trident is not communicating. On TS Instrument screen, bytes sent are increasing, bytes/packets received continue to be zero. |
1261 | 2021-09-17 | 13:14 | ACKN | Comments | Bus gaps continue. Trident 141 Instrument screen shows Transmit overruns of 3361 and errors of 3899 and receive overruns of 5708 after about 5 days since last reboot of TimeServer. |
1262 | 2021-09-17 | 13:16 | NWT | VSAT Check | Bandwidth increased for YNEN and GBLN, decreased for MCKN. |
1263 | 2021-09-17 | 13:51 | YMBN | MT Station | Rebooting Cygnus did not restart the MT instrument . MT Station is down. |
1264 | 2021-09-17 | 14:00 | YMBN | MT Station | Rebooted MT instrument by sending Ctrl-C on telnet port. Station restarted. |
No | Date | Time UT | Location | Activity | Comments |
1265 | 2021-09-29 | 16:00 | RDLO | Closed | Moved to new location RLKO on Nov. 19, 2004. |
1266 | 2021-10-06 | 16:40 | IGLN | Service Visit | To upgrade SSPB from 1 W to 2 Watt and re-point dish. |
1267 | 2021-10-06 | 16:41 | IGLN | SSPB Check | 2 Watt installed, 1 Watt removed. |
1268 | 2021-10-06 | 16:42 | IGLN | VSAT Check | Isolation 30.0 dB. Eirp increase on 7 dB on Hub 9. |
1269 | 2021-10-06 | 16:45 | IGLN | Gain Check | Unchaged at 1 and 40 sps. |
1270 | 2021-10-06 | 16:46 | IGLN | DC Filter Check | Unchaged at no DC filter. |
1271 | 2021-10-07 | 15:40 | YRTN | Service Visit | To replace seismometer, repoint |
1272 | 2021-10-07 | 15:41 | YRTN | Comments | Seismo vault flooded. seimo sub-merged. Vault re-built with raised pier. |
1273 | 2021-10-07 | 15:42 | YRTN | Hardware Swap | Cyg 217, Trd 070, ESP T3L28 installed. New sspb a0498a . |
1274 | 2021-10-07 | 15:43 | YRTN | VSAT Check | Faild sspb replaced. Power is good. Eirp inc at +2 dB. |
1275 | 2021-10-07 | 15:45 | YRTN | Gain Check | Unchaged at 1 and 40 sps. |
1276 | 2021-10-07 | 15:46 | YRTN | DC Filter Check | Unchaged at no DC filter. |
1277 | 2021-11-19 | 00:00 | OTTAWA | Comments | Activity log is out of date. Need update. |
1278 | 2021-11-19 | 15:36 | RLKO | New Station Code | Started recording at both Naqs in Ottawa Hub. |
1279 | 2021-12-14 | 16:00 | ALFO | Service Visit | To re-point and charge batteries. |
1280 | 2021-12-14 | 18:20 | ALFO | Re-Point | X/Co. 28/3 before, 32/2 dB after repoint. |
No | Date | Time UT | Location | Activity | Comments |
1281 | 2021-12-14 | 18:54 | ALFO | Comments | All is ok. Trident screen was checked. Overruns looks ok. |
1282 | 2021-12-14 | 18:55 | ALFO | Battery Check | Charging in progress. |
1283 | 2021-12-14 | 20:06 | ALFO | Service Visit | End of Serivice. Batteries were charged for about 4 houurs. All ok. |
1284 | 2021-12-16 | 16:00 | WLVO | Service Visit | To install DC power system. AC power not available any more. |
1285 | 2021-12-16 | 19:30 | WLVO | Reopened | Station started with solar power. No change in configuration. |
1286 | 2021-12-16 | 19:56 | WLVO | VSAT Check | No change at 30/0 dB. |
1287 | 2021-12-22 | 16:00 | CLWO | Service Visit | To clear antenna of ice and snow, check power. |
1288 | 2021-12-22 | 18:00 | CLWO | Comments | Too hard to walk alone to the site (KA) with no snow shoes. |
1289 | 2021-12-22 | 18:30 | CLWO | Power Check | Battery volts low, need sun for charging. |
1290 | 2021-12-31 | 20:00 | CLVB | Service Visit | To add a second battery box and 6 batteries. |
1291 | 2021-12-31 | 20:10 | CLVB | VSAT Check | Was good at 35/5 dB. |
1292 | 2022-01-13 | 18:00 | MEDO | Service Visit | To upgrade and fix power system |
1293 | 2022-01-13 | 18:10 | MEDO | Power Check | Added 2 batteries. LVD reset from 12 VDC to 11.5 VDC. |
1294 | 2022-01-13 | 18:20 | MEDO | Re-Point | Before x/c was 32/2, after repoint it is 32/1 dB. |
1295 | 2022-01-13 | 18:30 | MEDO | Comments | Electric fence installed by land owner. |
1296 | 2022-01-14 | 15:00 | ELFO | Service Visit | To install new charge controller and 2 batteries. |
No | Date | Time UT | Location | Activity | Comments |
1297 | 2022-01-25 | 14:00 | CARINA 92 | Comments | Ottawa Hub, Telesat antenna, stations on Hub 92 not working, see SOH graphs. |
1298 | 2022-01-25 | 15:00 | CARINA 92 | VSAT Check | Ottawa Hub, Telesat antenna, No RX. |
1299 | 2022-01-25 | 15:10 | CARINA 92 | SSPB Check | SSPB and LNB replaced. |
1300 | 2022-01-25 | 15:30 | CARINA 92 | Re-Point | Telesat antenna needed repoint, x/c before 5/25, after 0/32 dB. |
1301 | 2022-01-25 | 15:40 | CARINA 92 | Comments | Repoint needed early spring and mid-December. |
1302 | 2022-01-25 | 16:00 | CARINA 92 | VSAT Start | Ottawa Hub, Telesat antenna up and ok. |
1303 | 2022-01-25 | 16:10 | PTCO | Service Visit | To restore power and check VSAT. Ice cleanded off antenna and solar panels. |
1304 | 2022-01-25 | 16:20 | PTCO | VSAT Check | Co/x at 28/0 dB, no adjustment needed. |
1305 | 2022-01-25 | 16:30 | PTCO | Power Check | Loose wire fixed in charge controller. |
1306 | 2022-01-25 | 19:03 | PTCO | Restart | Station is back up. Signal looks good. |
1307 | 2022-01-31 | 17:30 | OTTAWA HUB | Reboot | Naqs2 was rebooted after service maintenance. |
1308 | 2022-02-01 | 18:31 | OTTAWA HUB | Reboot | Naqs1 war rebooted after service maintenance. |
1309 | 2022-02-01 | 19:12 | OTTAWA HUB | Reboot | NaqsServer 1.9 and required DLL installed on Naqs1. Restarted at 19:12 UT. |
1310 | 2022-03-16 | 14:30 | TYNO | Service Visit | To improve VSAT. |
1311 | 2022-03-16 | 14:45 | TYNO | VSAT Check | Co/X improved to 28/0 dB. Power level up by 5 dB. |
1312 | 2022-03-16 | 15:00 | TYNO | Cold-Boot | System restarted. |
No | Date | Time UT | Location | Activity | Comments |
1313 | 2022-03-16 | 17:30 | WLVO | Service Visit | To improve VSAT. |
1314 | 2022-03-16 | 17:45 | WLVO | VSAT Check | Co/X improved to 27/0 dB. Power level up by 6 dB. |
1315 | 2022-03-16 | 18:00 | WLVO | Cold-Boot | Station restarted. |
1316 | 2022-03-20 | 22:16 | MH2B | Closed | End of Operation. |
1317 | 2022-03-22 | 23:48 | MH1B | Closed | End of Operation. |
1318 | 2022-03-29 | 15:45 | OTTAWA HUB | Comments | DataServer Stopped on Naqs2. |
1319 | 2022-03-29 | 16:16 | OTTAWA | Comments | DataServer Version 1.05.01 Running on Naqs1, port 29001. |
1320 | 2022-03-29 | 16:17 | OTTAWA | Comments | NaqsServer 1.90 continues to run. |
1321 | 2022-03-29 | 16:18 | OTTAWA | Comments | NptoNMXP Version 1.00.08 is running. Converting Taurus data to NaqsStreams. |
1322 | 2022-03-29 | 16:19 | OTTAWA | Comments | Data from Taurus 204 in name 204T are streamed to Naqs1. |
1323 | 2022-04-08 | 16:28 | OTTAWA HUB | Re-Point | Improved Co-pol by 3 and X-Pol by 2 dB for Telesat antenna. Isolastion is now 35 dB. |
1324 | 2022-04-08 | 16:58 | OTTAWA | Restart | Re-point of Telesat antenna completed, Carina 92 restarted. |
1325 | 2022-04-12 | 15:00 | OTTAWA | Re-Point | SES-AM antenna re-pointed. Improved Co/X from 27/0 to 30/0 dB. |
1326 | 2022-04-15 | 17:48 | OTTAWA | Comments | Multi-cast IP on Ottawa Hub6, Carina 63 was incorrect at 224.1.1.1, corrected to 225.1.1.1 |
1327 | 2022-04-19 | 00:30 | STCO | Require GPS Lock | New Instrument installed. |
1328 | 2022-05-04 | 18:00 | OTTAWA | Restart | Test vault with Taurus 204 restarted in UBC-type vault. |
No | Date | Time UT | Location | Activity | Comments |
1329 | 2022-05-04 | 18:10 | OTTAWA | Seismo Check | ESP s/n T3F12 is installed in Taurus vault. |
1330 | 2022-05-05 | 11:40 | OTTAWA | Comments | Taurus 204 with ESP running ok. Mass-centered. |
1331 | 2022-05-05 | 13:00 | SILO | VSAT Check | IFL loss faked from 2 to 5 dB. Eirp increase is at 9 dB in attempt to reduce ReTX. |
1332 | 2022-05-10 | 19:11 | TWGB | Closed | End of Operation. |
1333 | 2022-05-12 | 23:30 | CLVB | Service Visit | Station was down. |
1334 | 2022-05-12 | 23:31 | CLVB | Wildlife | Cables were chewed by bear. All cables including power, bus, RF, etc were repaired or replaced. |
1335 | 2022-05-12 | 23:45 | CLVB | VSAT Check | X/Co pol 2/34 dB. |
1336 | 2022-05-12 | 23:47 | CLVB | Restart | After cold boot. |
1337 | 2021-05-19 | 04:46 | CBRQ | Closed | Station closed. |
1338 | 2021-05-19 | 15:00 | OTTAWA HUB | Re-Point | SES-AM antenna re-pointed. Improved Co/X from 28/10 to 30/0 dB. |
1339 | 2021-05-24 | 15:00 | MRHQ | Service Visit | To replace bad seismometer. |
1340 | 2021-05-24 | 15:10 | MRHQ | Seismo Swap | Seismometer was flooded. Replaced with ESP T3K95. |
1341 | 2021-05-24 | 15:15 | MRHQ | Power Check | Braker was incorrectly installed. Corrected. |
1342 | 2021-05-24 | 15:20 | MRHQ | VSAT Check | Antenna gain was set incorrectly to 45.5, was corrected to 46.5 dB. |
1343 | 2021-05-24 | 15:25 | MRHQ | VSAT Check | SSPB profile corrected. |
1344 | 2021-05-24 | 15:30 | MRHQ | Re-Point | Before 32/8, now 35/0 dB Co/X-pol. |
No | Date | Time UT | Location | Activity | Comments |
1345 | 2021-05-25 | 15:00 | LONDON HUB | Re-Point | Co/X improved by 2 dB to 30/0 dB. |
1346 | 2021-05-30 | 17:37 | OTTAWA HUB | Comments | NaqsServer 1.90 is installed and started on Naqs2. |
1347 | 2021-06-07 | 16:00 | LDIO | Service Visit | To replace seismo and repoint. |
1348 | 2021-06-07 | 16:30 | LDIO | Re-Point | X/co 0/28 dB, 2 dB improvement. |
1349 | 2021-06-07 | 16:45 | LDIO | Seismo Swap | 3T seismometer removed due to LP noise, 3ESP s/n T3K42 installed. |
1350 | 2021-06-07 | 16:57 | LDIO | Seismo Check | Mass centred and calibrated ok. |
1351 | 2021-06-17 | 17:00 | PKRO | Re-Point | X/Co 30/0, 6 dB improved by clearing weeds and re-pointing. |
1352 | 2021-06-17 | 22:00 | TYNO | VSAT Check | Co/x was good at 32/0 dB. Station need sspb temperature sensor. |
1353 | 2021-06-17 | 23:45 | STCO | VSAT Check | SSPB mounting screw is damaged, cound not adjust SSPB. X/Co 35/2, 2 dB improvement. |
1354 | 2021-06-17 | 23:50 | STCO | Service Visit | New SSPB A0304A, LNB 16201509. |
1355 | 2021-06-18 | 00:01 | IGLN | Restart | Station is back after down during winter months due to snow blcking the VSAT antenna. |
1356 | 2021-06-20 | 12:45 | OTTAWA HUB | Reboot | Naqs2 needed a re-boot to clear some data TX problems. |
1357 | 2021-06-20 | 14:56 | VSAT | Comments | New GE Test Frequencies: Tx 14031.0 MHz, Rx 11731.0 MHz reported by KA. |
1358 | 2021-06-20 | 16:18 | IGLN | Seismo Check | BHE component off-center. Signal on all channels look odd. |
1359 | 2021-06-23 | 16:00 | O51T | New Station Start | At Wemindji. Recording in O51T ringbuffers at Ottawa. |
1360 | 2021-06-23 | 16:25 | WEMQ | New Station Code | Started on Naqs2 at Ottawa around 16:25 and on Naqs1 around 16:31. |
No | Date | Time UT | Location | Activity | Comments |
1361 | 2021-06-23 | 16:30 | WEMQ | Hardware List | Cyg601, trd 833 ts 737, ESP, S/N T3P32, Sspb a0017a. |
1362 | 2021-06-23 | 16:31 | WEMQ | VSAT Check | 32 dB isolation but very difficult to start. Had to increase Eirp to +10 dB, fake sspb gain by 3 dB and IFL Loss to 5 dB. |
1363 | 2021-06-23 | 16:32 | WEMQ | Gain Check | SPS 100, Gain 1.0, Vpp 40. |
1364 | 2021-06-23 | 16:33 | WEMQ | DC Filter Check | No DC filter. |
1365 | 2021-06-23 | 16:35 | WEMQ | Site Geology | Solid bedrock. |
1366 | 2021-06-26 | 19:51 | O50T | New Station Start | Recording at YOSQ in O50T ringbuffers. |
1367 | 2021-06-26 | 20:47 | YOSQ | New Station Code | Started on Naqs2 at Ottawa around 20:47 and on Naqs1 around 20:52 UT. |
1368 | 2021-06-26 | 20:50 | YOSQ | Hardware List | Cyg-599, trd-832, ts 834, esp s/n t3n25, sspb a0018a 2 watt. |
1369 | 2021-06-26 | 20:51 | YOSQ | Gain Check | SPS 100, Gain 1.0, Vpp 40. |
1370 | 2021-06-26 | 20:52 | YOSQ | DC Filter Check | No DC filter. |
1371 | 2021-06-26 | 20:53 | YOSQ | VSAT Check | Antenna was not aligned correctly. Had to re-define sweet spot. |
1372 | 2021-06-26 | 20:55 | YOSQ | Site Geology | Seismometer vault in sand over solid bedrock |
1373 | 2021-06-27 | 17:35 | WEMQ | Service Visit | To check VSAT antenna. No problems noted. |
1374 | 2021-07-15 | 17:41 | CAMN | Closed | End of operation. |
1375 | 2021-07-16 | 19:30 | KNDN | Service Visit | To restore dead station. |
1376 | 2021-07-16 | 19:40 | KNDN | Wildlife | Damage to power cable was fixed. |
No | Date | Time UT | Location | Activity | Comments |
1377 | 2021-07-16 | 19:50 | KNDN | Restart | Station restarted |
1378 | 2021-07-17 | 15:51 | NODN | Closed | End of Operation. |
1379 | 2021-07-18 | 11:30 | FEDNOR | Comments | Started field work for installation of FedNor 3 stations. |
1380 | 2021-07-18 | 15:42 | IHLN | Closed | End of operation. |
1381 | 2021-07-19 | 20:59 | SNPN | Service Visit | MT electordes changed. |
1382 | 2021-07-19 | 20:59 | SNPN | VSAT Check | Repoint from 15/0 to 30/1 dB. |
1383 | 2021-07-19 | 21:00 | SNPN | Comments | GPS antenna connector was loose. Fixed and antenna mast installed. |
1384 | 2021-07-19 | 21:01 | SNPN | Comments | Large blase planned at 6 local time at Snap Lake. |
1385 | 2021-07-19 | 21:01 | SNPN | Comments | Staion looked good, couple of loose bolts fixed. |
1386 | 2021-07-20 | 14:40 | MGTN | Service Visit | To fix MT equipment. |
1387 | 2021-07-20 | 14:41 | MGTN | Wildlife | Sever damage, beyone repair, to MT equipment. |
1388 | 2021-07-20 | 14:45 | MGTN | Comments | Seismic is ok, MT can not be repair. Closed. |
1389 | 2021-07-20 | 17:50 | O64T | New Station Start | New station started near Cobalt. |
1390 | 2021-07-20 | 17:51 | O64T | Hardware List | Cyg 561, Trident 754, sspb A0014a, esp T3P33. |
1391 | 2021-07-20 | 17:52 | O64T | Comments | At 47.37082 N, 79.66566 W, 306 m. |
1392 | 2021-07-20 | 17:52 | O64T | VSAT Check | Co/x 32/2 dB. |
No | Date | Time UT | Location | Activity | Comments |
1393 | 2021-07-20 | 17:54 | O64T | Gain Check | 1.0 at 100 sps. |
1394 | 2021-07-20 | 17:54 | O64T | Site Geology | Solid bedrock. |
1395 | 2021-07-20 | 17:55 | O64T | Comments | Require GPS Lock is on. |
1396 | 2021-07-20 | 17:55 | O64T | DC Filter Check | On at 0.001 Hz. |
1397 | 2021-07-20 | 17:56 | O64T | Cold-Boot | To restart and capture screens. |
1398 | 2021-07-20 | 18:00 | MCKN | Service Visit | To restore seismic and service MT. |
1399 | 2021-07-20 | 18:01 | MCKN | Hardware Swap | Cygnus 550, Trdident 245 installed on Hub 6. |
1400 | 2021-07-20 | 18:20 | MCKN | Comments | Servicing MT equipment. |
1401 | 2021-07-20 | 18:50 | HSMO | New Station Code | New Station started on both Ottawa Hub NaqsServer. Switched from O64T to HSMO. |
1402 | 2021-07-20 | 19:28 | HSMO | SSPB Check | SSPB profile corrected. All modem parameters checked, were ok. Screen captures done. Leaving the site. |
1403 | 2021-07-20 | 20:15 | MCKN | Gain Check | Confirm gain is at 1.0. |
1404 | 2021-07-20 | 20:16 | MCKN | Comments | Finished MT, replace electords and checked Mag orientation and replace gps antenna mast. GPS was locked on MT. |
1405 | 2021-07-20 | 20:16 | MCKN | Comments | Require GPS Lock is on. |
1406 | 2021-07-20 | 20:16 | MCKN | DC Filter Check | Is not one. |
1407 | 2021-07-20 | 20:17 | MCKN | Comments | Team leaving site for Snap Lake. Due to leave for YK tomorrow. |
1408 | 2021-07-23 | 14:55 | O53T | New Station Code | Starting new station at Ottawa Hub, Moosonee location. |
No | Date | Time UT | Location | Activity | Comments |
1409 | 2021-07-23 | 14:56 | O53T | Hardware List | Cyg 606, trd 851, seismo 3esp t3p31, sspb a0016a |
1410 | 2021-07-23 | 14:58 | O53T | Site Geology | Soft sand, about 100 m from Airport runway. |
1411 | 2021-07-23 | 14:58 | O53T | VSAT Check | 30/0 dB. |
1412 | 2021-07-23 | 15:17 | MSNO | New Station Code | Started at Ottawa Hub Naqs2, Moosonee. |
1413 | 2021-07-23 | 15:23 | MSNO | New Station Start | Started at Ottawa Hub Naqs1, Moosonee. |
1414 | 2021-07-23 | 15:26 | MSNO | Gain Check | SPS 100, Gain 1. |
1415 | 2021-07-23 | 15:27 | MSNO | Comments | Require GPS lock to lock is on. |
1416 | 2021-07-23 | 15:27 | MSNO | DC Filter Check | DC filer on at 1 mHz. |
1417 | 2021-07-23 | 15:28 | MSNO | Comments | Finished installation, team to install Timmins Sunday. |
1418 | 2021-07-23 | 16:05 | DSMN | Closed | End of operation. |
1419 | 2021-07-24 | 15:30 | GBLN | Service Visit | Seismo vault tipped on its side. MT equipment damaged but may be fixable. Repair in progress. |
1420 | 2021-07-24 | 15:40 | GBLN | Hardware Swap | TRD 143 removed, may be damagd. |
1421 | 2021-07-24 | 15:40 | GBLN | Hardware Swap | Trd 513 installed. |
1422 | 2021-07-24 | 16:00 | GBLN | Seismo Check | Three attempts failed to mass-center seismometer. May be damaged. |
1423 | 2021-07-24 | 16:30 | GBLN | Seismo Swap | Bad (damaged?) seismometer T3K45 replaced. ESP T3J66 installed. |
1424 | 2021-07-24 | 16:55 | GBLN | Seismo Check | New seismometer responded ok to mass-center and calibration. |
No | Date | Time UT | Location | Activity | Comments |
1425 | 2021-07-24 | 16:57 | GBLN | Comments | MT repair in progress. |
1426 | 2021-07-25 | 18:55 | O62T | New Station Start | Started Timmins stations in O62T at Ottawa Hub. |
1427 | 2021-07-25 | 18:56 | O62T | Hardware List | Cyg 610, Trd 227, ESP T3F12, SSPB A0013A |
1428 | 2021-07-25 | 19:05 | O62T | Seismo Check | Bedrock at 48.46587 N 81.30321 W, 392 m |
1429 | 2021-07-25 | 19:06 | O62T | Gain Check | 100 SPS at 1.0. |
1430 | 2021-07-25 | 19:06 | O62T | VSAT Check | 30/0 dB. IP 79. |
1431 | 2021-07-25 | 19:07 | O62T | Comments | Require GPS Lock is on. |
1432 | 2021-07-25 | 19:07 | O62T | DC Filter Check | On at 0.001 Hz. |
1433 | 2021-07-25 | 19:08 | O62T | Seismo Check | Mass centred and calibrated ok. |
1434 | 2021-07-25 | 19:26 | TIMO | New Station Code | O62T is now started in TIMO at Ottawa Hub Naqs 2. |
1435 | 2021-07-25 | 19:28 | TIMO | New Station Code | Started at Ottawa Hub Naqs1 in name TIMO. |
1436 | 2021-07-25 | 19:29 | NAQS1 | Comments | Dataserver restarted on Ottawa Hub Naqs1. |
1437 | 2021-07-26 | 18:25 | HLSB | Closed | End of operation. |
1438 | 2021-07-27 | 01:40 | TIMO | Cables | Problem with RF out at SSPB end fixed. Station restarted. |
1439 | 2021-07-27 | 15:25 | HRMB | Closed | End of operation. |
1440 | 2021-07-28 | 01:38 | PIMB | Closed | End of operation. |
No | Date | Time UT | Location | Activity | Comments |
1441 | 2021-07-28 | 14:58 | HBDB | Closed | End of operation. |
1442 | 2021-07-29 | 11:53 | O47T | New Station Start | Started at Neskantaga First Nation Airport. |
1443 | 2021-07-29 | 12:44 | O47T | Hardware List | Cyg 268, trd 396, sspb a0015a, esp seismo t3p34 |
1444 | 2021-07-29 | 13:16 | O47T | Gain Check | Sps 100, gain 1.0 |
1445 | 2021-07-29 | 13:16 | O47T | VSAT Check | 22/2 dB. |
1446 | 2021-07-29 | 13:17 | O47T | Comments | Require GPS Lock is on. |
1447 | 2021-07-29 | 13:17 | O47T | Gain Check | On at 0.001 Hz. |
1448 | 2021-07-29 | 13:17 | O47T | Site Geology | Soft sandy soil. |
1449 | 2021-07-29 | 13:18 | O47T | Comments | Location 52.19650 N 87.93053 W 241 m |
1450 | 2021-07-29 | 13:22 | O47T | Comments | Site is 100 m from runway at small airport with about 5 daily flights. |
1451 | 2021-07-29 | 13:33 | NSKO | New Station Code | Started at Ottawa Hub Naqs 2. |
1452 | 2021-07-29 | 13:36 | NSKO | New Station Code | Started at Ottawa Hub Naqs 1. |
1453 | 2021-07-29 | 13:37 | OTTAWA HUB | Comments | Dataserver restarted on Ottawa Hub Naqs1. |
1454 | 2021-08-02 | 18:28 | O54T | New Station Start | Started at kasabonika Lake, KASO recording at Ottawa Hub. |
1455 | 2021-08-02 | 18:30 | O54T | Hardware List | Cyg 607, Trd 852, TS 748, sspb a0019a, esp seismo T3P61, IP 78 |
1456 | 2021-08-02 | 18:40 | O54T | VSAT Check | 35/5 dB. |
No | Date | Time UT | Location | Activity | Comments |
1457 | 2021-08-02 | 18:50 | O54T | Comments | Require GPS Lock is on. |
1458 | 2021-08-02 | 18:50 | O54T | DC Filter Check | On at 0.001 Hz. |
1459 | 2021-08-02 | 18:50 | O54T | Gain Check | SPS 100, Gain 1.0, Vpp 40. |
1460 | 2021-08-02 | 18:55 | O54T | Comments | About 80 m from runway. |
1461 | 2021-08-02 | 18:55 | O54T | Comments | Site is at small airport with 5-7 flights a day. Expect high wind and snow. |
1462 | 2021-08-02 | 18:55 | O54T | Site Geology | Seismometer in sand, at 53.52785N 88.64139W 192 m. |
1463 | 2021-08-03 | 15:03 | COQB | Closed | End of operation. |
1464 | 2021-08-03 | 15:17 | KASO | New Station Code | O54T is now recording in KASO ringbuffers at Ottawa Hub Naqs2. |
1465 | 2021-08-03 | 15:21 | KASO | New Station Code | O54T is now recording in KASO ringbuffers at Ottawa Hub Naqs1. |
1466 | 2021-08-03 | 15:22 | OTTAWA HUB | Comments | Dataserver restarted on Ottawa Hub Naqs1. |
1467 | 2021-08-04 | 15:19 | CLVB | Closed | End of operation. |
1468 | 2021-08-04 | 19:13 | O62T | New Station Code | New station started at Aroland Nakina, recording at Ottawa Hub. |
1469 | 2021-08-04 | 19:27 | O62T | Hardware List | Cyg 395, TRD 853, TS 480, sspb A0403A, ESP seismo T3K12. |
1470 | 2021-08-04 | 19:28 | O62T | Gain Check | SPS 100, Gain 1.0, Vpp 40. |
1471 | 2021-08-04 | 19:28 | O62T | VSAT Check | 31/0 dB. |
1472 | 2021-08-04 | 19:29 | O62T | Comments | Require GPS Lock is on. |
No | Date | Time UT | Location | Activity | Comments |
1473 | 2021-08-04 | 19:29 | O62T | DC Filter Check | On at 0.001 Hz. |
1474 | 2021-08-04 | 19:29 | O62T | Site Geology | Bedrock at 50.35429 N, 86.96841 W, 309 m |
1475 | 2021-08-04 | 19:30 | O62T | Comments | Site is about 1/2 hours from the Band. |
1476 | 2021-08-04 | 19:44 | NANO | New Station Code | Started recording with code NANO at Aroland/Nakina in Ottawa Hub Naqs2. |
1477 | 2021-08-04 | 19:48 | NANO | New Station Code | Started recording with code NANO at Aroland/Nakina in Ottawa Hub Naqs1. |
1478 | 2021-08-04 | 19:49 | OTTAWA HUB | Comments | Dataserver restarted on Ottawa Hub Naqs1. |
1479 | 2021-08-04 | 19:50 | NANO | Comments | End of FedNor3 installation. |
1480 | 2021-08-10 | 21:30 | RLKO | Service Visit | To fix lack of data problem. |
1481 | 2021-08-10 | 21:31 | RLKO | Hardware Swap | Cyguns 621, trd 867, seismo T3P07 installed. |
1482 | 2021-08-10 | 21:33 | RLKO | Hardware List | Cygnus 536, trd 695 removed. |
1483 | 2021-08-10 | 21:34 | RLKO | Comments | Old cygnus was not providing power on the bus. Cable was tested ok, cygnus faulty. |
1484 | 2021-08-10 | 21:35 | RLKO | VSAT Check | 30/0 dB. No adjustment needed. |
1485 | 2021-08-10 | 21:47 | RLKO | Comments | Site undisturbed. |
1486 | 2021-08-10 | 21:57 | RLKO | Comments | Finished service. New ESP seismo mass-centered ok. |
1487 | 2021-08-10 | 21:58 | OTTAWA HUB | Comments | Naqs1, 2 and Dataserver restarted. |
1488 | 2021-08-10 | 22:51 | RLKO | Gain Check | 100 SPS, Gain 0.5 |
No | Date | Time UT | Location | Activity | Comments |
1489 | 2021-08-10 | 22:52 | RLKO | Comments | Require GPS lock to lock is on. |
1490 | 2021-08-10 | 22:52 | RLKO | DC Filter Check | DC filer on at 1 mHz. |
1491 | 2021-08-11 | 14:55 | RLKO | Seismo Swap | Bad HHE component with new T3P07 seismo, replaced with original T3762, 3T seismo. |
1492 | 2021-08-11 | 15:02 | RLKO | Comments | End of Service visit. Site is back to normal. |
1493 | 2021-08-11 | 18:28 | MH4B | New Station Start | Started at MSH site close to volcano |
1494 | 2021-08-11 | 18:29 | MH4B | Hardware List | Cyg 619, Trd 064, TRD 137, sspb a0406a, seismometer 40T, T4777 |
1495 | 2021-08-11 | 18:31 | MH4B | Comments | Seismometer is oriented to North. |
1496 | 2021-08-11 | 18:32 | MH4B | Gain Check | TRD 064, SPS 40, Gain 0.5 |
1497 | 2021-08-11 | 18:32 | MH4B | VSAT Check | 36/5 dB. |
1498 | 2021-08-11 | 18:36 | MH4B | DC Filter Check | TRd 064, DC filter in not on. |
1499 | 2021-08-11 | 18:37 | MH4B | DC Filter Check | Trd 137, DC is not on. |
1500 | 2021-08-11 | 18:37 | MH4B | Gain Check | Trd 137, sps 40, Gain 0.5 |
1501 | 2021-08-11 | 18:38 | MH4B | Require GPS Lock | Require GPS Lock is on on both Tridents. |
1502 | 2021-08-11 | 18:38 | MH4B | Site Geology | Volcanic ash. Seismo on cement block. |
1503 | 2021-08-11 | 18:40 | MH4B | Comments | Team is taking GPS locations of all sensors. |
1504 | 2021-08-11 | 19:20 | MH4B | Comments | GPS location of seismo 46.30791 N 122.26601 W 1051 m |
No | Date | Time UT | Location | Activity | Comments |
1505 | 2021-08-11 | 19:21 | MH4B | New Station Code | Started at both Naqs1 and Naqs2, Ottawa Hub at18:27. |
1506 | 2021-08-13 | 23:14 | MH3B | New Station Start | Started recording at Ottawa Hub Naqs1 and 2. |
1507 | 2021-08-13 | 23:15 | MH3B | Hardware List | Cyg 616, Trd 854, Trd 856, sspb a0401a, 40T T4776. |
1508 | 2021-08-13 | 23:15 | MH3B | VSAT Check | 40/3 dB. |
1509 | 2021-08-13 | 23:16 | MH3B | Comments | Seismometer is oriented to True North. |
1510 | 2021-08-13 | 23:17 | MH3B | Gain Check | Trd 854, Gain 0.5, sps 40. |
1511 | 2021-08-13 | 23:18 | MH3B | DC Filter Check | Unticked on Trd 854. |
1512 | 2021-08-13 | 23:18 | MH3B | Gain Check | Trd 856, Gain 0.5, sps 40. |
1513 | 2021-08-13 | 23:19 | MH3B | DC Filter Check | Trd 856, unticked. |
1514 | 2021-08-13 | 23:20 | MH3B | Require GPS Lock | Require GPS lock to lock is on on both Tridents. |
1515 | 2021-08-13 | 23:20 | MH3B | Site Geology | Riverbed silt, seismo on concrete slab. |
1516 | 2021-08-13 | 23:21 | MH3B | Comments | Golf course site. |
1517 | 2021-08-13 | 23:21 | MH3B | Comments | Seismo is at 46.30846 N 119.33798 W 117 m |
1518 | 2021-08-14 | 00:12 | MH3B | Comments | Wind sensor data are received at the Hub after a power cycle of the WS425 Wind Sensor. Team about to leave the site. |
1519 | 2021-08-17 | 22:06 | AP3N | New Station Start | Started new Hudson Bay station at Apex 3 site, recording on Naqs2, Ottawa Hub. |
1520 | 2021-08-17 | 22:07 | AP3N | Gain Check | Gain of 1.0, sps 40. |
No | Date | Time UT | Location | Activity | Comments |
1521 | 2021-08-17 | 22:07 | AP3N | Hardware List | Cyg 253, Trd 158, sspb a001a, 2w, seismo ESP T3B50. |
1522 | 2021-08-17 | 22:08 | AP3N | VSAT Check | Telesat, first site on new bandwidth, 35 dB separation. |
1523 | 2021-08-17 | 22:09 | AP3N | DC Filter Check | DC filter is not on. |
1524 | 2021-08-17 | 22:10 | AP3N | Require GPS Lock | Require GPS Lock is on. |
1525 | 2021-08-17 | 22:11 | AP3N | Site Geology | Solid bedrock at 69 27 31.1 N 84 27 41.7 W 437 m. |
1526 | 2021-08-17 | 22:20 | AP3N | Comments | Started on Naqs1 at Ottawa Hub. Dataserver restarted. |
1527 | 2021-08-18 | 17:56 | LAIN | New Station Start | HB, Lailor River, recording on Naqs1 and Naqs2, Ottawa Hub. |
1528 | 2021-08-18 | 17:59 | LAIN | DC Filter Check | DC filter is not on. |
1529 | 2021-08-18 | 17:59 | LAIN | Gain Check | Gain of 1.0, sps 40. |
1530 | 2021-08-18 | 17:59 | LAIN | Hardware List | Cyg 261, Trd 248, sspb A0011A, 2 W, seismo, ESP, T3b27. |
1531 | 2021-08-18 | 18:00 | LAIN | Comments | Com port is not working on Cyg 261 but is not needed here. |
1532 | 2021-08-18 | 18:01 | LAIN | Comments | Mass centred and calibrated ok. |
1533 | 2021-08-18 | 18:01 | LAIN | Require GPS Lock | Require GPS Lock is on. |
1534 | 2021-08-18 | 18:01 | LAIN | VSAT Check | On telesat, 30 dB of isolation. |
1535 | 2021-08-18 | 18:02 | LAIN | Comments | Team are finishing off seismometer vault. |
1536 | 2021-08-18 | 18:02 | LAIN | Site Geology | Bedrock site. |
No | Date | Time UT | Location | Activity | Comments |
1537 | 2021-08-19 | 17:03 | GOWB | Closed | End of operation. |
1538 | 2021-08-19 | 21:48 | GIFN | New Station Start | Gifford Fjord, Baffin Island, recording at Naqs2, Ottawa. |
1539 | 2021-08-19 | 21:52 | GIFN | New Station Start | Gifford Fjord, Baffin Island, recording at Naqs1, Ottawa. |
1540 | 2021-08-19 | 21:53 | GIFN | Hardware List | Cyg 609 Trd 848 sspb a0002a, Seismometer esp T3p20. |
1541 | 2021-08-19 | 21:54 | GIFN | Gain Check | Gain of 1.0, sps 40. |
1542 | 2021-08-19 | 21:55 | GIFN | DC Filter Check | DC filter is not on. |
1543 | 2021-08-19 | 21:56 | GIFN | VSAT Check | 30 dB of isolation on Telesat. |
1544 | 2021-08-19 | 21:57 | GIFN | Site Geology | Sand. |
1545 | 2021-08-21 | 21:09 | SRLN | New Station Start | At Sarcpa Lake old Dewline station, recording on Naqs2, Ottawa. |
1546 | 2021-08-21 | 21:12 | SRLN | New Station Start | At Sarcpa Lake old Dewline station, recording on Naqs1, Ottawa. |
1547 | 2021-08-21 | 21:14 | SRLN | Hardware List | Cyg244, Trd136, Seismometer Esp T3B37(confirm with pics). |
1548 | 2021-08-21 | 21:15 | SRLN | Gain Check | Gain of 1.0, sps 40. |
1549 | 2021-08-21 | 21:16 | SRLN | DC Filter Check | DC filter is not on. |
1550 | 2021-08-21 | 21:17 | SRLN | Require GPS Lock | Require GPS Lock is on. |
1551 | 2021-08-21 | 21:18 | SRLN | Site Geology | Solid bedrock. |
1552 | 2021-08-21 | 21:19 | SRLN | Comments | Antenna and parts of solar panel frames bolted to rock. |
No | Date | Time UT | Location | Activity | Comments |
1553 | 2021-08-22 | 14:56 | GIFN | Require GPS Lock | Require GPS Lock is on. |
1554 | 2021-08-22 | 14:57 | GIFN | Seismo Check | New ESP type with auto mass lock. |
1555 | 2021-08-22 | 20:16 | IGLN | Closed | Moved, due to flooding of seismometer vault, to ILON. |
1556 | 2021-08-23 | 21:14 | ILON | Comments | Old code of IGLN is retired due to change of location. |
1557 | 2021-08-23 | 21:17 | ILON | Hardware List | CYG620, TRD871, ESP Seismo T3P18 (auto-mass-lock) |
1558 | 2021-08-23 | 21:18 | ILON | DC Filter Check | DC filter is not on. |
1559 | 2021-08-23 | 21:18 | ILON | Gain Check | Gain of 1.0, sps 40. |
1560 | 2021-08-23 | 21:18 | ILON | Site Geology | Solid bedrock a couple of feet below seismometer pier. |
1561 | 2021-08-23 | 21:19 | ILON | Seismo Check | Waiting for pier to dry out, seismometer not connected yet. |
1562 | 2021-08-23 | 21:20 | ILON | Require GPS Lock | Require GPS Lock is on. |
1563 | 2021-08-23 | 21:20 | ILON | VSAT Check | 30 dB of isolation on Telesat. |
1564 | 2021-08-23 | 23:00 | ILON | New Station Start | AT new location, Igloolik Airport Weather Station, recording at Ottawa Hub Naqs2 and Naqs1. |
1565 | 2021-08-24 | 13:30 | ILON | Seismo Swap | T3P18 has bad Z component, removed, T3K44 installed. |
1566 | 2021-08-24 | 13:49 | ILON | Comments | End of new installation. |
1567 | 2021-08-26 | 18:30 | TWBB | Service Visit | To replace broken Cygnus 177. |
1568 | 2021-08-26 | 18:35 | TWBB | Hardware Swap | Cyg177 removed, Cyg285 installed. |
No | Date | Time UT | Location | Activity | Comments |
1569 | 2021-08-26 | 18:40 | TWBB | VSAT Check | Co/x 18/0, low. |
1570 | 2021-08-26 | 18:42 | TWBB | Comments | Site is back to normal. |
1571 | 2021-08-30 | 17:14 | ENGB | Closed | End of operation. |
1572 | 2021-09-03 | 21:11 | PLCB | Comments | Renamed to PLBC, later. |
1573 | 2021-09-03 | 21:11 | PLCB | DC Filter Check | Off. |
1574 | 2021-09-03 | 21:11 | PLCB | Gain Check | 1.0 at 100 sps. |
1575 | 2021-09-03 | 21:11 | PLCB | Hardware List | Cyg 277, Trd, 276, esp s/n t3n05 |
1576 | 2021-09-03 | 21:11 | PLCB | Require GPS Lock | On. |
1577 | 2021-09-03 | 21:11 | PLCB | VSAT Check | 30 db Isolation on Telesat. |
1578 | 2021-09-16 | 21:55 | PLCB | Closed | Changed name to PLBC. |
1579 | 2021-09-16 | 21:57 | PLBC | New Station Code | PLCB is renamed to PLBC, Ottawa, Naqs2 data starts. |
1580 | 2021-09-16 | 22:05 | PLBC | Comments | Ottawa Hub, Dataserver restarted. |
1581 | 2021-09-16 | 22:05 | PLBC | Name Change | PLCB is renamed to PLBC, Ottawa, Naqs1 data starts. |
1582 | 2021-09-29 | 02:15 | TIGG | New Station Start | Tignish, PEI, first of re-deployment. |
1583 | 2021-10-01 | 19:18 | MADG | New Station Start | At Ile d'Entree, near Magdelan Is., NS. |
1584 | 2021-10-02 | 23:31 | TIGG | Recording Hub | London. |
No | Date | Time UT | Location | Activity | Comments |
1585 | 2021-10-02 | 23:32 | TIGG | Gain Check | Gain 1.0, sps 100. |
1586 | 2021-10-02 | 23:33 | TIGG | DC Filter Check | DC filter is off. |
1587 | 2021-10-02 | 23:33 | TIGG | Require GPS Lock | On. |
1588 | 2021-10-02 | 23:34 | TIGG | Hardware List | Cyg172, TRD78, sspb h0018a, esp s/n t3d95, IP 172. |
1589 | 2021-10-02 | 23:35 | TIGG | VSAT Check | 0/16 dB. Poor. |
1590 | 2021-10-02 | 23:37 | MADG | Recording Hub | London. |
1591 | 2021-10-02 | 23:38 | MADG | DC Filter Check | DC filter is off. |
1592 | 2021-10-02 | 23:38 | MADG | Gain Check | Gain of 1.0 at 100 sps. |
1593 | 2021-10-02 | 23:38 | MADG | SR Change | Incorrectly reported 100 sps, it remains at 40 sps until Oct. 21 at 12:48 UT. |
1594 | 2021-10-02 | 23:39 | MADG | Require GPS Lock | On. |
1595 | 2021-10-02 | 23:39 | MADG | VSAT Check | 2/30 dB. |
1596 | 2021-10-02 | 23:40 | MADG | Hardware List | Cyg254, TRD767, sspb h0019a, ip 254. |
1597 | 2021-10-02 | 23:41 | MADG | SSPB Check | Soil/clay. |
1598 | 2021-10-02 | 23:41 | MADG | Seismo Check | Seismometer T3A81 is damaged, needs replacement. |
1599 | 2021-10-02 | 23:42 | MADG | Comments | Antenna is mounted on King Post. Site is solar-powered. |
1600 | 2021-10-06 | 12:53 | CODG | New Station Start | Started at London Hub, station Codroy, NF. |
No | Date | Time UT | Location | Activity | Comments |
1601 | 2021-10-06 | 12:56 | CODG | Hardware List | Cyg 262, Trd 293, esp seism. T3B19. |
1602 | 2021-10-06 | 12:57 | CODG | Gain Check | Gain 1.0, SPS 100. |
1603 | 2021-10-06 | 13:02 | CODG | DC Filter Check | Off. |
1604 | 2021-10-06 | 13:02 | CODG | Require GPS Lock | On. |
1605 | 2021-10-06 | 13:03 | CODG | Comments | IP 62. We had trouble with Cygnus 283 at this location. Phase lock timing was not working. Used another cygnus, 262 instead. |
1606 | 2021-10-06 | 13:04 | CODG | VSAT Check | 0/16 dB. Poor. |
1607 | 2021-10-06 | 13:37 | CODG | SSPB Check | H0003A. Not sure if profile is correct. |
1608 | 2021-10-06 | 13:39 | CODG | Comments | 47.84056 N 59.25352 W 50 m (seismo location). |
1609 | 2021-10-06 | 13:40 | CODG | Comments | DC site, 8 solar panels, 6 batteries. |
1610 | 2021-10-06 | 13:40 | CODG | Site Geology | Clay. |
1611 | 2021-10-06 | 13:41 | CODG | Comments | Poor VSAT signal, needed high Erip increase. |
1612 | 2021-10-12 | 23:21 | OTTAWA | Comments | New (third) 3.8 M Hub antenna installed. Served by SatMex. |
1613 | 2021-10-13 | 19:56 | SABG | New Station Start | Started recording from Sable Island at London Hub. |
1614 | 2021-10-15 | 14:30 | MALG | New Station Start | Started station Malagash, NS recording at London Hub. |
1615 | 2021-10-17 | 12:33 | GBN | New Station Start | Started new Atlantic station at Guysborough, recording at London Hub. |
1616 | 2021-10-17 | 12:34 | GBN | Comments | Dataserver restarted on London Hub, Naqs1. |
No | Date | Time UT | Location | Activity | Comments |
1617 | 2021-10-18 | 19:29 | GBN | Hardware List | Cygnus 535, Trident 214, Carina 29, IP 39, Seismo ESP, s/n T3D80, sspb H0002A. |
1618 | 2021-10-18 | 19:33 | GBN | DC Filter Check | Off. |
1619 | 2021-10-18 | 19:33 | GBN | Gain Check | Gain 1.0, SPS 100. |
1620 | 2021-10-18 | 19:33 | GBN | Require GPS Lock | On. |
1621 | 2021-10-18 | 19:34 | GBN | Site Geology | Solid bedrock. |
1622 | 2021-10-18 | 19:34 | GBN | VSAT Check | 26/0 dB. |
1623 | 2021-10-18 | 19:35 | GBN | Comments | This is AC site with NPM mount bolted to bedrock. |
1624 | 2021-10-19 | 15:32 | CHEG | New Station Start | Ottawa Hub, Naqs2, Telesat, Cheticamp, NS |
1625 | 2021-10-19 | 15:34 | CHEG | Hardware List | Cygnus 562, Trd 760, sspb H0014a 1w, esp seismo. |
1626 | 2021-10-19 | 15:38 | CHEG | DC Filter Check | Off. |
1627 | 2021-10-19 | 15:38 | CHEG | Gain Check | 1.0 at 100 sps. |
1628 | 2021-10-19 | 15:38 | CHEG | Require GPS Lock | On. |
1629 | 2021-10-19 | 15:38 | CHEG | VSAT Check | 37 dB isolation. |
1630 | 2021-10-19 | 15:39 | CHEG | Site Geology | Seis on patch of bedrock. |
1631 | 2021-10-19 | 15:40 | CHEG | Comments | VSAT on mount bolted to bedrock. |
1632 | 2021-10-19 | 15:40 | CHEG | Seismo Check | Seismo at 46.80836 N 60.67447 W 1463 ft |
No | Date | Time UT | Location | Activity | Comments |
1633 | 2021-10-19 | 15:42 | CHEG | Comments | Mass centred and calibrated ok. |
1634 | 2021-10-19 | 15:50 | CHEG | Comments | Started recording at Ottawa Naqs1. |
1635 | 2021-10-19 | 15:51 | CHEG | Comments | Dataserver restarted on London Hub, Naqs1. |
1636 | 2021-10-19 | 15:51 | CHEG | Comments | Used Telesat here since GE had poor coverage. |
1637 | 2021-10-19 | 19:01 | MALG | Hardware List | Cygnus 398, Trd 418, seismo ESP, s/n T3B20. |
1638 | 2021-10-19 | 19:03 | MALG | DC Filter Check | Off. |
1639 | 2021-10-19 | 19:03 | MALG | Gain Check | Gain 1.0, SPS 100. |
1640 | 2021-10-19 | 19:03 | MALG | VSAT Check | 29/0 dB. |
1641 | 2021-10-19 | 19:04 | MALG | Require GPS Lock | On. |
1642 | 2021-10-19 | 19:12 | SABG | Gain Check | Gain 1.0, SPS 100. |
1643 | 2021-10-19 | 19:13 | SABG | Comments | Small 1.2 m antenna used. |
1644 | 2021-10-19 | 19:13 | SABG | DC Filter Check | Off. |
1645 | 2021-10-19 | 19:13 | SABG | VSAT Check | 27/0 dB. |
1646 | 2021-10-19 | 19:14 | SABG | Seismo Check | ESP, s/n TBD at 43.93124 -60.00843 0. |
1647 | 2021-10-21 | 12:48 | MADG | SR Change | From incorrect 40 sps to correct 100 sps. |
1648 | 2021-10-22 | 14:25 | MADG | Seismo Swap | Seismo with bad Z component swapped with good seismo. |
No | Date | Time UT | Location | Activity | Comments |
1649 | 2021-10-22 | 14:25 | MADG | Seismo Swap | T3A81 removed, T3B22 installed. |
1650 | 2021-10-22 | 17:43 | BATG | New Station Start | Bathurst, NB, recording at London Hub. |
1651 | 2021-10-22 | 18:08 | BATG | Hardware List | Cygnus 605, Trd 269, sspb 1w A0126A, seismo esp s/n T3D78. |
1652 | 2021-10-22 | 18:44 | BATG | DC Filter Check | Off. |
1653 | 2021-10-22 | 18:44 | BATG | Gain Check | 1.0 at 100 sps. |
1654 | 2021-10-22 | 18:44 | BATG | VSAT Check | 26/0 dB. |
1655 | 2021-10-22 | 18:45 | BATG | Power Check | AC site. |
1656 | 2021-10-22 | 18:45 | BATG | Require GPS Lock | On. |
1657 | 2021-10-22 | 18:45 | BATG | Site Geology | On edge of small dam bolted to bedrock. |
1658 | 2021-10-25 | 14:42 | GASG | New Station Start | Gaspe, Quebec, on Telesat, recording at Ottawa Hub. |
1659 | 2021-10-25 | 14:47 | GASG | Hardware List | Cyg 419, Trd 296, seismo esp T3D85, SSPB H001a 1w, TS 547. |
1660 | 2021-10-25 | 14:48 | GASG | Comments | 48.94626 N 66.11613 W 855 ft. |
1661 | 2021-10-25 | 14:49 | GASG | Site Geology | Glacial. |
1662 | 2021-10-25 | 14:50 | GASG | Comments | Movement of material in nearby gravel pit for a couple of days. |
1663 | 2021-10-25 | 14:50 | GASG | Gain Check | 1.0 at 100 sps. |
1664 | 2021-10-25 | 14:51 | GASG | DC Filter Check | Off. |
No | Date | Time UT | Location | Activity | Comments |
1665 | 2021-10-25 | 14:51 | GASG | Require GPS Lock | On. |
1666 | 2021-10-25 | 14:51 | GASG | VSAT Check | 33 dB isolation on Telesat. |
1667 | 2021-10-25 | 14:52 | GASG | Power Check | DC site on 6 batteries and 10 panels, but limited view of sun. |
1668 | 2021-10-26 | 16:42 | ALFO | Service Visit | Station has been dead, found Cygnus to be dead. |
1669 | 2021-10-26 | 16:43 | ALFO | Hardware Swap | Dead Cygnus 175 removed, Cygnus 269 installed. |
1670 | 2021-10-26 | 16:48 | ALFO | Comments | Dataserver restarted at London Hub (ALFO is served by this Hub). |
1671 | 2021-11-01 | 21:00 | RLKO | Comments | New seismometer T3P06 has electonics mass unlock. |
1672 | 2021-11-01 | 21:00 | RLKO | Seismo Swap | T3762 removed, ESP T3P06 installed. |
1673 | 2021-11-01 | 21:00 | RLKO | Service Visit | To replace seismometer with bad ns component. |
1674 | 2021-11-26 | 15:00 | NATG | Comments | Recording at Ottawa Hub, Telesat. |
1675 | 2021-11-26 | 15:00 | NATG | New Station Start | Last of the Atlantic 2005, about 40 km from Natashquan, Quebec. |
1676 | 2021-11-28 | 14:32 | NATG | DC Filter Check | Off. |
1677 | 2021-11-28 | 14:32 | NATG | Gain Check | 1.0 at 100 sps. |
1678 | 2021-11-28 | 14:32 | NATG | Require GPS Lock | On. |
1679 | 2021-11-28 | 14:33 | NATG | Hardware List | Cyg 161, Trd 913, TS 075, esp T3P19 with auto mass-lock. |
1680 | 2021-11-28 | 14:33 | NATG | Site Geology | Solid Bedrock. |
No | Date | Time UT | Location | Activity | Comments |
1681 | 2021-11-28 | 14:33 | NATG | VSAT Check | 34 dB isolation, Telesat. |
1682 | 2021-11-28 | 14:34 | NATG | Comments | King Post bolted to rock face. Need more bolts on next visit. |
1683 | 2021-11-28 | 14:34 | NATG | Power Check | AC site. No grounding in supplied power. |
1684 | 2021-11-28 | 14:35 | NATG | Comments | Cygnus and charging system in battery box with two AC-type batteries. |
1685 | 2021-12-07 | 19:00 | DELO | Service Visit | To upgrade power system solar panels. |
1686 | 2022-02-24 | 18:21 | KLNB | Service Visit | Due to bad seismic data. |
1687 | 2022-02-24 | 18:22 | KLNB | Comments | Flooded seismo, water level 5 cm above connector. Cleaned up, changed seismo cable, restarted. |
1688 | 2022-02-24 | 18:23 | KLNB | Seismo Check | Mass-Centered ok. |
1689 | 2022-02-24 | 18:26 | KLNB | Seismo Check | Calibrated OK. End of service visit. |
1690 | 2022-04-28 | 12:55 | ELGO | Closed | End of Operation. |
1691 | 2022-04-29 | 19:00 | HGVO | Closed | End of Operation. |
1692 | 2022-04-30 | 15:30 | PTCO | Closed | End of Operation. |
1693 | 2022-05-03 | 01:22 | LDGN | Closed | End of Operation. |
1694 | 2022-05-04 | 14:20 | LINO | Closed | End of Operation. |
1695 | 2022-05-13 | 12:52 | MGCB | Closed | End of Operation. |
1696 | 2022-05-14 | 10:26 | TWKB | Closed | End of Operation. |
No | Date | Time UT | Location | Activity | Comments |
1697 | 2021-06-02 | 18:05 | OTTAWA | Comments | Activity File moved to new home. |
1698 | 2021-06-16 | 17:11 | SSIB | Closed | End of Operation. |
1699 | 2021-06-18 | 23:50 | TSJB | Closed | End of Operation. |
1700 | 2021-06-20 | 18:06 | TWBB | Closed | End of Operation. |
1701 | 2021-06-22 | 15:55 | KLNB | Closed | End of Operation. |
1702 | 2021-07-04 | 19:09 | SILB | Closed | End of Operation. |
1703 | 2021-07-09 | 15:59 | YNEN | Closed | End of Operation. |
1704 | 2021-07-10 | 14:48 | YMBN | Closed | End of Operation. |
1705 | 2021-07-11 | 14:40 | ACKN | Closed | End of Operation. |
1706 | 2021-07-14 | 14:04 | EKTN | Closed | End of Operation. |
1707 | 2021-07-17 | 14:21 | DVKN | Closed | End of Operation. |
1708 | 2021-07-17 | 16:25 | LGSN | Closed | End of Operation. |
1709 | 2021-07-21 | 23:30 | JERN | New Station Code | LUPN moved to this location. Station started with same hardware but on Telesat. |
1710 | 2021-07-22 | 01:05 | JERN | Seismo Check | Mass-center and calibration ok. |
1711 | 2021-07-24 | 13:58 | JERN | Gain Check | Gain is set to 1.0 |
1712 | 2021-07-24 | 13:58 | JERN | Hardware List | CYG389 TRD437 CMG-3ESP T3E89. |
No | Date | Time UT | Location | Activity | Comments |
1713 | 2021-07-24 | 13:59 | JERN | DC Filter Check | Off. |
1714 | 2021-07-24 | 13:59 | JERN | Require GPS Lock | On. |
1715 | 2021-07-24 | 13:59 | JERN | VSAT Check | Isolation 30 dB. On Telesat Hub 9. |
1716 | 2021-07-24 | 14:00 | JERN | Comments | Operation at 40 SPS. |
1717 | 2021-07-25 | 17:34 | GBLN | Service Visit | Site was down but nothing was broken. Feed horn was replaced. |
1718 | 2021-07-25 | 17:40 | GBLN | Comments | Raining at the site. |
1719 | 2021-07-25 | 17:40 | GBLN | SSPB Check | New SSPB is A0176A. |
1720 | 2021-07-25 | 17:40 | GBLN | VSAT Check | 25/0 dB. |
1721 | 2021-07-25 | 17:41 | GBLN | Comments | Sensor calibration ok. |
1722 | 2021-07-31 | 15:05 | O50T | Hardware List | Cyg 665, Trd 747. |
1723 | 2021-07-31 | 15:05 | O50T | New Station Code | New test station started for MACE replacement. |
1724 | 2021-08-02 | 14:00 | OTTAWA | Hub Swap | Linux firewall replaced with a pair of SonicWall Pro 300 firewalls. |
1725 | 2021-08-04 | 17:05 | PLIO | SSPB Check | Damaged sspb repaired. |
1726 | 2021-08-04 | 17:05 | PLIO | Service Visit | To restore dead station. |
1727 | 2021-08-04 | 17:10 | PLIO | Hardware Swap | New SSPB is A0165A, LNB 12200294. |
1728 | 2021-08-04 | 17:20 | PLIO | VSAT Check | 26/0. |
No | Date | Time UT | Location | Activity | Comments |
1729 | 2021-08-04 | 17:30 | PLIO | Seismo Check | Seismo is about 15 cm above water level. |
1730 | 2021-08-04 | 17:40 | PLIO | Comments | First site serviced by Bicyle (KA, BD). |
1731 | 2021-08-12 | 22:37 | ARVN | New Station Start | Started new station at Arviat, HB, Ottawa Hub Naqs2, Telesat Carina 116. |
1732 | 2021-08-12 | 22:38 | ARVN | VSAT Check | 31 db isolation. |
1733 | 2021-08-12 | 22:39 | ARVN | Hardware List | Same as DVKN, Cyg243, Trd239, TS 056, 40T/T4187. |
1734 | 2021-08-12 | 22:40 | ARVN | Comments | Libra VSAT IP is 10.15.255.243 |
1735 | 2021-08-12 | 22:41 | ARVN | DC Filter Check | Off. |
1736 | 2021-08-12 | 22:41 | ARVN | Gain Check | Gain is 1.0 at 40 sps. |
1737 | 2021-08-12 | 22:42 | ARVN | Carina Hubs | Both Cyg and Carina needed Multicast Forwarding at Naqs to work. |
1738 | 2021-08-12 | 22:42 | ARVN | Require GPS Lock | On. |
1739 | 2021-08-12 | 22:44 | ARVN | Comments | This is an AC station. |
1740 | 2021-08-12 | 22:56 | ARVN | New Station Start | Started new station at Arviat, HB, Ottawa Hub Naqs1. |
1741 | 2021-08-12 | 22:57 | ARVN | Comments | Naqs1 dataServer restarted. |
1742 | 2021-08-17 | 01:58 | BULN | New Station Start | Bullion, Hudson Bay, starting Ottawa Hub Naqs1, Naqs2. |
1743 | 2021-08-17 | 01:59 | BULN | Hardware List | Cyg246, Trd235, TS105, ESP T3A86, SSPB H013A, redeployed from YMBN. |
1744 | 2021-08-17 | 02:01 | BULN | Battery Check | Only 11 batteries (new) were available. |
No | Date | Time UT | Location | Activity | Comments |
1745 | 2021-08-17 | 02:01 | BULN | VSAT Check | 31 dB Isolation, Telesat, Carina 116. |
1746 | 2021-08-17 | 02:02 | BULN | DC Filter Check | Off. |
1747 | 2021-08-17 | 02:02 | BULN | Gain Check | Gain 1, sp 40 sps. |
1748 | 2021-08-17 | 02:02 | BULN | Require GPS Lock | On. |
1749 | 2021-08-17 | 02:03 | BULN | Site Geology | 200 m off runway. |
1750 | 2021-08-18 | 01:43 | STLN | New Station Start | Stellar Camp, Hudson Bay, Ottawa Naqs2, Telesat. |
1751 | 2021-08-18 | 01:44 | STLN | Hardware List | Cyg 242, Trd 183, TS 138, ESP T3A67, redeployed LGSN. |
1752 | 2021-08-18 | 01:45 | STLN | New Station Start | Stellar Camp, Hudson Bay, Ottawa Naqs1. |
1753 | 2021-08-18 | 01:45 | STLN | VSAT Check | 36 dB isolation! |
1754 | 2021-08-18 | 01:49 | STLN | Comments | Dataserver restarted on Naqs1. |
1755 | 2021-08-18 | 12:33 | STLN | DC Filter Check | Off. |
1756 | 2021-08-18 | 12:33 | STLN | Gain Check | Gain of 1 at 40 sps. |
1757 | 2021-08-18 | 12:33 | STLN | Require GPS Lock | On. |
1758 | 2021-08-18 | 15:05 | STLN | Site Geology | On Esker deposits, 1 Km from camp, south of airport. |
1759 | 2021-08-18 | 15:06 | STLN | Battery Check | 12 New battereies. |
1760 | 2021-08-18 | 15:07 | STLN | Hardware List | SSPB A0028A. |
No | Date | Time UT | Location | Activity | Comments |
1761 | 2021-08-20 | 20:36 | WAGN | New Station Code | Wager Bay, Hudson Bay, Naqs2, Ottawa Hub, Telesat. |
1762 | 2021-08-20 | 20:47 | WAGN | Hardware List | Cyg232, TS188, Trd179, ESP T3A89, SSPB H0017A, redeployed from YNEN. |
1763 | 2021-08-20 | 20:48 | WAGN | DC Filter Check | Off. |
1764 | 2021-08-20 | 20:48 | WAGN | Gain Check | 1.0 at 40 sps. |
1765 | 2021-08-20 | 20:48 | WAGN | Require GPS Lock | On. |
1766 | 2021-08-20 | 20:48 | WAGN | VSAT Check | 30 dB isolation on Telesat. |
1767 | 2021-08-20 | 20:49 | WAGN | Battery Check | 16 new batteries, 12 solar panels. |
1768 | 2021-08-20 | 20:49 | WAGN | Site Geology | On Esker near small stream 50 m off runway. |
1769 | 2021-08-20 | 20:50 | WAGN | New Station Code | Wager Bay, Hudson Bay, Naqs1, Ottawa Hub, Telesat. |
1770 | 2021-08-20 | 20:54 | WAGN | Comments | Dataserver started on Naqs1. |
1771 | 2021-08-21 | 11:59 | WAGN | Comments | Site location is at Sila Lodge air strip, Wager Bay. |
1772 | 2021-08-23 | 16:53 | MALG | Comments | Switched from SES-AM to Telesat. |
1773 | 2021-08-23 | 16:54 | MALG | Comments | F-Connector repaired on LNB. |
1774 | 2021-08-23 | 16:54 | MALG | VSAT Check | 34.5 dB isolation. |
1775 | 2021-08-23 | 17:30 | SNO | New Station Start | First Taurus Internet station started at SNO Lab, Sudbury. |
1776 | 2021-08-24 | 15:42 | SEDN | New Station Code | Sedna, Hudson Bay, on Ottawa Hub Naqs2, Telesat. |
No | Date | Time UT | Location | Activity | Comments |
1777 | 2021-08-24 | 15:52 | SEDN | New Station Code | Sedna, Hudson Bay, on Ottawa Hub Naqs1, Telesat. |
1778 | 2021-08-24 | 17:08 | SEDN | Hardware List | Cyg245, Trd423, TS190, ESP T3A90, SSPB 1645-0503, redeployed from LDGN. |
1779 | 2021-08-24 | 17:09 | SEDN | Site Geology | Bedrock at the end of runway strip. |
1780 | 2021-08-24 | 17:10 | SEDN | DC Filter Check | Off. |
1781 | 2021-08-24 | 17:10 | SEDN | Gain Check | 1.0 at 40 sps. |
1782 | 2021-08-24 | 17:10 | SEDN | Require GPS Lock | On. |
1783 | 2021-08-24 | 17:11 | SEDN | Battery Check | 16 new batteries, 12 solar panels. |
1784 | 2021-08-24 | 17:11 | SEDN | VSAT Check | 32 dB isolation. |
1785 | 2021-08-24 | 18:10 | GBN | Comments | Switched from SES-AM to Telesat. |
1786 | 2021-08-24 | 18:11 | GBN | VSAT Check | At least 28 dB isolation, very strong signal. |
1787 | 2021-08-24 | 20:06 | FLLB | New Station Code | Fletcher Lake Big Creek, Nechako Basin, BC, Ottawa Hub Naqs1,2, Telesat. |
1788 | 2021-08-24 | 20:07 | FLLB | Hardware List | Cyg266, Trd76, 3ESP T3D84, redeploed from TSJB. |
1789 | 2021-08-24 | 20:08 | FLLB | DC Filter Check | Off. |
1790 | 2021-08-24 | 20:08 | FLLB | Gain Check | Gain of 1 at 100 sps. |
1791 | 2021-08-24 | 20:09 | FLLB | Battery Check | 6 new batteries, 8 solar panels. |
1792 | 2021-08-24 | 20:09 | FLLB | Require GPS Lock | On. |
No | Date | Time UT | Location | Activity | Comments |
1793 | 2021-08-24 | 20:09 | FLLB | Site Geology | River bed deposits. |
1794 | 2021-08-24 | 20:09 | FLLB | VSAT Check | 29.5 dB isolation. |
1795 | 2021-08-24 | 20:10 | FLLB | Cold-Boot | Was done before all this. |
1796 | 2021-08-25 | 11:50 | JOSN | New Station Code | Second Sedna station, started at Ottawa Hub, Naqs1,2, Telesat. |
1797 | 2021-08-25 | 11:52 | JOSN | Hardware List | Cyg392, Trd141, TS 373, 40T T47773, SSPB A0208A 1W. |
1798 | 2021-08-25 | 11:54 | JOSN | DC Filter Check | Off. |
1799 | 2021-08-25 | 11:54 | JOSN | Gain Check | 1.0 at 40 sps. |
1800 | 2021-08-25 | 11:54 | JOSN | Require GPS Lock | On. |
1801 | 2021-08-25 | 11:55 | JOSN | Comments | All hardware except for Cygnus redeployed from EKTN. |
1802 | 2021-08-25 | 11:55 | JOSN | VSAT Check | 31 dB isolation. |
1803 | 2021-08-25 | 17:34 | SABG | Comments | Switched from SES-AM to Telesat, Carina 116. |
1804 | 2021-08-25 | 17:35 | SABG | VSAT Check | Signal from this 1.2 m antenna is strong at ERIP increase of 0 dB. |
1805 | 2021-08-27 | 00:48 | THMB | New Station Code | Started Thunder Mountain, BC, Ottawa Hub Naqs1,2 on Telesat. |
1806 | 2021-08-27 | 00:49 | THMB | Hardware List | Cyg271, Trd256, TS230, ESP T3D79, SSPB A0159A. |
1807 | 2021-08-27 | 00:51 | THMB | Battery Check | 6 new batteries, 8 solar panels. |
1808 | 2021-08-27 | 00:51 | THMB | DC Filter Check | Off. |
No | Date | Time UT | Location | Activity | Comments |
1809 | 2021-08-27 | 00:51 | THMB | Gain Check | Gain of 1.0 at 100 sps. |
1810 | 2021-08-27 | 00:51 | THMB | Require GPS Lock | On. |
1811 | 2021-08-27 | 00:51 | THMB | VSAT Check | 28 dB isolation. |
1812 | 2021-08-27 | 00:52 | THMB | Comments | Initial problem with GPS, long wait before fine lock. |
1813 | 2021-08-27 | 00:52 | THMB | Site Geology | Soft soil. |
1814 | 2021-08-27 | 00:53 | THMB | Comments | Bad seismometer cable, replaced with good one. |
1815 | 2021-08-27 | 21:12 | YBKN | New Station Code | Baker Lake, Hudson Bay, Ottawa Hub, Naqs1,2, Telesat. |
1816 | 2021-08-27 | 21:13 | YBKN | Hardware List | Cyg283, Trd936, TS187, ESP T3B15, SSPB, 1W H0026A. |
1817 | 2021-08-27 | 21:14 | YBKN | Gain Check | 1.0 at 40 sps. |
1818 | 2021-08-27 | 21:15 | YBKN | DC Filter Check | Off. |
1819 | 2021-08-27 | 21:15 | YBKN | Require GPS Lock | On. |
1820 | 2021-08-27 | 21:15 | YBKN | Site Geology | Gravel, airport area. |
1821 | 2021-08-27 | 21:15 | YBKN | VSAT Check | 27 dB Isolation. |
1822 | 2021-08-27 | 21:16 | YBKN | Battery Check | 2 DC-type batteries used in this AC site. |
1823 | 2021-08-27 | 21:17 | YBKN | Comments | Dish in low-lying area, 20 m away from Env. Canada Bldg. |
1824 | 2021-08-27 | 21:17 | YBKN | Power Check | AC Power to be connected in a day or so. |
No | Date | Time UT | Location | Activity | Comments |
1825 | 2021-08-28 | 00:54 | TALB | New Station Code | Tatla Lake, BC, Ottawa Hub, Naqs1,2, Telesat. |
1826 | 2021-08-28 | 00:55 | TALB | Hardware List | Cyg285, Trd83, TS194, 3ESP T3E02, SSPB 1W A0124A. |
1827 | 2021-08-28 | 00:56 | TALB | DC Filter Check | Off. |
1828 | 2021-08-28 | 00:56 | TALB | Gain Check | 1.0 at 100 sps. |
1829 | 2021-08-28 | 00:57 | TALB | Battery Check | 6 new batteries, 8 solar panels. |
1830 | 2021-08-28 | 00:57 | TALB | Require GPS Lock | On. |
1831 | 2021-08-28 | 00:57 | TALB | Site Geology | Soft soil |
1832 | 2021-08-28 | 00:57 | TALB | VSAT Check | 28 dB isolation. |
1833 | 2021-08-28 | 14:32 | CODG | Feed Horn Box | Replaced. |
1834 | 2021-08-28 | 14:32 | CODG | Service Visit | To repaire animal damage. |
1835 | 2021-08-28 | 14:33 | CODG | Cables | LNB cable repaired, SSPB cable replaced. |
1836 | 2021-08-28 | 14:33 | CODG | Carina Swap | Switched from SES-AM to Telesat, Ottawa Hub, Carina 116. |
1837 | 2021-08-28 | 14:34 | CODG | VSAT Check | 39 dB isolation!, running hot. |
1838 | 2021-08-29 | 19:10 | CHEG | Service Visit | All ok, solar panel rack leg fixed. |
1839 | 2021-08-29 | 19:11 | CHEG | SSPB Check | Profile adjusted from 45 to 50 dB flat, EIRP increase adjusted from -5 to 0 to balance. |
1840 | 2021-08-29 | 19:11 | CHEG | VSAT Check | 31.5 dB isolation. |
No | Date | Time UT | Location | Activity | Comments |
1841 | 2021-08-29 | 19:12 | CHEG | Comments | Site in very good condition, bolts on solar panel secured. |
1842 | 2021-08-30 | 00:00 | ARVN | Comments | Station moved by about one km to new locatoin for beeter AC connection. Old data to be junked, station code to be kept the same. |
1843 | 2021-08-30 | 01:32 | ALRB | New Station Code | Anahim Lake, Lehmann Ranch, BC, recording at Ottawa Hub, Naqs1,2, Telesat. |
1844 | 2021-08-30 | 01:36 | ALRB | Hardware List | Cyg192, Trd91, TS37, SSPB A0177A, ESP T3B25. |
1845 | 2021-08-30 | 01:37 | ALRB | DC Filter Check | Off. |
1846 | 2021-08-30 | 01:37 | ALRB | Gain Check | 1.0 at 100 sps. |
1847 | 2021-08-30 | 01:38 | ALRB | Require GPS Lock | On. |
1848 | 2021-08-30 | 01:38 | ALRB | VSAT Check | 29.2 dB isolation. |
1849 | 2021-08-30 | 01:39 | ALRB | Comments | Site installed without analyzer, helped by Telesat. |
1850 | 2021-08-30 | 20:58 | TIGG | Carina Swap | Switched from SES-AM to Telesat at Ottawa Hub. |
1851 | 2021-08-30 | 20:59 | TIGG | SSPB Check | Gain adjusted to 50 dB, EIRP adjusted to match. |
1852 | 2021-08-30 | 20:59 | TIGG | VSAT Check | 28 dB isolation while raining at the site. |
1853 | 2021-08-31 | 11:57 | LAIN | Comments | Unable to do scheduled service. Trident and /or seismometer are broken. No seismic data for months. |
1854 | 2021-08-31 | 12:02 | LAIN | TDMA Halt | Removed from TDMA (Cygnus 261, carina 115) until next visit. Station down. |
1855 | 2021-08-31 | 12:17 | MGTN | Comments | Unable to do scheduled service. Station has been down since 2005. |
1856 | 2021-08-31 | 12:18 | MGTN | TDMA Halt | Removed from TDMA (Cygnus 559, Carina 62) until next visit. Station down. |
No | Date | Time UT | Location | Activity | Comments |
1857 | 2021-08-31 | 21:07 | BATG | Carina Swap | Switched from SES-AM to Telesat on Ottawa Hub. |
1858 | 2021-08-31 | 21:08 | BATG | Hardware Swap | Cygnus 422 installed, 605 removed. |
1859 | 2021-08-31 | 21:09 | BATG | VSAT Check | 30 dB. |
1860 | 2021-08-31 | 21:15 | BATG | Comments | Polarization was off by 90 deg during Telesat move, resulted in confusion. Fixed. |
1861 | 2021-09-01 | 22:00 | CLSB | New Station Code | Cack Lake, Smith's Ranch, SW of Quesnel, BC. Started on Ottawa Hub, Telesat. |
1862 | 2021-09-02 | 14:38 | CLSB | Hardware List | Cyg273, Trd425, TS170, ESP T3B26, SSPB H0020A. |
1863 | 2021-09-02 | 14:39 | CLSB | VSAT Check | 30 dB, strong signal on Telesat. |
1864 | 2021-09-02 | 14:40 | CLSB | DC Filter Check | Off. |
1865 | 2021-09-02 | 14:40 | CLSB | Gain Check | 1.0 at 100 sps. |
1866 | 2021-09-02 | 14:40 | CLSB | Require GPS Lock | On. |
1867 | 2021-09-02 | 14:41 | CLSB | Comments | GPS slow to connect. Bad seismo spotted, replaced. |
1868 | 2021-09-02 | 14:41 | CLSB | Site Geology | Packed clay, trees nearby. |
1869 | 2021-09-03 | 21:04 | RAMB | New Station Code | South West of Quesnel, Ramsey Rd, BC, Ottawa Hub, Telesat. |
1870 | 2021-09-03 | 21:05 | RAMB | Hardware List | Cyg182, Trd63, TS53, seismometer 40T, s/n T4769, sspb A0160A. |
1871 | 2021-09-03 | 21:06 | RAMB | DC Filter Check | Off. |
1872 | 2021-09-03 | 21:06 | RAMB | Gain Check | 1.0 at 100 sps. |
No | Date | Time UT | Location | Activity | Comments |
1873 | 2021-09-03 | 21:07 | RAMB | Battery Check | 6 new batteries, 8 panels. |
1874 | 2021-09-03 | 21:07 | RAMB | Require GPS Lock | On. |
1875 | 2021-09-03 | 21:07 | RAMB | VSAT Check | 31 db isolation. |
1876 | 2021-09-03 | 21:08 | RAMB | Site Geology | Silt. |
1877 | 2021-09-05 | 12:46 | SULB | New Station Code | South of Vanderhoof, BC, Ottawa Hub, Telesat. |
1878 | 2021-09-05 | 12:47 | SULB | Hardware List | Cyg219, Trd080, TS090, SSPB A0156A, 3ESP T3D96. |
1879 | 2021-09-05 | 12:49 | SULB | DC Filter Check | Off. |
1880 | 2021-09-05 | 12:49 | SULB | Gain Check | 1.0 at 100 sps. |
1881 | 2021-09-05 | 12:49 | SULB | Require GPS Lock | On. |
1882 | 2021-09-05 | 12:50 | SULB | Site Geology | Gravel pit. |
1883 | 2021-09-05 | 12:50 | SULB | VSAT Check | 27.5 dB isolation on Telesat. |
1884 | 2021-09-05 | 12:51 | SULB | Comments | This is the 7th and last of the Nechako Basin. |
1885 | 2021-09-05 | 14:55 | SULB | Comments | Trident SN in Remote Cygnus Ringbuffer corrected to TRD080. |
1886 | 2021-09-06 | 14:43 | MADG | Carina Swap | Switched from SES-AM to Telesat, Carina 92, London to Ottawa Hub. |
1887 | 2021-09-06 | 14:44 | MADG | Comments | Corrosion problem, feed horn need to be replaced on next service. |
1888 | 2021-09-06 | 14:44 | MADG | VSAT Check | 30 dB isolation, strong signl. |
No | Date | Time UT | Location | Activity | Comments |
1889 | 2021-09-06 | 14:45 | MADG | Comments | Cell phone coverage from site, but poor line. |
1890 | 2021-09-06 | 16:58 | ALRB | Service Visit | To fix bus gap problems. |
1891 | 2021-09-06 | 16:59 | ALRB | Comments | Bus and grounding cables checked, all ok. |
1892 | 2021-09-06 | 16:59 | ALRB | Gain Check | 1.0 at 100 sps, as before. |
1893 | 2021-09-06 | 16:59 | ALRB | Hardware Swap | Trident 91 removed, Trd695 installed. |
1894 | 2021-09-06 | 17:00 | ALRB | Comments | This is the end of Nechako basin installation/service. |
1895 | 2021-09-06 | 17:00 | ALRB | DC Filter Check | Off, as before. |
1896 | 2021-09-06 | 17:00 | ALRB | Require GPS Lock | On, as before. |
1897 | 2021-09-07 | 17:52 | OTTAWA | Comments | Data directories on drive D: on Naqs Servers 1 and 2 cleaned up. Interent and Test site data ringbuffers moved to drive C: due to lack of space on D: |
1898 | 2021-09-07 | 17:54 | OTTAWA | Comments | SOH ringbuffers for Internet and Test sites remain on drive D:. |
1899 | 2021-09-07 | 18:27 | BULN | VSAT Check | IFL loss faked to 4.0 db to increase power. EIRP inc is already at max. 10 dB. This improved performance. |
1900 | 2021-09-08 | 18:00 | GASG | Service Visit | Routine maintenance. |
1901 | 2021-09-08 | 18:10 | GASG | VSAT Check | Repointing gained 2 dB, isolation is 40 dB. |
1902 | 2021-09-08 | 22:00 | OTTAWA | Comments | Storm in Ottawa caused temporary outage on GBLN, GLWN, ILKN, MCKN, SNPN and SILO. Indicating TX power on these is low? |
1903 | 2021-09-10 | 16:44 | NATG | Service Visit | To fix bad VSAT. |
1904 | 2021-09-10 | 16:45 | NATG | Hardware Swap | LNB replaced before bad Cygnus was spotted. |
No | Date | Time UT | Location | Activity | Comments |
1905 | 2021-09-10 | 16:45 | NATG | Hardware Swap | TimeServer problem with Cygnus 161. Replaced with Cygnus 605. |
1906 | 2021-09-10 | 16:46 | NATG | Seismo Check | Calibrated OK. |
1907 | 2021-09-10 | 16:46 | NATG | VSAT Check | 32.5 dB, 2 dB improvement. |
1908 | 2021-09-10 | 16:47 | NATG | Comments | VSAT bolts to rock were ok but secured. |
1909 | 2021-09-12 | 20:00 | INUQ | New Station Code | Internet, Ottawa Hub, at Inukjuak, Hudson Bay. |
1910 | 2021-09-12 | 20:05 | INUQ | Internet Station | On VPN. |
1911 | 2021-09-12 | 20:10 | INUQ | Hardware List | Taurus 270, ESP T3B16, IP 86.34 |
1912 | 2021-09-12 | 20:20 | INUQ | 40 sps, Gain 1, No DC filter, Require GPS Lock | Checked. |
1913 | 2021-09-15 | 14:00 | AKVQ | New Station Code | Internet, Ottawa Hub, at Akulivik, Hudson Bay. |
1914 | 2021-09-15 | 14:05 | AKVQ | Internet Station | On VPN. |
1915 | 2021-09-15 | 14:10 | AKVQ | Hardware List | Taurus 192, ESP T3D60, IP 86.18. |
1916 | 2021-09-15 | 14:20 | AKVQ | 40 sps, Gain 1, No DC filter, Require GPS Lock | Checked. |
1917 | 2021-09-15 | 14:30 | AKVQ | Comments | Internet TX problems, gaps in data. |
1918 | 2021-09-16 | 22:00 | IVKQ | New Station Code | Internet, Ottawa Hub, at Ivujivik, Hudson Bay. |
1919 | 2021-09-16 | 22:10 | IVKQ | Hardware List | Taurus 201, ESP T3P18. |
1920 | 2021-09-16 | 22:20 | IVKQ | 40 sps, Gain 1, No DC filter, Require GPS Lock | Checked. |
No | Date | Time UT | Location | Activity | Comments |
1921 | 2021-09-16 | 22:25 | IVKQ | Internet Station | On VPN. |
1922 | 2021-09-20 | 04:00 | SNQN | New Station Code | Internet, Ottawa Hub, at Sanikiliuq, Hudson Bay. |
1923 | 2021-09-20 | 04:10 | SNQN | Hardware List | Taurus 207, ESP T3P17. |
1924 | 2021-09-20 | 04:20 | SNQN | 40 sps, Gain 1, No DC filter, Require GPS Lock | Checked. |
1925 | 2021-09-20 | 04:25 | SNQN | Internet Station | On VPN. |
1926 | 2021-10-03 | 13:50 | RLKO | Service Visit | TimeServer failure, both Cygnus and Trident replaced. |
1927 | 2021-10-04 | 13:50 | RLKO | VSAT Check | X-pol 29 dB. |
1928 | 2021-10-05 | 14:00 | NANO | Service Visit | Vandalism, station turned off. Electronics replaced. |
1929 | 2021-10-05 | 14:10 | NANO | VSAT Check | 30 dB xpol. |
1930 | 2021-10-13 | 18:00 | ELFO | Internet Station | Converted to Taurus Internet at old location, served by Ottawa Hub. |
1931 | 2021-10-13 | 18:10 | ELFO | Hardware List | Taurus 121, ESP T3J88 (as before). |
1932 | 2021-10-13 | 18:20 | ELFO | Gain Change | Gain is 1.0 (was 0.5 before) at 100 sps. |
1933 | 2021-10-13 | 18:30 | ELFO | DC Filter Check | On at 0.001 Hz. |
1934 | 2021-10-13 | 18:40 | ELFO | Require GPS to Lock | On. |
1935 | 2021-10-13 | 18:50 | ELFO | Internet Station | On VPN. |
1936 | 2021-10-18 | 16:00 | PLVO | Service Visit | To connect MT station. |
No | Date | Time UT | Location | Activity | Comments |
1937 | 2021-10-18 | 16:10 | PLVO | Comments | MT battery box and recorder missing. |
1938 | 2021-10-18 | 16:20 | PLVO | VSAT Check | 25/0 dB, improved by 3.5 dB. |
1939 | 2021-10-26 | 19:36 | PLVO | MT Station | Started MT recording, live to London Hub. |
1940 | 2021-10-27 | 12:51 | ARVN | TDMA Halt | Removed from TDMA (CYG243, Carina 115) until next visit. Station down due to antenna shift. |
1941 | 2021-10-27 | 12:56 | A3PN | TDMA Halt | Removed from TDMA (CYG253, Carina 115) until next visit. Station down due to antenna shift. |
1942 | 2021-10-31 | 01:00 | LONDON HUB | Hub Swap | All stations moved from Carina 28 to Carina30. Carina 28 powered down. |
1943 | 2021-10-31 | 15:10 | LONDON | Hub Swap | All stations but TOBO moved from Carina 29 to Carina30 and 31. |
1944 | 2021-11-01 | 19:27 | PLIO | MT Station | Started MT recording, live to London Hub. |
1945 | 2021-11-02 | 21:42 | CLWO | MT Station | Started MT recording, live to London Hub. |
1946 | 2021-11-04 | 12:17 | TOBO | Hub Swap | Finished hub swap. TX disabled on carina 28 and 29. |
1947 | 2021-11-09 | 17:15 | DSNO | Internet Station | Started at deep SNO, Ottawa Hub Naqs1,2. |
1948 | 2021-11-09 | 17:15 | SSNO | Internet Station | Started at shallow SNO, Ottawa Hub Naqs1,2. |
1949 | 2021-11-09 | 17:16 | DSNO | Hardware List | Taurus 294. Trillium 120P s/n to be reported. |
1950 | 2021-11-09 | 17:16 | SSNO | Hardware List | Taurus 476. Trillium 120P s/n to be reported. |
1951 | 2021-11-09 | 17:17 | DSNO | Gain Check | Gain 1.0 at 200 sps. |
1952 | 2021-11-09 | 17:17 | SSNO | Gain Check | Gain 1.0 at 200 sps. |
No | Date | Time UT | Location | Activity | Comments |
1953 | 2021-11-09 | 17:18 | DSNO | DC Filter Check | Off. |
1954 | 2021-11-09 | 17:18 | SSNO | DC Filter Check | Off. |
1955 | 2021-11-09 | 17:19 | DSNO | Require GPS to Lock | Off. No GPS at this site. |
1956 | 2021-11-09 | 17:19 | SSNO | Require GPS to Lock | on. |
1957 | 2021-12-04 | 00:00 | O66T | Comments | Started long-term test of Trillium 120P in Ottawa insulated vault. |
1958 | 2021-12-06 | 14:38 | O66T | Comments | Trident moved into the insulated cavity near seismometer to register seismometer temperature more accurately. |
1959 | 2021-12-06 | 14:38 | O66T | Hardware List | Trillium s/n is 121. |
1960 | 2021-12-06 | 19:51 | OTTAWA HUB | VSAT Check | SES-AM Antenna steady at 29.5 dB x-pol. |
1961 | 2021-12-06 | 19:53 | OTTAWA | VSAT Check | Bad burst Rx may continue on Telesat, even after repoint. |
1962 | 2021-12-06 | 19:53 | OTTAWA | VSAT Check | Telesat Antenna co-pol improved by 0.5 dB, isolation is now 32.5 dB. |
1963 | 2021-12-11 | 17:00 | PLVO | MT Station | North e-line was cut by animale bite. Added 25' splice, soldered, heat tape applied and bagged in silicon. |
1964 | 2022-01-01 | 00:00 | OTTAWA | MT Station | New Scaling corrections started for Ontario MT stations. |
1965 | 2022-02-06 | 18:00 | OTTAWA | Comments | Naqs1 and Naqs2 server MTU restored to 1500. |
1966 | 2022-02-15 | 17:15 | OTTAWA | Battery Check | All 10 UPS batteries were replaced. |
1967 | 2022-02-15 | 17:20 | OTTAWA | Comments | Inverter glitch due to power disconnect. |
1968 | 2022-03-15 | 17:34 | OTTAWA | VSAT Check | Satmex antenna re-point, x-pol: before 25, now 32 dB. |
No | Date | Time UT | Location | Activity | Comments |
1969 | 2022-04-02 | 15:04 | MPPO | Service Visit | Charge controller was broken, replaced. |
1970 | 2022-04-02 | 15:06 | MPPO | VSAT Check | 30/0 dB. |
1971 | 2022-04-02 | 15:10 | MPPO | Comments | Mass-center and calibration ok. |
1972 | 2022-04-04 | 18:30 | MRHQ | Service Visit | Needed repoint. |
1973 | 2022-04-04 | 18:35 | MRHQ | VSAT Check | 26/0 dB. |
1974 | 2022-04-04 | 18:36 | MRHQ | SSPB Check | New 2 W SSPB. |
1975 | 2022-04-04 | 18:37 | MRHQ | Power Check | LVD added to AC power system. |
1976 | 2022-04-04 | 18:40 | MRHQ | Comments | Snowing heavily at the site. |
1977 | 2022-04-05 | 17:25 | O66T | VSAT Check | Needed repoint. Now 29/0 dB, improved by 10 dB. |
1978 | 2022-04-05 | 17:30 | O66T | Comments | Testing Trillium 120P here, antenna movement consistent with 120P showing effect of tilt. |
1979 | 2022-04-10 | 17:00 | LONDON HUB | VSAT Check | X-pol 30 dB, power improved by 2 dB. |
1980 | 2022-04-12 | 18:23 | O66T | Comments | Trident digitizer configuration corrected from +12V, active low, High, High to +5V, active High, Low, Low, Low to match Trillium seismometer. |
1981 | 2022-04-12 | 18:25 | O66T | Comments | Naqs Staion file at Ottawa corrected for Trident connected to Trillium. |
1982 | 2022-04-12 | 18:40 | SILO | VSAT Check | Changed modulation from 112 to 64KB to improve transmission. |
1983 | 2022-04-18 | 15:30 | OTTAWA-HUB | Comments | Cables, switches and Internet connection upgraded. 8 and 24 managed switches installed. |
1984 | 2022-05-02 | 18:00 | OTTAWA-HUB | VSAT Check | Satmex antenna re-point, x-pol: improved by 16 dB, now it is 35 dB. |
No | Date | Time UT | Location | Activity | Comments |
1985 | 2022-05-07 | 15:58 | OTTAWA-HUB | VSAT Check | Satmex antenna re-point, x-pol: before 26, now 32 dB. |
1986 | 2022-05-07 | 15:59 | LITE | VSAT Check | X-pol is off by 5 dB. Site visit is required. |
1987 | 2022-05-07 | 16:00 | LITE | Comments | Eirp Inc. changed from -2 to -7 dB. |
1988 | 2022-05-12 | 12:35 | OTTAWA-HUB | Comments | NPtoNMP software v 2.2.1 installed on Naqs1. |
1989 | 2021-06-11 | 21:26 | PLBC | Comments | Stopped POLARIS VSAT, restarted as CNSN. |
1990 | 2021-06-19 | 17:55 | ORHO | Internet Station | New internet station at Orleans (Heritage park) Ontario, Ottawa Naqs1. |
1991 | 2021-06-19 | 17:57 | ORHO | Internet Station | 100 sps, 40Vpp, Software Gain 1, No DC filter, Low Impedance. |
1992 | 2021-06-19 | 18:01 | ORHO | Hardware List | Taurus 267, Guralp T3, T3762. |
1993 | 2021-06-19 | 18:01 | ORHO | Require GPS to Lock | On. |
1994 | 2021-06-19 | 18:54 | ORHO | Comments | Station off-line until June 22. |
1995 | 2021-06-20 | 20:42 | DRWO | Internet Station | New station, Darlington West, Ontario, Ottawa Naqs1. |
1996 | 2021-06-20 | 20:43 | DRWO | Internet Station | 100 sps, No DC filter, 40 Vpp, Low Impedance, Software Gain 1. |
1997 | 2021-06-20 | 20:45 | DRWO | Hardware List | Taurus 1013, KS-2000M seismometer, S/N 569. |
1998 | 2021-06-20 | 20:46 | DRWO | Require GPS to Lock | On. GPS always on, No Alignment. |
1999 | 2021-06-20 | 20:47 | DRWO | Comments | Darlington Vault. |
2000 | 2021-06-22 | 19:51 | DREO | New Station Start | Darlington East, Libra VSAT recording at Ottawa Hub Naqs1. |
No | Date | Time UT | Location | Activity | Comments |
2001 | 2021-06-22 | 19:53 | DREO | VSAT Check | 34.5 dB isolation for Telesat. |
2002 | 2021-06-22 | 19:55 | DREO | Gain Check | Software gain 1.0, Vpp 40, at 100 sps. |
2003 | 2021-06-22 | 19:56 | DREO | DC Filter Check | Off. |
2004 | 2021-06-22 | 19:56 | DREO | Hardware List | Cygnus 802, TrD 1144, seismometer Geotech, 568. |
2005 | 2021-06-22 | 19:56 | DREO | Require GPS to Lock | On. |
2006 | 2021-06-22 | 19:58 | DREO | Comments | N43.87244 w 78.70396 309 FT |
2007 | 2021-06-22 | 19:59 | DREO | Comments | A0017498 sspb, 2 w |
2008 | 2021-06-22 | 20:01 | DREO | Power Check | 16 batteries, 12 solar panels. |
2009 | 2021-06-23 | 03:00 | ORHO | Restart | Station restarted after Internet connection was serviced. |
2010 | 2021-06-26 | 17:46 | FLLB | Comments | Mass-centered OK. |
2011 | 2021-06-26 | 17:46 | FLLB | Service Visit | Replacing bad seismometer. |
2012 | 2021-06-26 | 17:47 | FLLB | Cold-Boot | OK. |
2013 | 2021-06-26 | 18:21 | FLLB | Comments | Calibration OK. |
2014 | 2021-06-26 | 18:23 | FLLB | Hardware Swap | T3D84 removed, T3K45 installed. |
2015 | 2021-06-26 | 18:24 | FLLB | Hardware Swap | Seismometer cable changed due to condensation in connector. |
2016 | 2021-06-26 | 18:25 | FLLB | Comments | All is OK. |
No | Date | Time UT | Location | Activity | Comments |
2017 | 2021-06-26 | 18:25 | FLLB | Comments | Padlock changed. |
2018 | 2021-06-26 | 18:25 | FLLB | VSAT Check | 30 dB, no change. |
2019 | 2021-06-26 | 18:26 | FLLB | Gain Check | Gain 1 at 40 Vpp 100 sps. |
2020 | 2021-06-26 | 18:29 | FLLB | DC Filter Check | Off. |
2021 | 2021-06-27 | 14:57 | ORHO | Seismo Swap | T3762 (3T) removed, T3A81 (3ESP) installed. |
2022 | 2021-06-27 | 15:00 | ORHO | Seismo Check | Mass-centered OK. |
2023 | 2021-06-27 | 20:00 | THMB | Service Visit | To improve VSAT. |
2024 | 2021-06-27 | 20:01 | THMB | Seismo Check | Re-levelled. Cable connector cleaned up. |
2025 | 2021-06-27 | 20:01 | THMB | VSAT Check | 36 dB separation. |
2026 | 2021-06-27 | 20:03 | THMB | Comments | Mass-centered and calibrated OK. |
2027 | 2021-06-27 | 22:48 | ORHO | Comments | Auto Mass-centering is turned on, on Red at 4 V. |
2028 | 2021-07-09 | 12:50 | O66T | VSAT Check | 30 dB. Station is up. |
2029 | 2021-07-09 | 12:51 | O66T | Comments | Mag connected to this station. |
2030 | 2021-07-10 | 21:41 | KASO | Service Visit | Just to adjust VSAT. |
2031 | 2021-07-10 | 22:18 | KASO | Comments | Service continues. |
2032 | 2021-07-11 | 00:46 | KASO | VSAT Check | Isolation improved from 10 to 16 dB. |
No | Date | Time UT | Location | Activity | Comments |
2033 | 2021-07-11 | 00:47 | KASO | Comments | Feedhorn arms were upside down, corrected. |
2034 | 2021-07-11 | 00:49 | KASO | Comments | General condition of site was fine. |